A Bibliography of Publications in Computer Languages

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 January 2019 Version 2.37

Title word cross-reference

ifdefs [BGM^{+}18].

8_{1/2} [Mic96]. ^{+} [NL95]. C [MVT^{+}16]. k [AG17]. μ [CCRS18]. O(n) [BS92].

-DSU [CCRS18]. -inductive [AG17].

0 [Led99b, Led99c, Led99d, Led99e, Led99a]. 0-471-19233-3 [Led99b]. 0-471-19572-3 [Led99e]. 0-471-29549-3 [Led99c]. 0-471-31502-8 [Led99a]. 0-471-32734-4 [Led99d].

1 [Sul75]. 14 [HM18]. 16th [DL17].

2 [DNR90, Sal92]. 2-offset [CAS08]. 2001 [Ano01b]. 2011/12 [LP16]. 2013 [HM18].

2013/14 [HM18]. 2013/2014 [HB16]. 2015 [GAS17]. 25 [Ano99]. 26 [Ano00]. 27th [MB13]. 28th [MB14]. 29th [Bry15].

3 [Led99b, Led99c, Led99e]. 3-Tier [Led99a]. 30th [Bry16]. 370 [FF75]. 3rd [HB16].

4 [Led99d]. 4th [HB16].

6.0} [Led99c]. 67 [Sch78]. 6th [CPPV15].

7th [CPPV15].

8 [Led99a]. 80 [GL95].

’92 [CB93b]. 95 [GSX99].

ABC [AC18]. abductive [CLMT01]. Abstract [Bai87, DL17, GZ87, HC12, BZ88, CZ11, CCF15, FW87, Jal92, Liu93, Log09, McL77, Noo85, PC15, RK93, UM17]. Abstracting [HF87]. abstraction [KG17, ØK00, ZP04]. Abstractions [Coo81, SS79, BEL77, Ber77, DMVY17, DNR90]. Access [SC94, AMF13, KK18, DOZ06]. accessibility [CY02]. accumulative [Mor16]. Achieving [GHD18]. ACM [Bry15, Bry16, HB16, MB13, MB14]. acquired [SK18]. Action [DS93]. Active [YF98]. Activity [Sal92, ABS17]. actor [CHS16, DMVD16]. actors [HB16, Ric16, VMD09, VBS^{+}14]. acyclic [VS93]. Ada [BJS93, CMM85, EL87, EHMO91, GSX99, Geh82, MZGT85, SC94]. Ada-95 [GSX99]. ADABTPL [SS92]. Adaptable [RS83]. Adapting [RDT08]. adaptive [PPK11, RBY^{+}05]. adding [MZC10]. address [FAHC17]. advanced [GSX99]. affine [AG17]. affix [HSS88]. Agent [BIMP17, ACZ05, ABL17, KGS17, WHKK17]. Agent-oriented [BIMP17, ACZ05]. agents [CLMT01, HB16, Ric16]. AGERE [HB16]. aggregates [BCR11]. agile [WHKK17]. ahead [JPB^{+}08]. ahead-of-time [JPB^{+}08]. aid [ZP04]. algebra [BLM93, MH07]. Algebraic [RH94, GSP17, Rus87]. ALGOL [SB79, CHH02, NK90]. Algol-like [NK90, CHH02]. Algorithm [Man78, AC18, CAS08, Dha90, FNRR16, Noo85, Pai16, Yan00]. Algorithms [PB84, Sal83, War78, Ban17, PS10, SIK09]. Allocation [CAC^{+}81, BM95, LCC07, jLtCxH09, Zob93]. allocator [ÜA15, ÜA16]. allocators [HC05]. Altering [Cov93]. alternatives [GG09]. ALua [URI02]. Amake [Buf17]. Ambient [AKPG02]. ambients [BCF02, BCF^{+}04, BC02]. AmbientTalk [VBS^{+}14]. Among [Pet78, CLMT01]. analyses [BGH13, KHO14]. Analysis [Liu88, MM82, Ree84, Sha80, TSF^{+}87, Wad80, ALS^{+}18, ABL17, BC93, Ban17, BCF^{+}04, BGM^{+}18, BC10, CCB15, GDD12, Hor17, HV93, KDM03, LDG09, LR17, MCC17, MT05, OK18, Ozt11, Rid79b, RD78, SF89, YTC02, ZLZ^{+}18]. analyzability [HG93]. analyzer [ZGE85]. analyzers [Yan96]. Anatomy [Ree84]. AND- [HC96]. and-cross [ALS^{+}18]. AND-parallel [PGT^{+}96]. Android [CC15, CC16]. animation [Bay76]. Ann [CSdL16]. annotation [CV14, LDN18]. annotation-based [LDN18]. annotations [BC10, CSdL16, SNP16]. Announcement [Ano06a]. anomaly [MW96]. answer [PLS10]. Answering [KP78]. ant [PM18]. Ants [PM18]. APL [GFK81, SW77]. Apple [KN85]. applicability [YTC02]. Application [BKG^{+}08, CJ80, KJTA17, Orm83, OAB^{+}18, Sch78, BMN^{+}18, GAGdL17, KS90, PLDD15]. Application-specific [BKG^{+}08, PLDD15]. Applications [CMM85, Ken78, Led99b, Led99c, AA09, BDL^{+}12, CS16, CJD17, KKG92, Man01, MZC10, MGLFCP12, MP92, PLDD15, PJ91, RGP98, RDT08, VBS^{+}14, Zak88]. Applicative [GS86, Sal83]. Applied [Bry15, Bry16, MB13, MB14, Zav86]. Applying [DQ09, Cov93]. Approach [CCRS18, CJ80, Rid79a, Sha80, Zav86, ABS17, ABL17, Bas75, BMN^{+}18, CJD17, CC15, CC16, CO89, DTXP13, DCA^{+}15, DCA^{+}16, Guo16, Hoa75, Ier93, Lee05, LB06, MZC10, OAB^{+}18, PFH16, Rid79b, RK18, SM94]. Approaches [BBT15, BBT16]. Approximate [Spr79]. apps [CC15, CC16, RK18]. Arabic [AA89, AAH95, ZA87]. ARABLAN [AAH95]. architectural [SMB15, SMB16]. Architecture [Ano07a, KKG15, SK14]. architectures [CC95, OK18, PC15, VS95, VLC98, WMP^{+}08, vOKF01]. Argos [MR01]. arithmetic [PS94a]. Arity [Bre18]. Array [CPD93, JG89, LR17]. arrays [DK92, DLP07, Lus02]. artificial [AC18]. ary [CBTR18]. ascent [Hor93]. ascent-descent [Hor93]. Ash [Led99c]. asm.js [VSN^{+}17]. Aspect [PCB^{+}18]. Aspects [HH06, Was79, Alj16, DGU91]. Assembling [Tay96]. assertions [Jay92]. assessment [IS18]. Assignment [Sam79, Dha88, Dha90, SIK09]. assistance [ÁdLNW18]. assistants [AA09]. Associated [Fle84]. Associative [CRPP00]. Athena [ONB^{+}18]. atomic [DLP15, YF98]. Atomistic [NN17]. attaching [AA09]. attribute [CY02, DPP10, Geh77, MS89, Yan00]. attribute-grammar [CY02]. Attributed [CBTR18]. Attributes [Tai79, SH15]. augmenting [Li96]. Author [Ano05a, Ano05g, Ano06c]. automata [KG17, KB75, PTJM16]. automata-based [KG17]. Automated [ÁdLNW18, CBTR17, DKK18, GAGdL17, Guo16, KKG92]. Automatic [AG17, CYS^{+}15, HL08, IS18, LBGA18, MT82, Man01, BM95, CM11, DPP10, PNF^{+}18, SSB94, Wet77]. Automatically [BC89, Ear75]. automating [yCH92]. Automatized [SP18]. automaton [MR01]. automaton-based [MR01]. autonomous [DMT10]. AVX [HMHS18]. aware [BDL^{+}12, BGM^{+}18, SSS17]. Axiomatic [BEH86, Hoa75].

B [BMN^{+}18]. Babbage [Fri92]. Babel [Fri92]. back [SIK09]. back-translation [SIK09]. backtrack [Sar94]. backtracking [KPP93]. backward [LCFÁ10]. BaLinda [FWY96, YF98]. ball [Vai04]. banker [McK75]. base [McL77]. Based [BKL18, CCRS18, CLM83, GS86, AD07, ACZ05, AC17, ALR15, Ban17, BLM93, BSW15, Bou04, Bou08, BMN^{+}18, CDW09, CBTR17, CJD17, CCJ93, CG96, FBDH12, FL92, Gan89a, HB16, HSS88, HGC^{+}09, KG17, KPN17, LDN18, LR17, Mal17, MR01, MCC17, OAB^{+}18, OK18, PSW95, RR99, SRRB10, SRT17, SLS18, VMD18, Wan92, WPR06, WBGM10, Zak88, FAHC17, KS90].

Bases [BC84]. Basic [Led99c]. basis [JG89, Zob93]. bee [AC18]. beginners [Hug85, Mor16]. Behavior [Rid79a, Sar93, SJW94]. behavioral [KKP^{+}15, RDT08, Zdu06]. Behavioural [BC02]. benchmarks [EvdSV^{+}15]. Bergeon [Led99d]. best [HC05]. best-fit [HC05]. BETA [ØK00]. better [KY75, Yan96]. between [FBDH12, SSM10, VMD09]. beyond [Fri92]. Bidirectional [KDM03]. Binary [HT13, CYS^{+}15, MLW05]. Binding [Sam79, VF82]. bio [ABS17]. bio-inspired [ABS17]. biomedical [Zak88]. Black [Ber91b]. Black-box [Ber91b]. blend [GBZ09]. block [DKK18]. Blocks [Pag79]. Board [Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano05b, Ano05c, Ano05d, Ano09a, Ano10a, Ano18a, Ano18b, Ano18c]. bottom [BDB90]. bottom-up [BDB90]. bound [KJ12]. boundaries [BCF02, DKK18]. Bounded [KKNS14, KLIN15]. bounds [BJ14]. box [Ber91b]. BPEL [KJ12]. Branching [RGP98]. Branching-time [RGP98]. bridge [FBDH12]. Bridging [YD78]. brief [Fri92]. Bringing [CV14]. Broadcasting [Bro88, PS94b]. browser [SB04]. Buffering [Bro88]. build [CJD17]. Building [Led99b, Led99c, Li96]. builds [Buf17]. bulk [MH07]. bulk-synchronous [MH07]. business [LLvdW^{+}01, PLDD15, RK18]. bytecode [DDT06, JPB^{+}08]. bytecode-to-C [JPB^{+}08].

C [Ano88, Bud82, CHS16, CL89, EP89, ECB12, JPB^{+}08, KS90, LC02, MP92, Pen05, PE88, ZT17]. C# [Fru10]. C-Flavours [KS90]. cached [Buf17]. Cactus} [RGP98]. Calculus [GS86, Abd75a, Abd75b, AMF13, BL92, DLP07, AKPG02]. Calendar [WPR06]. Call [Ano07a, Ano07b, Bre18, Kır02]. call-tracking [Kır02]. candidates [FT15, FT16]. capabilities [CGG^{+}09]. card [SK14]. Carla [CC95]. Case [DKK18, Zav86, Alj16, Bli94, BJ14, BMN^{+}18, MKPW06, NPS17, RT18]. CASL [WMP^{+}08]. CCS [NN09]. CDL [LS90, LS94]. cellular [VLC98]. centric [LDG09]. chaining [HGC^{+}09, VS93]. Chains [Ken78]. challenges [PBDF12]. change [Ban17, BGM^{+}18]. change-aware [BGM^{+}18]. changing [Pun01]. channel [Fis88]. Characterization [DK83]. Checking [Bai86, DL17, CCT08, DQ09, GSP17, Ier93, JKK^{+}16, JL96, KKP^{+}15, MS93, MVT^{+}16, MP17, MP92, Pen05, Pen14, PRR12, Sis04, ZP04]. Chinese [TC81]. Choosing [MT82]. CHR [BKL18]. CHR-Based [BKL18]. CIaaS [ONB^{+}18]. circular [SH15]. CL ARRAY [ZT17]. Class [Log09, BDNW05, JD94, Wal89]. Classboxes [BDNW05]. classes [Ban17, VBDPM16]. classical [Har97]. classification [BKSW09, WD04]. clause [KG17]. Client [Led99a, CJD17]. client-side [CJD17]. Client/Server [Led99a]. clone [FT15, FT16]. cloning [CHK93]. closer [FF86]. Closure [FL92]. Closures [FL87]. cloud [ZLZ^{+}18]. clustering [AC18]. Co [MKPW06, LCC07]. co-allocation [LCC07]. Co-evolving [MKPW06]. COBOL [Tha77, Pet78]. Cocke [Man78]. Cocke-Younger-Kasami [Man78]. Code [Ano88, BT86, CJ80, DK83, DH86, FL87, JRSB85, RS82, BDB90, BBRR12, BC13, BM95, CAS08, CBTR17, CCJ93, Dha88, Gan89a, GDD12, Hat91, HV93, IS18, Kha10, Kha11, MT05, MKPW06, PM18, SNP16, SLS18]. Coding [PM18]. COLD [CKM^{+}18]. collaboration [MGLFCP12]. colony [AC18, PM18]. Coloring [CAC^{+}81]. COM [Led99c, Led99b]. Combination [FW78]. Combinator [JRSB85]. Combinators [MO83]. Combining [SA16, BM95]. Commands [Bai86]. comments [AA89]. common [RW09]. Communicating [DH86, DMVD16]. Communication [Bro88, AKPG02]. communications [CC95]. compact [HS03]. comparing [EvdSV^{+}15]. Comparison [Fle84, SIK09, Tha77]. Compilation [Sch78, BRB07, VMD18]. compile [BGM^{+}18, FL92]. Compiler [Ano07a, Ano07b, MB85, HSS88, Hat91, JPB^{+}08, KMLS15, MB75]. Compiler-Architecture [Ano07a]. Compilers [Sha80, ZT17]. Compiling [PMS15, PMS16, WF78]. complete [GL95]. completely [RH18]. Complex [Spr79]. complexity [BZ88, IPF82, Ste84]. compliant [MZC10]. Component [WBGM10, CC15, CC16, FDH08, FBDH12, PSW95]. Component-based [WBGM10, FBDH12, PSW95]. component-level [CC15, CC16]. component-oriented [FDH08]. components [CV16, PSW^{+}13, Tay96, Zdu06]. composable [LMR93]. composed [MW82]. composing [HRW18, RDB15]. composition [BBT15, BBT16, Bou04, BRT99, DSW05, PPK11, PCB^{+}18, RPB09, Zdu06]. Compositional [GSX99]. comprehensible [FT15, FT16]. Computation [CIF84, Nag79, AJ93, CAS08, MST14, PT09]. computational [HT13, LCC07, jLtCxH09, ONB^{+}18]. Computationally [RS87]. computations [DLP07, PRD02]. Computer [BS78, CF02, HR91, Rin91, Jos78, Nym95, Zak88]. computer-based [Zak88]. computers [BZ88, PS94b]. Computing [Ano07b, Bry15, Bry16, MB13, MB14]. ConC [GR91]. concept [MT05]. Concepts [DCA^{+}15, DCA^{+}16, GAS17]. conceptual [GWDD06, Rod15]. Concern [CKM^{+}18]. Concern-oriented [CKM^{+}18]. concerns [SNP16]. Concur [SBF80]. Concurrency [Geh82, KPP93, FO02, KH12]. Concurrent [MMC15, MMC16, SBF80, Sal83, CS03, CGG^{+}09, CO98, Dre96, GR91, GMMP89, LfL00, MP17, MW96, Rom97, Tal93a, Tal93b]. concurrent-write [CS03]. condition [SSM10]. conditions [PNF^{+}18, SSM10]. Conference [DL17, GAS17]. Conferences [CPPV15]. configurable [BGM^{+}18]. configuration [PMK^{+}18, Zdu06]. connected [PS94b]. connectives [Kor15, Kor16]. connectors [PPK11]. Considered [Sym85]. consistency [KKP^{+}15]. Constant [Tai79]. constrained [KJTA17]. Constraint [YG93, ZCM^{+}17, HHLv89, LfL00, Zim86]. Constraint-driven [YG93]. constraints [Luq93]. Construct [ECB12]. Constructors [MW82]. Constructs [BGMT82, Abd75a, MP00]. consuming [BER00]. consumption [Ozt11]. container [McC91]. containers [ZT17]. content [LR17]. Contents [Ano02d, Ano05f, Ano05g, Ano06c]. Context [BS92, Cel81, HWM13, BC93, BDL^{+}12, CDF18, IvdS17, Seb89, SK18]. context-aware [BDL^{+}12]. Context-Free [Cel81, BS92, BC93, Seb89, SK18]. Context-sensitive [HWM13]. contextual [SP18]. contiguous [KR95, LR17]. continuation [Wan92]. continuation-based [Wan92]. Continuations [HFW86, WF78, DH89, JD94]. Continuous [SBF80]. Contract [KPN17]. Contract- [KPN17]. contraction [CKS83]. contractual [OK18]. Control [CG84, LS84, AG17, AL85, AMF13, CKS83, DNR90, HB16, MC96, OM92, OM91, PSW^{+}13, Ric16, SC94, YF98]. Controller [TC81]. controlling [BDNW05, NH93]. conversation [CG93]. Conversations [Rom95]. convex [KG17]. coordinating [CLMT01]. coordination [CG96, CFG00, PPK11, SRRB10]. copies [BC13]. Copyright [Jos78]. Core [dLZ12]. corecursion [Anc13]. corollaries [Sch75b]. Coroutines [HFW86, KS90]. Correction [FM80]. Correctness [Ber77, YD78, Liu93]. correctors [Wet77]. Cost [DMT10]. Cost-driven [DMT10]. Costing [EL07]. costs [Lou07]. count [NH93]. Countdown [Led99d]. counting [CCGC12]. coupled [SRRB10]. coupling [ECB12]. covariant [CCT08]. Cover [Ano02c, Ano03c, Ano04b, Ano05b, Ano05c, Ano05d, Ano03a, Ano03b]. Creating [BDPW08, FF89]. critical [PMS15, PMS16]. critique [Fis88]. cross [ALS^{+}18, CBTR17]. cross-platform [CBTR17]. CSP [PB84]. CSP-S [PB84]. custom [FO10]. Customizing [Mal10, NPS17]. cycle [Hoo89]. Cyclic [CCGC12].

DAGs [KR95, Kes98]. Data

[Bai87, BF78, BC84, CS03, Fle78, GZ87, Geh79, Han78, KK18, KJ12, MO83, PBG84, YD78, BT91, BEL77, Ber77, BMZM92, BC13, CNGW09, DOZ06, DQ09, Ear75, FC18, FW87, FF89, Geh77, HG93, HC96, Jal92, JG89, JO11, KDM03, McL77, Mic96, MP17, MP00, Nil90, OAB^{+}18, OM91, PRD02, SJW94, SP18, ZLZ^{+}18, dlVGSZS18]. Data-Bases [BC84]. Data-bound [KJ12]. Data-Flow [MO83, MP00]. data-parallel [Mic96]. data-parallelism [HC96]. Data-race [CS03]. Database [Orm83, PC85, DCA^{+}15, DCA^{+}16, HC12]. databases [BL92, HHLv89]. Dataflow [Wei85, Ozt11]. Datatype [Wei85]. Debugger [CDGN15]. debuggers [CDGN15]. Debugging [Joh81, COHW95, FC18]. decentralized [HB16]. declarations [SC94]. Declarative [SH15, ZTLM13, CL97, CFG00, Mic96, NL95, SNA18]. decorator [Alj16]. Deducing [Sch75a]. Deep [Sam79, Kha11, SA16, ZH18]. DeepDSL [ZH18]. define [BG84]. defined [DNR90]. defining [yCH92, RDB15]. definite [GG09]. Definition [BF78, BSW15, yCH92, CG84, Ken78, CDF18, CRPP00, KB75, McL77, RTMRK18, Thi82]. definitional [Fal97]. delayed [VS95]. delayed-load [VS95]. Delving [MT05]. democratisation [dlVGSZS18]. denotational [Ier93, Mal93]. denotations [HS03]. dense [DLP07]. dependence [BC13, SSM10]. dependencies [PS10]. dependency [ZP18]. Dependent [JO11]. deployment [MLW05]. Derivation [PS86, SvE16, RR99]. Deriving [MB85, HRW18]. DesCaRTeS [MRO03]. descent [Hor93, MPS90]. Description [KP78, PB84, Rid79a, Bay76, Hor17, SMdSB09]. Design [AAH95, ABG^{+}05, ESG16, FFMB11, KN85, Mic96, RS83, Sch78, TC81, VLC98, ZA87, Zak88, ZH18, Alj16, Bas75, CS16, CDW09, CSdL16, COHW95, DCA^{+}15, DCA^{+}16, FBDH12, FM04, FWY96, KS90, LDN18, LP97, LS94, MST14, MSRG10, MKPW06, Run89, Sco91, SSS17, Tuc75]. Designing [HG93, Ear75]. destructive [HV94]. detailed [KHO14]. Detection [Pai16, FM04]. determinism [OM92]. deterministic [Lee05, PTJM16, PRD02, RP98]. determinization [PTJM16]. developers [RT18]. Developing [BB91]. Development [CDGM80, GG82, HR91, Bai90, BAG18, BDPW08, BIMP17, CBTR17, yCH92, CKM^{+}18, ESG16, Mal10, MAGD^{+}16, MŽ05, PFH16, RK18, Rot92, SK14, SSS17, VC15, WHKK17, WD04, dlVGSZS18]. Developments [Cro79, Fle78]. Devices [Sym85]. DFL [PBG84]. diagram [ABS17]. diagrams [DKK18, Her76]. dialects [CHHP91]. dialogue [Nym95]. different [Coo98, Sli17]. digitaled [HLJ76]. Dijkstra [Bai86]. DILOG [HLJ76]. DILOG-digitaled [HLJ76]. dimensional [FT15, FT16]. Direct [MB75]. Directed [LBR81, DS93, Har97, Kha10, Nil90, OWG93, VS93]. Discrete [BB91, Bli94, Hoo87]. dispatch [KA07]. Dispel [Joh81]. Display [MOT84, NK90]. distance [Dai94]. Distributed [BT91, BGMT82, CLSM96, Coo81, Kır02, Led99e, PB84, Tal93a, YF83, CNGW09, DRT97, LS94, MY17, NJLS12, PLS10, PJ91, Sco91, SRRB10, Tay96, Whi77, ZTLM13]. do [RT18]. documentation [LBGA18]. documents [CNGW09]. Domain [MMC16, SNA18, CJD17, CDGN15, CSdL16, CCF15, FFMB11, GAGdL17, LDN18, PSW^{+}13, RTMRK18, SvdBV18, SA16, dlVGSZS18, MMC15]. Domain-specific [MMC16, SNA18, CDGN15, CSdL16, FFMB11, GAGdL17, PSW^{+}13, RTMRK18, SvdBV18, SA16, dlVGSZS18, MMC15]. Domains [DMVD16, McL77]. dominance [AC18]. Dragonfly [AMR18]. Driven [ACG18, BF78, DdLP17, AC18, ABL17, BAG18, BIMP17, DMT10, KGS17, LDN18, jLtCxH09, Rod15, SK14, SRT17, WHKK17, YG93, ZCM^{+}17]. DRL [DRT97]. DSL [BAG18, PFH16, PLDD15, SvdBV18, ZH18]. DSLs [CV16, LBGA18, MAGD^{+}16, NPS17]. DSML4CP [MMC15, MMC16]. DSU [CCRS18]. Dynamic [BB91, BRT99, CCRS18, GG09, BKSW09, BG84, BS18, FF90, GBZ09, HDN09, LC02, LDG09, Pen05, PRD02, PLS10, RH18, RN09, RT18]. dynamically [Ber11, Pun01].

Early [MOT84, CS16]. easytime

[FFMB11]. edge [Dha90]. editing [Thi82]. Edition [Led99a]. Editor [Ano01a, DP09]. Editorial [Ano01a, DW04, LP16, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano05b, Ano05c, Ano05d, Ano09a, Ano10a, Ano18a, Ano18b, Ano18c]. Edwards [Led99a]. EE [MCC17]. Effect [GFK81, IR95]. Effective [DMVY17, Flo78, HMHS18, CSdL16]. effectiveness [DTXP13]. efficiency [PGT^{+}96]. Efficient [BDB90, JRSB85, JPB^{+}08, PDK^{+}09, PTJM16, CCJ93, FF89, Hat91, Li96, Lia92, Pai16, PT09]. Effort [CIF84]. Elements [Pet78, Whi77]. eliminating [RW09]. Elimination [BC13, Dem75]. Embedded [Ano07a, ABG^{+}05, HL08, JPB^{+}08, MRO03, NPS17, PDK^{+}09, Wan92]. embedding [KMLS15, SA16]. Emerald [HHS90]. emphasis [ALS^{+}18]. Empirical [Ban17, SW77, SJW94, VBDPM16]. Employing [Sis04]. enabled [PPK11]. end [LBGA18]. enforcement [IF16]. engineered [Hug85]. Engineering [ACG18, CPPV15, DdLP17, MAGD^{+}16, SSJB96, BMN^{+}18, CKM^{+}18, HRW18, KGS17, Mal17, Man01, Rod15, ZCM^{+}17]. Engines [DH89, HF87]. Enhancement [DOZ06]. Enhancements [ZL81]. Entity [SS79, DCA^{+}15, DCA^{+}16]. entity-relationship [DCA^{+}15, DCA^{+}16]. entry [MC96, OM92]. Environment [MOT84, RS83, DGU91, JD94, KA07, PJ91, PSW95]. Environments} [Led99e, PRD02]. ENVISAGER [DGU91]. epsilon [FL92]. EQL [Nag79]. equational [Hat91]. equivalence [Tze12]. Error [CB93a, FM80, Dai94, HRS84, LCFÁ10, Wet77]. errors [DP98, RD78]. escape [DLP15]. Evaluating [EvdSV^{+}15, KR98]. Evaluation [CD81, GFK81, ABG^{+}05, DCA^{+}15, DCA^{+}16, DPP10, FW87, Jay92, KHO14, LRB^{+}11, MC96, MS89, NS93, PBDF12, PS94a, SIK09, Tre00, TM00]. evaluations [KR95]. Event [BB91, DMVD16, Ric16, SRRB10, VMD09, BMN^{+}18]. Event-B-based [BMN^{+}18]. event-based [SRRB10]. event-loop [DMVD16]. evolution [OAB^{+}18]. Evolutionary [KR17]. evolving [MKPW06]. Exception [DG94, LS90, BKYV80, CM11, CC15, CC16, CD82, HO90, JPB^{+}08, Rom97]. exceptions [BJ90]. exchanging [FF89]. executable [CIP^{+}00, HZ96, KJ12, PCG16]. execute [FKR75]. Execution [LS84, ALR15, BSW15, BJ14, CPD93, GMMP89, LfL00, MB75, PLS10, ZLZ^{+}18]. Executions [BKL18]. exercise [Sal92]. exercises [GAGdL17]. Existing [EvdSV^{+}15, AA09]. expecting [DG94]. Experience [Wei85, Sco91]. Experiences [GAS17, MOT84]. experimental [Ste75]. Experiments [HV94]. expert [SNA18]. explicit [OM92]. exploiting [KKG15]. Exploration [SvdBV18, GHD18]. explorative [KR17]. export [FF86]. exposed [VBDPM16]. Expression [Tai79, KR95, Kes98, LRB^{+}11, PS94a, SvdBV18, VS95]. Expressions [WF78, GGK^{+}11, Pai16]. Extended [Cel81, RS87, SS79, DCA^{+}15, DCA^{+}16]. Extended-Entity-Relationship [SS79]. Extending [BL99, LCFÁ10, MZC07, Sul75]. extensible [KMLS15, RH18, RDB15]. Extension [Nag80, BRS90, Sal92, ZT17]. extensions [BDNW05, Dre96, FO10]. external [MAGD^{+}16]. Extracting [PSW95].

Facet [BC93]. Facilitating [CBTR17]. Facility [BF78, Nag80, BJS93, FF90]. Factorizations [WF78]. Fail [Dre96, Wet77]. Fail-safety [Dre96]. Fairness [OBGK02]. Familial [Orm83]. fast [HZ96, MLW05]. fault [CL89, DTXP13, ZP18]. fault-localization [DTXP13]. fault-tolerant [CL89]. Faust [JO11]. Feature [MM82, MLW05, VC15]. feature-oriented [VC15]. feature-rich [MLW05]. Features [Hoo87, CM75, Hoo89, RT18]. Feedback [Kha10]. Feedback-directed [Kha10]. Ferret [BKSW09]. fields [CCGC12]. file [BGM^{+}18]. Financial [Hor17]. finest [Yan00]. First [JD94, KH12, Wal89]. First-class [JD94, Wal89]. First-order [KH12]. fit [HC05]. FLANDM [dlVGSZS18]. flavour [KS90]. Flavours [KS90]. FLEX [LN91]. Flexary [Kor15, Kor16]. Flexibility [SRRB10, URI02]. Flexible [DdLP17, LN91, BBRR12, DCB^{+}17, PT09, VMD18, ZCM^{+}17]. floating [Won99]. floorplanning [Bou08]. floorplanning-placement [Bou08]. Flow [MO83, PBG84, RS82, Sha80, CHH02, IF18, KDM03, LDG09, MP00, NN17]. Flowcharts [CL78, Pet78, Coh78]. fly [RDB15]. FOBS [dLZ12]. Folding [Tai79]. fork [MZC10, RS94]. fork/join [MZC10]. Formal [AMA98, MY17, Pag78, Sha81, ZA87, ALS^{+}18, Bou08, DLP15, KB75, MT05, OK18, Wil80, Nag80]. formalization [BDNW08]. formalizing [vOKF01]. Formatter [ZA87, AA89]. forms [FKR75, SIK09]. formulation [PNF^{+}18]. FORTH [AL85]. FORTRAN [Tha77]. Forward [HGC^{+}09, Man01]. Fostering [CKM^{+}18]. Foundations [FDH08]. FP [AC18, BC84, Fle86]. FP-ABC [AC18]. Fp-Style [Fle86]. fragment [WHKK17]. Frame [PNF^{+}18]. framework [ACZ05, AA09, BBRR12, BM95, CS16, CY02, CDGN15, HT13, JM96, KMLS15, NN09, Rot92, SSB94, VC15, VMD18, ZLZ^{+}18, ZT17, dlVGSZS18]. framework-based [ACZ05]. Free [Cel81, BC93, BS92, IR95, Seb89, SK18]. freedom [CS03]. Front [Ano02c, Ano03a, Ano03b, Ano03c, Ano04b, Ano05b, Ano05c, Ano05d]. Fully [PC15]. fun [Mor16]. Functional [AD07, Bai87, BJ90, Fle86, FW78, HM18, LP16, MO83, Bai90, FC18, HV94, Lia92, Mal10, Mal93, Thi93, dLZ12]. functionalities [PLDD15]. functionality [MRO03]. Functions [Mic86, IR95, SS09, UM17, CCF15]. Fundamental [Sym85]. future [EvdSV^{+}15]. Fuzzy [AC18, Dja88]. Fuzzy-Pareto [AC18].

G [Bai87, Pla91]. Galois [NN17]. GALS [MSRG10]. Gap [YD78, FBDH12]. general [BM95]. Generalized [Car78, LS84, PC85]. Generated [Pet78]. Generating [KR95, BC89, Noo85]. Generation [BKL18, FL87, Wad80, BDB90, BM95, CAS08, CNGW09, DPP10, FL92, Gan89a, GAGdL17, Guo16, Hat91, Hor90, SLS18]. Generative [GAS17, Mor16, SSS17]. generators [Bud82, Gan89b]. Generic [Bai87, CGG^{+}09, Geh80, Bai90, Bou08, FC18, ZT17]. generics [EHMO91, TKH99]. genetic [FNRR16]. genuinely [BJS93]. genuinely-lazy [BJS93]. geometry [RH18]. Global [BT86, Zob93]. goal [Har97, Nil90, OWG93]. goal-directed [Har97, Nil90, OWG93]. goals [Lee05]. GPCE [GAS17]. GPGPU [KPN17]. GPU [PM18]. Grammar [BSW15, BEH86, RP98, Sar94, CY02, RDB15, SK18]. Grammar-based [BSW15]. Grammars [BF78, Cel81, CF79, Mic86, Pag79, BC89, Dem75, HSS88, JGM98, NS93, Seb89, Yan00, BC93]. grammatical [Nym95]. Graph [BF78, BRS90, BJ14, SSM10]. graphical [RK18]. graphics [Zak88]. Graphs [LBR81, MO83, RS82, BC13, VS93]. GRAS [BM95]. Green [dOG06, dOG09]. Gregorian [WPR06]. Grid [Geh79, LCC07, jLtCxH09]. grids [GH07]. Guarded [Bai86]. Guest [Ano01a, DP09].

HALO [HGC^{+}09]. handle [BL99, PRD02]. Handling [Bai86, GG82, BKYV80, CM11, CO98, CB93a, CD82, Dai94, DP98, DG94, HO90, JM96, JPB^{+}08, LW75, LS90, Rom97]. hardware [DPP10]. Harmony [AC17]. Hebrew [NB84, NB84]. height [PTJM16]. height-deterministic [PTJM16]. hemodialysis [BMN^{+}18]. Heterogeneous [PC85, ÁdLNW18]. Heuristic [VS93]. heuristics [MP17]. Hierarchical [LBR81, BZ88, Bai90, Bou08, PSW^{+}13]. hierarchy [GSP17]. High [CIF84, Ear75, Geh79, McL77, BEL77, Ber77, CCJ93, Lou07, MB75, Sch75b, Sch75a, Tuc75]. high-level [Lou07, MB75, Tuc75]. higher [Fal97, KH12, RW09, SvE16]. higher-order [KH12, RW09, SvE16]. historical [BL92, Fel87]. history [Fri92, HGC^{+}09]. history-based [HGC^{+}09]. Hoisting [CJ80]. Horn [KG17]. host [NPS17]. Huhu [NB84]. human [Nym95]. human-computer [Nym95]. hybrid [BMN^{+}18, dLZ12]. hybridizations [Mal17].

IBM [FF75]. ICCL [CB93b]. Icon [Gri83, OWG93, Wal89]. IDE [NPS17]. identify [Ban17]. IFC [Ano04a]. II [Abd75b, Ber77, Sch75a]. image [WDCL08]. imageSegment [PBDF12]. imperfect [ZP18]. Implementable [BEH86]. Implementation [CMM85, GZ87, Geh80, HMHS18, MT82, PB84, RS83, TC81, ZL81, AA89, ABG^{+}05, BAK89, Bud82, CL97, CDF18, FBDH12, FWY96, FFMB11, FW87, GWDD06, HGC^{+}09, KJTA17, Lia92, MC96, Mic96, OWG93, RM93, VLC98, ZH18]. Implementations [Sal83, CKS83]. Implementing [Alj16, BF78, Gri83, KNW94, VSN^{+}17]. implicit [IvdS17]. import [FF86]. imprecise [BL99]. Improved [Man78, CCT08]. Improving [ALS^{+}18, Kha11, PGT^{+}96, Ten83, DTXP13]. inclusion [Sch75a]. Incremental [Hor90, MŽ05, MS89, MPS90, Li96, SB04, VS94]. incrementally [NJLS12]. Independent [BT86, FM04, IF16, PGT^{+}96, VF82]. Index [Ano00, Ano01b, Ano05a, Ano05e, Ano05g, Ano99, Ano06c]. Induced [TBKG04]. induction [PC78]. induction-inference [PC78]. inductive [AG17]. Inference [CF79, UM17, PC78, Pun01, SvE16, ZCM^{+}17]. inferencing [KDM03]. Information [CHH02, ÁdLNW18, Ano09a, Ano10a, DCA^{+}15, DCA^{+}16, Hor17, IF18, KKG15, LDG09, PR10, ZTLM13]. infrastructure [GDD12]. Inheritance [SS92, Bou04, MW96, TKH99]. inlining [HWM13, KR98]. input [BER00]. input-consuming [BER00]. insertions [NN17]. inspired [ABS17]. instructions [Dha90]. integrated [KA17, LCFÁ10]. Integrating [HHLv89, HHS90, PT09, ÁdLNW18]. Integration [Sha81, ACZ05, LP97, MY17, SP18, Tal93b, WD04]. integrity [NN17]. Intel [HMHS18]. intelligence [HLJ76, ONB^{+}18]. intelligent [ÜA15, ÜA16]. intensional [MKPW06]. intentional [TBKG04]. Inter [GWDD06, MC96, OM92, FO02]. Inter-entry [MC96, OM92]. Inter-language [GWDD06]. inter-program [FO02]. Interacting [YF83]. Interaction [Ano07a]. Interactive [GG82, LfL00]. interdiagram [KKP^{+}15]. Interface [MP92, CNGW09, Tay96, Thi82, Zak88]. interfaces [Pun01]. Intermediate [BT86, McC91, BG84, FC18, MB75]. International [CPPV15, DL17, GAS17]. interpretation [CZ11, HC12, Log09, RK93, UM17, DL17]. Interpreter [GS86, BBT15, BBT16, PT09, VSN^{+}17, Zim86]. Interpreters [Mic86, IvdS17, RR99]. Interprocedural [CD81]. intervals [BL99]. Intra [KKP^{+}15]. Intra- [KKP^{+}15]. Introduction [Ano01a, BW05, CB93b, DP09, HR91, HR92, Lou07, Rin91, SD06]. intrusively [MZC10]. invariance [GSP17]. invariants [AG17, Log09]. inversion [SM89]. investigation [PLS10]. invocation [CO98, OBGK02]. invocations [GH07]. involving [LRB^{+}11, ZP18]. ISBN [Led99b, Led99c, Led99d, Led99e, Led99a]. Isolating [FO10]. Issue [CPPV15, GAS17, LP16, ACG18, Bry15, Bry16, CB93b, DL17, DdLP17, HM18, HB16, KGS17, Lou07, Mal17, MB13, MB14, SD06, Zuc04]. issues [CL89, COHW95]. Iteration [MP00]. iterators [Ear75].

J [Fel87, KMLS15]. J-operator [Fel87]. JADE [BIMP17]. JADEL [BIMP17]. Jager [Led99d]. Java [ACZ05, BCR11, CV14, CY02, CSdL16, HWM13, IS18, IF16, JPB^{+}08, KMLS15, MCC17, PT09, Rez12, TKH99, VBDPM16]. JavaBean [MZC10]. JavaLog [ACZ05]. Jeri [Led99a]. jLab [PT09]. John [Led99b, Led99c, Led99d, Led99e, Led99a]. join [MZC10]. JR [CGG^{+}09]. Jumps [Abd75b]. Just [dACSAP14, VMD18]. Just-in-time [dACSAP14, VMD18].

Kasami [Man78]. kernels [KKG15]. Keyword [Ano05e, Ano05g, Ano06c]. know [Sch76]. knowledge [CBTR18, SNA18].

LAILA [CLMT01]. Lambda [GS86, WF78, Abd75a, Abd75b, FL92]. Lambda-Calculus [GS86, Abd75a, Abd75b]. Lambda-Expressions [WF78]. Landin [Fel87]. Language [Ano07b, BS78, Bai87, BT86, Bar82, BEL77, BGMT82, BC84, CV16, CCRS18, CPPV15, DGU91, FM04, GS86, GO88, Hoo87, Hoo89, Hul87, Joh81, KN85, KP78, MT82, MMC15, MMC16, MO83, MM82, Nag79, Nag80, Orm83, PBG84, PC85, RBY^{+}05, Rin91, SBF80, ZL81, AL85, AAH95, ALR15, Bas75, BL92, Bay76, BIMP17, BKSW09, BAK89, Bou08, BG84, CIP^{+}00, CGG^{+}09, yCH92, CLMT01, CFG00, CC95, CL89, CKM^{+}18, CSdL16, CHHP91, DRT97, Dja88, EL87, EvdSV^{+}15, FDH08, FBDH12, FFMB11, GR91, GAGdL17, dOG06, dOG09, GWDD06, HDN09, HV94, HHS90, HZ96, HRW18, Hor17, Hug85, JD94, KKG92, KA17, KNW94, LMR93, LDN18, LP97, LB89, Liu93, LS94, Luq93, MSRG10, Mal10, Mal93, MR01, MŽ05, MB75, Mic96, ND77, NL95, OWG93, ØK00]. language [PGM84, Pla91, PE88, PSW^{+}13, RN09, RTMRK18, RGP98, Run89, RH94, Sal92, Sco91, SS92, SMdSB09, Ste75, Tuc75, Tze12, VC15, VLC98, WMP^{+}08, Wan92, WDCL08, Zdu06, Zim86, dLZ12, Bai86, yCH92, RS94]. Language-And [BT86]. language-based [Bou08]. Language-independent [FM04]. language-integrated [KA17]. Languages [CIF84, CG84, Cro79, HR91, HR92, MB13, MB14, Was79, vOKF01, Abd75a, Abd75b, Bai90, BC88, BLM93, BL99, Ber11, BEL77, Ber77, Bry15, Bry16, BW90, CL97, CJD17, CO98, Cia92, CHH02, CG93, CF02, CDF18, COHW95, CRPP00, Fri92, HC12, HHLv89, HG93, HRW18, Ier93, IR95, KPN17, LCFÁ10, LfL00, Lia92, MP00, NK90, OM91, PC78, PCB^{+}18, RDB15, Rot92, Rus87, Sch75b, Sch75a, SNA18, SvdBV18, SA16, YG93, dlVGSZS18]. languages-value [Sch75b]. LARA [PCB^{+}18]. large [HRW18, LRB^{+}11, MP92, SJW94]. Lass [Bar82]. layered [MR04, OM91]. layers [Vai04]. Lazy [Han97, BJS93, FC18, HV94, Jay92, Tre00]. learning [ZH18]. Lenient [Tre00, TM00]. Level [BEH86, CIF84, Geh79, MO83, Pag79, BEL77, Ber77, Buf17, CCJ93, CC15, CC16, Ear75, Lou07, MSRG10, McL77, MB75, Sch75b, Sch75a, Tuc75]. Leveraging [FAHC17, MAGD^{+}16]. lexical [Yan96, YTC02]. lexically [FF90]. libraries [FF75]. library [CDW09, KPN17, ZT17]. library-based [KPN17]. life [Hoo89]. Lightweight [MVT^{+}16, PCG16, CC15, CC16]. like [CHH02, MRO03, Was79, NK90]. LINDA [SS93, RPB09]. line [KR17, PMK^{+}18, SSS17]. linear [Dha90]. linearly [PS10]. Lines [MAGD^{+}16, FAHC17]. Linguistic [PRR12, VMD09]. Lisp [FWY96]. List [Ano09b, Ano10b, Ano11a, AJ93, FKR75, War78]. listless [Jay92]. Lists [Wad80, Lus02]. literature [MAGD^{+}16]. LL [BC89, Li96, Sli17]. LLLR [Sli17]. load [Dha90, VS95]. local [CDF18, FNRR16]. localization [DTXP13]. Logic [ACS96, HS03, RS87, BER00, BKG^{+}08, BRT99, CLSM96, CLMT01, Cia92, CG96, DRT97, FL01, GG09, Har97, HC96, HGC^{+}09, HLJ76, JM96, JGM98, KNW94, LMR93, LP97, LfL00, NJLS12, RGP98, Tal93b, KPP93]. Logical [CIF84, TSF^{+}87, IPF82, JG89]. longest [YTC02]. longest-match [YTC02]. look [FF86]. Lookahead [SC87, Ber91a]. loop [KKG15, DMVD16, SF89, VMD09]. Loops [DK83, Bli94, Ric16]. loosely [SRRB10]. Low [MO83]. Low-Level [MO83]. LR [BC88, Ber91a, Cel81, CB93a, DP98, Dem75, Hor90, Sli17, VS94, WBGM10]. Lynx [Sco91]. LySa [BC10].

Machine [BT86, Pet78, BMN^{+}18, HLJ76]. Machine-Generated [Pet78]. Machine-Independent [BT86]. machines [ALS^{+}18, RM93, Yan96]. Macro [Nag80]. Macro-Oriented [Nag80]. macros [FKR75]. Macrospace [Flo78]. made [Mor16]. Maintaining [NJLS12]. Making [Flo78]. malleable [MZC10]. management [DOZ06, LC02, OOB18]. Manipulating [GO88]. manipulation [Mal93]. many [AC18]. many-objective [AC18]. Mapping [FNRR16, SLS18]. Martin [Led99c]. match [YTC02]. Matching [Gri83, Liu88, BDB90, CF88, Nil90]. Mathematical [KP78, KKG92]. MATLAB [AMR18]. MATLAB/Octave [AMR18]. Matrix [HMHS18]. maximisation [FT15, FT16]. MDL [yCH92]. ME [BAG18]. Mealy [Yan96]. measure [Geh77, IPF82, Ste84]. measurement [ECB12]. Mechanical [Ric80]. Mechanism [YF83, Bai90, CC15, CC16, CD82, HO90, RDB15]. Mechanisms [CCT08, CO98, MC96, OM92].

Mechanizing [McK75]. medium [SJW94]. membership [Sch75a]. Memory [BKL18, LRB^{+}11, DMVY17, HC05, KKNS14, KKG15, LC02, Ozt11, PLS10, RN09, ÜA15, ÜA16]. Memory-optimal [LRB^{+}11]. merge [ACS96]. mergesort [SJW94]. merging [Dem75]. mesh [PS94b]. meshes [PS94a]. Message [CLM83, MR04]. meta [GDD12]. meta-model [GDD12]. metaclass [Bou04, DSW05]. metaclasses [RBY^{+}05]. metadata [CNGW09, RH18]. Metalanguage [BEH86, BKG^{+}08]. MetaMod [SvdBV18]. metamodel [FAHC17]. metamodeling [CCB15]. metaprogramming [GBZ09]. method [Ber11, Dai94, Ear75]. methodologies [FW87]. Methodology [BB91, CLM83, Bou08, CHK93, KKG15, LP97, yCH92]. Methods [Wil81, KK18, WHKK17]. metrics [OOB18]. Micro [CCRS18, WMP^{+}08]. micro-architectures [WMP^{+}08]. Micro-Language [CCRS18]. Microcomputers [ZA87, ZGE85]. microprogrammed [MB75]. Microsoft [Led99b]. middleweight [IF16]. MIMD [VLC98]. minimal [Kes98]. minimalist [LB06]. minimizing [PS10]. minimum [Dai94]. mining [MP17, dlVGSZS18]. misconfigurations [MCC17]. missions [OAB^{+}18]. mixed [HHS90]. mixed-language [HHS90]. mixin [Bou04]. mixin-based [Bou04]. mixins [BS18]. Mizar [Kor15, Kor16]. MMFilter [BKL18]. Mobile [Led99e, BCF02, BCF^{+}04, HL08, RK18, VBS^{+}14]. mobility [DMT10]. moded [BER00]. Model [ACG18, ABL17, DL17, DdLP17, MCC17, Pen14, Rod15, SS79, ZP04, Abd75a, Abd75b, BDL^{+}12, BIMP17, Ber91a, BSW15, BMZM92, BS18, CM11, CAS08, CV14, DCB^{+}17, DQ09, ESG16, GHD18, GAGdL17, GDD12, GWDD06, JKK^{+}16, KGS17, DMVD16, MP17, OAB^{+}18, OM91, Rus87, SK14, SP18, Sis04, SRT17, WPR06, WD04, Yan96, ZLZ^{+}18, ZCM^{+}17, FAHC17]. Model-based [MCC17, FAHC17]. model-checking [DQ09]. Model-Driven [ACG18, Rod15, BIMP17, KGS17, SRT17, ZCM^{+}17]. Modeling [MZGT85, MMC15, MMC16, Spr79, CBTR18, DCB^{+}17, DCA^{+}15, DCA^{+}16, FAHC17, PLDD15, RK18, ZP18]. modelling [ÁdLNW18, GVvdP^{+}01, Rid79b]. Models [BS78, BKL18, Fle84, GZ87, KP78, BKG^{+}08, DMVY17, GSX99, JKK^{+}16, KKP^{+}15, Lou07, OAB^{+}18, PCG16, PNF^{+}18, RBY^{+}05]. modern [WMP^{+}08]. modification [RDB15]. modifies [PNF^{+}18]. Modula [DNR90, Sal92]. Modula-2 [DNR90, Sal92]. Modular [EHMO91, IvdS17, KN85, BKYV80, CV16, RP98]. modularity [SvdBV18]. module [AC18, PC15]. modules [BRT99, LMR93]. MOF [CCB15]. Moldable [CDGN15]. Monaco [PSW^{+}13]. monitoring [IF18]. monotone [NN09]. morality [McK75]. Motion [CJ80, BC13]. motivation [Sco91]. movilog [MZC07]. MSC [Man01]. mud [Vai04]. MudPie [Vai04]. Multi [FFJ90, MOT84, ABL17, AMR18, DCA^{+}15, DCA^{+}16, KGS17, WHKK17]. multi-agent [ABL17, KGS17, WHKK17]. Multi-Display [MOT84]. multi-paradigm [DCA^{+}15, DCA^{+}16]. multi-platform [AMR18]. Multi-way [FFJ90]. Multicomputers [Geh82, SSB94]. multicore [FNRR16, HMHS18]. multidimensional [DLP07, ZT17]. multiparadigm [LP97, NL95, Pla91]. Multiple [ACS96, KA07, SS79, BKSW09, PCB^{+}18, PS94b, TKH99]. Multiplication [HMHS18]. Multiprocessing [CL78, Coh78]. Multiprocessor [Hul87, Tay96]. multirate [JO11]. multiview [KKP^{+}15]. multiway [CO89]. mutant [SMdSB09]. Mutation [BS85, GAGdL17]. mutual [FF90].

N [Led99c, CBTR18]. n-ary [CBTR18]. N-Tier [Led99c]. narrowing [CZ11, Han97]. natural [Run89]. need [Kes98]. neither [Tre00]. nested [DLP15, JG89]. Nesting [BCF^{+}04]. net [GSX99]. Nets [MZGT85, GMMP89]. Network [CF79, MRO03]. Neumann [RM93]. Neverlang [VC15]. Node [Wad80, War78]. Non [BC10, LR17, MZC10, NPS17, OM92, PRD02]. non-contiguous [LR17]. non-determinism [OM92]. non-deterministic [PRD02]. non-intrusively [MZC10]. non-programming [NPS17]. Non-repudiation [BC10]. Nondeterministic [Sal83, PGT^{+}96]. Noninterference [AMF13, IF16]. nonterminal [Dem75, SH15]. nor [Tre00]. notation [SRT17, Wil80]. note [Ano06b, Ano11b, Fel87]. notion [BW90]. NP [CIP^{+}00]. NP-SPEC [CIP^{+}00]. number [DK92]. numbers [Run89]. Numerical [Nag79].

Object [ACS96, BB91, GVvdP^{+}01, GG82, LP97, PBDF12, AC17, BS18, DGU91, FM04, Ier93, IR95, KPN17, KS90, LCFÁ10, LDG09, MW96, NL95, RBY^{+}05, YG93, dLZ12]. object-centric [LDG09]. object-models [RBY^{+}05]. Object-oriented [GVvdP^{+}01, AC17, DGU91, FM04, Ier93, IR95, KPN17, KS90, LCFÁ10, MW96, NL95, YG93, dLZ12]. objective [AC18]. objects [CLSM96, KJTA17, LRB^{+}11, Lus02, LB06, MW82, Rom95, YF98, DOZ06]. observation [FC18]. observer [Alj16]. Obtaining [HFW86]. Occam [AMA97, AMA98, Fis88, Hul87, Tal93a]. OCL [CCB15, PNF^{+}18]. Octave [AMR18]. offset [CAS08]. OmniBrowser [BDPW08]. on-the-fly [RDB15]. only [PNF^{+}18]. onto [FNRR16]. Ontological [PFH16]. ontologies [SNA18]. Open [HH06, Led99e, Ban17, DK92]. OpenCL [ZT17]. Operating [Cro79]. Operation [Sam79, CG93]. operation/procedure [CG93]. Operational [MB85, LS94, OWG93, OM91]. operations [CGG^{+}09, Dja88, Won99]. operator [Fel87]. Operators [GFK81, Sym85, BLM93, CZ11]. Optimal [RS82, KR95, LRB^{+}11, jLtCxH09, PRD02]. Optimisation [KA17, Sch75a, Sha75]. Optimization [BT86, DK83, LBR81, PM18, Sch75b, KR17, NK90]. Optimizing [SS09, Sha80, Thi93, Won99, Sar94]. oracle [Guo16]. order [CAS08, Fal97, KH12, RW09, SvE16]. Orderly [AKPG02]. ordinary [MZC10]. Orientation [ACS96]. Oriented [BB91, CLM83, Nag80, ACZ05, AC17, BIMP17, CS16, CG93, CKM^{+}18, DGU91, ESG16, FDH08, FM04, GVvdP^{+}01, GCH09, Ier93, IR95, KPN17, KS90, LCFÁ10, MW96, NL95, RK18, RS94, VC15, YG93, dLZ12]. Orthogonal [CM06, Rot92]. other [Jos78]. outline [PGM84]. Over-exposed [VBDPM16]. Overloading [EL87, Ber11]. OWL [CBTR18]. Own [Zav86].

Package [Ree84]. packaging [PSW95]. Paisley [Zav86]. Papers [Ano07a, Ano07b]. ParaAJ [Alj16]. paradigm [DCA^{+}15, DCA^{+}16]. Parallel [Cia92, Cro79, Hoa75, KPP93, LCC07, LN86, PS94a, PS94b, Sch78, VS94, AD07, AJ93, AMR18, BAK89, BC13, CB93a, CM06, DLP07, EL07, HZ96, Lou07, MH07, Mic96, PLS10, URI02, VLC98, PGT^{+}96]. parallelism [HC96, MZC10, Tal93b, TM00]. parallelization [SSB94, Zob93]. parameterised [BRT99]. parameterized [ZP04]. Parameters [Pag79, DK92]. Parametric [LMR93, Alj16]. parametrization [Lia92]. Parcels [MLW05]. Pareto [AC18]. PARLOG [Tal93b]. parse [Li96]. parser [CB93a, Gan89b]. parsers [BC89, Hor90, Li96, PDK^{+}09, Sar94, Sli17]. parses [BC88]. Parsing [Cel81, GFK81, LN86, BC93, BS92, Ber91a, Hor93, MS93, MPS90, PS94b, SM94, Sha75, Sli17, VS94, WBGM10]. parsing1 [RP98]. Partial [NS93, JD94, Lia92, RDT08]. partially [BCR11]. particular [ALS^{+}18]. Partitioning [PS86, KK18, RP98, Yan00]. partitions [LR17]. partly [Fel87]. PASCAL [KY75, CM06, Fle84, Ten83, Was79]. Pascal-Like [Was79]. passing [MR04]. Path [CD81]. Pattern [Gri83, Liu88, BDB90, CF88, ESG16, Nil90]. pattern-matching [CF88]. patterns [Alj16, CS16, FM04, SSS17, Sha75, Wal89]. PC [Ano88]. PEARL [GCH09]. peer [VBS^{+}14]. peer-to-peer [VBS^{+}14]. per-file [BGM^{+}18]. Performance [CL89, FW87, Spr79, Bli94, Kha11, Sar93]. performant [VSN^{+}17]. persistent [BFPR04]. personal [AA09]. personalization [PR10]. Personalized [PMK^{+}18]. Perspective [Rin91]. Peter [Led99e, Led99d]. Petri [GSX99, GMMP89, MZGT85]. phase [ZP18]. piecewise [AG17]. PL [Sul75, Tha77]. PL-I [Tha77]. PL/1 [Sul75]. placement [Bou08, Dha88, Dha90]. platform [ABG^{+}05, AMR18, CBTR17, DPP10]. platforms [FNRR16, PLS10]. pluggable [HDN09]. plus [FL92]. point [WPR06, Won99]. pointcuts [BKG^{+}08, HGC^{+}09]. pointer [BGH13, HG93]. Pointers [BEL77, Ber77, Pen05]. points [SSM10]. policies [IF18, MY17]. polyhedral [KG17]. Polymorphic [JL96, AMF13, KNW94]. polymorphism [DCB^{+}17]. portable [RM93, ZGE85]. positive [GSP17]. PostScript [HV93]. power [CM75, RK93]. PPL [JM96, WDCL08]. Practical [CDGN15, HDN09, Rom97, Dai94, SM94]. practice [KGS17]. practices [CCB15]. pragmatic [Pai16]. PRAM [RS94]. PRAM-language [RS94]. Precise [SSM10, LW75, Pai16]. predicates [Cov93, Gan89b]. predicative [AMF13]. prediction [Lee05]. Preemption [HF87]. Preface [CK08, De 08]. Preference [JGM98]. Preliminary [MP85]. presentation [FT15, FT16]. preserving [BRB07]. primitive [CCJ93]. primitive-based [CCJ93]. primitives [AJ93]. priority [BW90]. private [SC94]. probabilistic [SF89]. Problem [PS86, TSF^{+}87, LfL00, PC78]. problem-solving [LfL00]. problems [CIP^{+}00]. Procedural [Sym85, OWG93]. procedure [CG93, CHK93]. Procedures [Geh80, Pag79, Abd75b, FF89, MS89]. process [AL85, MH07, PSW^{+}13, RK18, RS94, SP18, WHKK17]. process-oriented [RK18]. Processes [DH86, SBF80, YF83, yCH92, KJ12, LLvdW^{+}01, PMK^{+}18, Tal93a]. Processing [Led99e, AJ93, CM06, RH94, Tay96, WDCL08, ZLZ^{+}18]. Processor [HMHS18]. processors [BS92]. producer [CAS08]. producing [CCJ93]. Product [FAHC17, MAGD^{+}16, KR17, PMK^{+}18, SSS17]. product-line [PMK^{+}18]. production [LBGA18]. Productions [FM80, Dem75]. profiling [BBRR12, Kha11]. Program [BZ88, BS85, CDGM80, DK83, Fal97, PR10, PS86, Sha81, SC87, Tze12, YD78, Zob93, FO02, PRR12, PSW95, Rot92, Ste84, SSM10]. programmed [Seb89]. Programmer [DNR90]. Programmer-defined [DNR90]. programmers [Sch76]. Programming [ACS96, Ano07b, BCR11, BKSW09, CG84, Coo81, Cro79, CLM83, GAS17, GO88, HM18, Hul87, LP16, MOT84, MMC16, MB14, Ric16, RS87, Sym85, TSF^{+}87, Abd75a, Abd75b, AAH95, AD07, ACZ05, BT91, BDL^{+}12, BIMP17, Bry15, Bry16, BW90, CBTR17, CGG^{+}09, CM75, CHS16, CO98, Cia92, CG93, Coo98, CHHP91, CRPP00, Fal97, Fri92, GR91, HB16, HHS90, HG93, Hoa75, Hug85, Ier93, JM96, JD94, KKG92, KNW94, LMR93, LP97, LfL00, Liu93, Lou07, Mal10, MRO03, MŽ05, MW96, MP00, ND77, NL95, NPS17, ØK00, OM91, PGM84, PSW^{+}13, RGP98, Rus87, SB04, Sco91, Ste75, Sul75, Tay96, URI02, VBS^{+}14, Whi77, YG93, ZCM^{+}17, MMC15, MB13]. Programs [Fle86, GG82, KN85, LN91, AG17, BJ90, BER00, CY02, CCJ93, CG96, EL07, FC18, FNRR16, FO02, FF89, GG09, HC96, HS03, IPF82, Jos78, LC02, MVT^{+}16, MP17, NJLS12, PMS15, PMS16, PLS10, Rom97, SW77, Sar93, SF89, Thi93]. Prograph [MP85]. projects [GCH09]. Prolog [ACZ05, Anc13, CF88, Gan89a, NH93, SNA18, Tal93b]. prone [Ban17]. Proof [Ric80, GSP17]. proofs [Ber77, Liu93]. propagation [IvdS17]. properties [BSW15, BFPR04, UM17]. Proposal [Car78, Fle84, Liu88, ACS96, BJS93]. Proposals [Ten83, BEL77]. Prosper [LB89]. protected [PC15]. protection [Jos78]. PROTOB [BB91]. protocols [KKNS14]. prototyping [CS16, CHHP91, DS93, FL01, HZ96, Hoo89, LB89, Luq93, PJ91, WMP^{+}08]. provability [Har97]. provides [Coo98]. Providing [MGLFCP12]. proving [Fru10, UM17]. pruning [BJ14]. Publication [Ano09a, Ano10a]. Publisher [Ano06b, Ano11b]. Purba [Led99b]. pure [NPS17]. Purely [MO83]. Purpose [HR92]. pushdown [PTJM16].

QAS [KP78]. QoS [OOB18, PPK11]. QoS-enabled [PPK11]. QoSMOS [OOB18]. quadratic [AG17]. qualitative [LW75]. quality [ACG18, BSW15, GHD18]. Quantitative [Liu88]. queries [DQ09, KA17]. Query [BC84, CNGW09, MM82, PC85, BL92, BLM93, BL99, HC12, KA17]. Query-By-Rule [PC85]. querying [BRS90]. Question [KP78]. queue [CAS08]. quicksort [SJW94, SSJB96].

race [CS03]. ranking [UM17]. rapid [CHHP91, FL01, Luq93, WMP^{+}08]. rapid-prototyping [WMP^{+}08]. RASP [Dja88]. Re [GH07]. Re-scheduling [GH07]. reactive [PMS15, PMS16, PSW^{+}13]. readable [Jos78, RTMRK18]. Real [BGMT82, CMM85, LN91, Luq93, ABL17, BW90, DGU91, DRT97, GCH09, HL08, LS94]. Real-Time [BGMT82, CMM85, LN91, Luq93, ABL17, BW90, DGU91, DRT97, GCH09, HL08, LS94]. Reasoning [MR04, CLMT01, KH12]. Rebeca [JKK^{+}16]. recognition [PS94b]. recommender [PMK^{+}18]. Reconciling [Ber11]. reconfigurable [PDK^{+}09, PS94a]. reconfigurations [SMB15, SMB16]. Recording [SNP16]. recovery [HRS84, LCFÁ10]. recursion [FF90, Mor16, Thi93]. Recursive [Hor93, KK18, MS89, MPS90, SS09]. REDOSPLAT [RTMRK18]. Reducing [Ozt11]. reduction [DTXP13, Lee05]. reductions [Sis04]. Redundant [DH86, Pai16]. reference [CCGC12]. Refinement [BJ14, MP17, KG17]. refinements [EL07]. reflection [GWDD06, RDT08]. Reflections [Fel87]. reflective [KA07]. Regime [LS84]. region [BGH13]. Register [CAC^{+}81, Dha88, BM95, Dha90, Kes98, PS10, Zob93]. registers [VS95]. Regular [Anc13, PC78]. Relating [HC96]. Relational [BC84, BL92, BLM93, BMZM92, HHLv89, McL77]. relational-calculus [BL92]. relations [BRS90, CBTR18]. Relationship [SS79, DCA^{+}15, DCA^{+}16]. relationships [LW75, Sch75a]. relaxed [DMVY17]. reliability [ZP18]. Reliable [Ano07b, MST14]. remodularization [AC17]. removal [McC91, ZP18]. Removing [Lia92]. rendezvous [CO89]. reordering [GG09]. repair [HRS84]. Report [MP85]. representation [CPD93, Ear75]. repudiation [BC10]. requirement [ABL17]. requirements [NM17, RTMRK18]. Resilient [ABG^{+}05]. resolution [Rom97, Tay96]. Resource [CLM83, JM96, LCC07, jLtCxH09]. resources [CBTR17]. responsive [HZ96, VBS^{+}14]. restrictive [EL87]. Result [Geh80, WG83]. results [EvdSV^{+}15]. retargetable [BDB90, Gan89a]. retrieving [CNGW09]. reusability [SvdBV18]. reusable [VS95]. reuse [CKM^{+}18]. Reverse [LS84, Man01]. review [MAGD^{+}16]. Reviewers [Ano08, Ano12, Ano13, Ano15, Ano17, Ano09b, Ano10b, Ano11a]. Revised [Led99a]. revisions [FAHC17]. Revisiting [CHS16]. rewriting [SH15]. rich [MLW05]. Richard [Led99d]. richer [CV14]. Ring [GDD12]. robust [CC15, CC16, KR17]. Rofail [Led99c]. Role [FM80]. RPC [GH07]. RT [LS90, LS94]. RT-CDL [LS94, LS90]. RTC [MVT^{+}16]. Ruby [RT18]. Rule [CG96, PC85, CC95, YTC02]. Rule-based [CG96]. rules [CD82, FO10, GSP17, SNA18, VF82, Wil80]. Run [Joh81, Sar93, IF18, MRO03, SJW94]. Run-Time [Joh81, Sar93, IF18, MRO03, SJW94]. Runtime [DDT06, KA07, MVT^{+}16, RDT08].

S [PB84]. Safe

[Bou04, DCB^{+}17, BC02, DSW05, KMLS15]. SafeGPU [KPN17]. Safety [GCH09, Dre96, Fru10, PMS15, PMS16]. safety-critical [PMS15, PMS16]. samples [SK18]. Sanjiv [Led99b]. Sapaty [Led99e]. SASL [Sar93]. SAwUML [OK18]. SC-SystemJ [PMS15, PMS16]. scalable [MST14, OAB^{+}18]. SCAN [BAK89]. scenarios [ABS17]. scheduled [PS10]. scheduler [IF16]. scheduler-independent [IF16]. Scheduling [Kes98, VS95, GH07, PLS10]. Schema [CMM85]. scheme [LfL00, NH93, FFJ90, FF90, JL96, KS90, VSN^{+}17, Wan92, Won99]. Scheme-based [KS90]. scientific [PT09]. scope [VF82]. scoped [FF90]. scoping [FO10, KR17]. scripting [Ber11, PT09]. SDAC [ZLZ^{+}18]. SDL [Man01]. Seamless [NM17]. search [AC17, Ban17, FNRR16, Mal17]. search-based [Mal17]. seas [KLIN15]. sections [DLP15]. Secure [Ano07b, ZTLM13]. Security [BRB07, BCF02, BFPR04, CF02, Kır02, MCC17, MY17]. Segment [Wad80]. selection [Lus02, MC96, OM92]. selector [ÜA15, ÜA16]. self [PPK11, RR99]. self-adaptive [PPK11]. self-interpreters [RR99]. Semantic [COHW95, Fle84, Gan89b, Pag79, Tai79, BC93, Guo16, KHO14, SB79, Wil80]. Semantics [BER00, BEH86, MB85, Pag78, Wil81, AMA97, AMA98, AD07, BJ90, BG84, DLP15, DS93, GL95, JM96, KB75, LS94, Log09, Lou07, Mal93, OWG93, PC15, RS94, ZLZ^{+}18]. semantics-directed [DS93]. semi [GSP17]. semi-algebraic [GSP17]. semistructured [CNGW09, DQ09]. sensitive [HWM13, IF18, NN17, SM94]. separation [Fal97]. Seque [GO88]. SequenceL} [Coo98]. Sequences [GO88, WG83, Nil90]. Sequential [DH86]. Server [Led99b, Led99a]. service [CS16, ONB^{+}18]. service-oriented [CS16]. Services [PPK11, GH07, MZC07]. servicing [OBGK02]. SESPOOL [ND77]. set [Dja88, PLS10, SK18]. sets [GSP17]. Shallow [Sam79, SA16]. shared [BT91, OBGK02]. Sharing [DMVD16, PLS10]. shell [GCH09]. Shellsort [SJW94]. shortcomings [EP89, PE88]. shot [AD07]. should [Sch76].

Side [IR95, CJD17]. Side-effect [IR95].

SIGPLAN [HB16]. Simple

[Abd75a, War78, FDH08, Tze12, CDGM80]. simplification [Han97]. SIMULA [PGM84, Sch78]. Simulating [TKH99]. Simulation [Hoo87, Hoo89, KS90, Sal92]. single [AD07, Dem75, MGLFCP12, SIK09]. single-user [MGLFCP12]. singleton [Alj16]. SIR [FO02]. size [SJW94, SS09]. Skeleton [AD07]. Skeleton-based [AD07]. sketch [RR99]. sketch-based [RR99]. skipping [KK18]. SL5 [Han78]. SLE [CPPV15]. SLIPS [GS86]. slope [DKK18]. SMALLTALK [GL95, ABG^{+}05, DOZ06, DDT06, GDD12, SD06]. SMALLTALK-80 [GL95]. smart [SK14]. Snobol4 [Gri83, Pag78]. Software [CCRS18, CPPV15, FAHC17, HRW18, HR91, MAGD^{+}16, RS83, Rid79a, RPB09, Zav86, AC17, AC18, Ban17, CBTR17, yCH92, CL89, ECB12, FO10, Hoo89, HL08, KR17, KKG15, KS90, Mal17, OK18, RH18, Rid79b, SK14, SSS17, TBKG04, Zdu06, ZP18, vOKF01]. Solmar [BS78]. solution [jLtCxH09, PC78, PSW^{+}13]. solutions [CJD17]. solve [DQ09]. Solver [BKL18]. solving [CIP^{+}00, EP89, LfL00]. Some [KY75, Ten83]. Sonar [BS78]. Sons [Led99b, Led99c, Led99d, Led99e, Led99a]. sophisticated [BDPW08]. Source [FF75, Ban17, GDD12, MT05, SNP16]. sources [ÁdLNW18]. space [OAB^{+}18, Ozt11]. Spaces [ACS96]. sparse [KHO14]. SPEC [CIP^{+}00, Ber91b]. Special [ACG18, Bry15, Bry16, CPPV15, DL17, DdLP17, GAS17, HM18, HB16, HR92, KGS17, LP16, Mal17, MB13, MB14, Zuc04, CB93b, Lou07, SD06]. Special-Purpose [HR92]. specialization [dACSAP14, Kha10, Kha11]. specific [BKG^{+}08, CJD17, CDGN15, CSdL16, ECB12, FFMB11, GAGdL17, LDN18, MMC16, PLDD15, PSW^{+}13, RTMRK18, SNA18, SvdBV18, SA16, dlVGSZS18, MMC15]. Specification [BS85, IF16, Jal92, Orm83, Pag79, Ber91b, BMZM92, CIP^{+}00, DGU91, GVvdP^{+}01, GCH09, Hat91, HZ96, LB89, MY17, Nym95]. specification-PEARL [GCH09]. specified [BCR11]. Specifying [Wil81, CY02, CC95, Wil80]. speeding [KKG15]. SPIN [OK18]. SPITBOL [Tha77]. Spy [BBRR12]. SQL [Led99b, BRS90, KMLS15]. SR [CO89, FO02, HO90, MRO03]. SR-like [MRO03]. SSA [Pai16]. stack [SS09]. STAPLE [Ste75]. State [Pun01, ALS^{+}18, DMVD16, Tze12]. Stateful [BDNW08]. Statement [Car78, FF75]. statements [FF86, PNF^{+}18]. Static [Bai86, HV93, Wil81, BC93, GBZ09, IF16, LR17, Pen05, PCG16, SIK09, Wil80]. Statistical [JKK^{+}16, RD78]. Step [CCF15]. steps [KY75]. stepwise [EL07]. Stochastic [Bar82]. store [Dha90]. stores [JD94]. story [WHKK17]. strategies [VF82]. strategy [CCJ93, HGC^{+}09, RW09]. stream [CDW09, FNRR16, Nil90]. streams [BJS93, FFJ90]. strict [Tre00]. strictness [SvE16]. String [CF88, Liu88, BGH13, KB75]. strongly [YG93]. strongly-typed [YG93]. Structural [Sha80, MP17, Thi93]. Structure [Geh79, Ear75, PRD02, PSW95, Zob93]. Structured [CL78, Coh78, Her76, SC87, Bas75, Ste75, Sul75]. Structures [Fle78, Han78, YD78, FC18, HG93, JG89, YF98]. Structuring [CG93, Fle86, JO11]. studies [DKK18]. Study [Zav86, BMN^{+}18, HC05, KR17, LC02, MKPW06, NPS17, SW77, Sar93, SJW94, SLS18, VBDPM16]. studying [VMD18]. Style [Pet78, PRR12, Fle86]. stylesheets [GGK^{+}11]. sub [SS93]. subexpressions [RW09]. Sublist [Jay92]. Subset [Pag78]. substring [CB93a]. suite [DTXP13, OOB18]. Supercombinator [SK18]. supervenience [Rez12]. Support [Ano07b, BKSW09, FO02, Hoo89, LCFÁ10, RBY^{+}05]. supported [Rod15]. Supporting [CG84, FT15, FT16, MZC07]. supports [Nil90]. survey [Cia92, HRS84, Rod15, Tal93b, ZP04]. swapping [PBDF12]. switched [AG17]. symbiosis [VMD09]. Symbolic [ALR15, GMMP89, CPD93, MST14, PNF^{+}18]. symbols [Dem75]. SymGridPar2 [MST14]. symmetry [Sis04]. Symposium [Bry15, Bry16, MB13, MB14]. Synchronisation [MW96]. Synchronization [DH86, ØK00, YF83]. synchronous [MR01, MH07]. Syntactic [FM80, HRS84]. syntactical [PC78]. Syntax [Sha75, AMA97, Dai94, DP98, Her76, Noo85, RD78, ZGE85]. synthesis [AG17, CBTR17, HL08, Man01, MS93]. Synthesizing [ABS17]. System [CDGM80, CMM85, KP78, MZGT85, Rid79a, CHHP91, DCA^{+}15, DCA^{+}16, dOG09, HSS88, HHS90, KS90, MSRG10, MRO03, McL77, Pen05, Rid79b, SS93, Whi77, FF75]. System/370 [FF75]. Systematic [SLS18, MAGD^{+}16]. SystemJ [MSRG10, PMS15, PMS16]. Systems [Ano07a, BB91, Bar82, BGMT82, Cro79, Hul87, Orm83, Spr79, AC17, ABL17, BGH13, BGM^{+}18, BMN^{+}18, DGU91, DPP10, Dre96, GMMP89, HZ96, HL08, JPB^{+}08, KGS17, MRO03, MR04, ND77, PMS15, PMS16, PMK^{+}18, PGT^{+}96, SRRB10, WHKK17, ZTLM13, ZP04].

tables [Li96]. Tailorable [Zdu06]. Taking [LDG09]. TaKo [MGLFCP12]. target [PCB^{+}18]. targets [Buf17]. Task [CMM85, MZGT85, PLS10]. tasking [GSX99]. Technique [Cel81, Sha81, KHO14, RR99]. Techniques [Spr79, DP98, Dha88, Dre96, FW87, KR98, Mal17]. technology [MLW05, PT09]. Telegram [TSF^{+}87]. template [SLS18]. template-based [SLS18]. Temporal [KJTA17, BL99, FL01, UM17]. Terminal [MOT84, TC81]. termination [Tal93a]. terms [NH93]. Test [CD81, ABS17, DTXP13, Guo16]. test-suite [DTXP13]. Testing [BS85, Ric80, Was79, Jal92, SRT17]. Text [Ree84, ZA87, AA89, MB75, SRT17, Thi82]. text-based [SRT17]. Thank [Ano17]. their [BDNW08, Dre96, Mal17, OK18]. Theory [Fle78, PS86, KGS17, SS92]. thread [DLP15]. threading [Li96]. threads [VMD09]. Three [War78]. Tier [Led99c, Led99a]. Time [BLM93, BGMT82, CMM85, Joh81, LN91, ABL17, BL99, BJ14, BW90, dACSAP14, DGU91, DRT97, GCH09, HL08, IF18, JPB^{+}08, LS94, Luq93, MRO03, RGP98, Sar93, SJW94, VMD18]. Time-based [BLM93]. Timed [HF87, JKK^{+}16]. timetabling [RTMRK18]. tolerant [CL89]. Tony [Led99c]. too [EL87]. tool [DCA^{+}15, DCA^{+}16, FL01, OOB18]. toolbox [AMR18]. Tools [Zav86, BDPW08, GDD12, Hor17, PRR12, RH94, SRT17, WD04]. top [Buf17]. top-level [Buf17]. trace [HWM13, Log09, PC15, VMD18]. trace-based [VMD18]. track [Bry15, Bry16, MB13, MB14]. tracking [Kır02]. trait [CDW09]. trait-based [CDW09]. Traits [CDW09, BDNW08, CV16]. Transactional [RN09, CM11]. Transformation [Sha81, ALR15, DDT06, Kha11]. transformational [SMdSB09]. transformations [BSW15, ESG16, Pen14, PR10, Sar94]. Transforming [GGK^{+}11]. Transition [CF79]. transitions [ALS^{+}18]. Translating [MO83]. Translation [MT82, CYS^{+}15, SIK09]. Translator [SM89]. transmission [Sch75b]. transparent [IF18, MGLFCP12].

Transputers [SS93]. Trapezoid [CCF15]. tree [BDB90, KG17, Mal93]. tree-manipulation [Mal93]. trees [HT13, LRB^{+}11, Li96, Noo85, VS95]. Trends [LP16, HM18]. Tuple [ACS96]. Two [BEH86, GZ87, Pag79, FT15, FT16, ZP18]. two-dimensional [FT15, FT16]. Two-Level [BEH86, Pag79]. two-phase [ZP18]. TXL [CHHP91]. Type [FC18, PS86, Bai90, BGH13, CCT08, FW87, Fru10, dOG09, Ier93, JL96, KMLS15, KDM03, MS93, MP92, Nil90, Pen05, SM94, ZCM^{+}17]. type-checking [CCT08, Ier93, JL96, MS93]. type-safe [KMLS15]. type-sensitive [SM94]. typed [Ber11, CCGC12, YG93]. Types [Bai87, Fle84, GZ87, BRB07, BMZM92, Fis88, HDN09, Jal92, JO11, SvE16]. typing [BC02, GBZ09, KNW94].

ubiquitous [HL08]. UML

[ABS17, GCH09, OK18, PCG16, PNF^{+}18]. UML-based [OK18]. UML/OCL [PNF^{+}18]. Unanticipated [RDT08, WD04]. unbounded [Nil90]. Undecidability [Geh80]. undecidable [CS03]. Understander [NB84]. unexpected [DG94]. Unified [Lus02, FDH08, Rod15]. Uniform [DSW05, CL97, PS10]. unifying [Ber91a, GDD12]. unit [SRT17]. Units [Geh77]. UNITY [RM93]. University [NB84]. unnesting [KA17]. untrusted [FO10]. updates [HV94]. Updating [CCRS18]. Usability [BAG18, MM82]. usage [PS10]. Use [Ken78, MO83, AL85, RT18, Thi82, BAG18]. Use-Definition [Ken78]. USE-ME [BAG18]. used [Sli17]. User [WHKK17, LBGA18, MGLFCP12]. User-story [WHKK17]. users [NPS17]. Using [BC84, CDF18, FL87, Pag79, Wad80, WF78, Alj16, ABS17, BZ88, BDB90, Bou04, BC13, BKG^{+}08, BC10, CV16, CDGN15, CNGW09, Dha88, Dha90, GVvdP^{+}01, GMMP89, KMLS15, LDN18, OK18, PM18, PCB^{+}18, RBY^{+}05, SSS17, SNP16, WD04]. usually [Dha90]. Utility [jLtCxH09]. Utility-driven [jLtCxH09]. Utilizing [BS78].

validation [CYS^{+}15, CSdL16, GHD18, PNF^{+}18]. value [dACSAP14, IF18, Kha11, Sch75b]. value-sensitive [IF18]. values [Nil90]. variability [FAHC17, SSS17, SP18]. variability-aware [SSS17]. Variable [Bai86, DK92]. variables [KJ12]. VDL [MB85]. Vector [HMHS18, CM06, JO11]. verifiable [BKYV80]. Verification [DL17, Was79, CD82, DMVY17, DKK18, EHMO91, GHD18, HL08, KG17, PCG16, PNF^{+}18, SMB15, SMB16]. Verifying [BGH13, BFPR04, PMS15, PMS16, vOKF01]. versa [GGK^{+}11]. Version [Man78]. Very [Tuc75, CCJ93, Sch75b, Sch75a]. via [CBTR17, CAC^{+}81]. vice [GGK^{+}11]. Vienna [KPP93]. view [Coo98, LDG09, SB79]. viewing [FL92]. viewpoint [Tuc75]. Views [SS79, MKPW06, TBKG04]. visibility [BDNW05]. visibly [PTJM16]. Visual [FL01, Led99c, CDF18, MP00, SRT17, AMA97, AMA98]. Visualising [LLvdW^{+}01]. visualization [Hor17]. visualization/analysis [Hor17]. visualizing [vOKF01]. VMCAI’03 [Zuc04]. Volume [Ano02d, Ano05f, Ano05g, Ano06c, Ano99, Ano00, Ano01b].

way [Coo98, FFJ90]. Weak [BKL18]. web [CBTR17, CJD17, MCC17, VSN^{+}17, Mal10, MZC07, PPK11]. web-based [CBTR17]. well [BER00]. well-moded [BER00]. whole [WDCL08]. whole-image [WDCL08]. Widening [CZ11]. Wiley [Led99b, Led99c, Led99d, Led99e, Led99a]. Within [Tai79]. work [CDW09, Led99a]. workbenches [EvdSV^{+}15]. works [Jos78]. workshops [HB16]. worst [Bli94, BJ14]. worst-case [BJ14]. write [CS03]. writing [HSS88]. Written [MB85]. WS [KJ12].

XML [CJD17]. XML-based [CJD17]. XQuery [GGK^{+}11]. XSL [Pen14]. XSLT [GGK^{+}11].

Y2K} [Led99d]. Younger [Man78].

Z [PE88]. Zero [GBZ09].

References

Abi-Akar:1989:ATF

[AA89] Ramez Abi-Akar. Arabic text formatter: comments and implementation. Computer Languages, 14(1):53–60, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Armentano:2009:FAP

[AA09] Marcelo G. Armentano and Analía A. Amandi. A framework for attaching personal assistants to existing applications. Computer Languages, Systems and Structures, 35(4):448–463, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                                    http://www.sciencedirect.
                                                    com/science/article/pii/
                                                    S1477842408000444.

Al-AAli:1995:DAP

[AAH95] Mansoor Al-A’Ali and Mohammed Hamid. Design of an Arabic programming language (ARABLAN). Computer Languages, 21(3-4):191– 200, October–December 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Abdali:1975:LMPa

[Abd75a] S. K. Abdali. A lambdacalculus model of programming languages. I. simple constructs. Computer Languages, 1(4):287–301, ???? 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Abdali:1975:LMPb

[Abd75b] S. K. Abdali. A lambdacalculus model of programming languages. II. jumps and procedures. Computer Languages, 1 (4):303–320, ???? 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Andersen:2005:DIE

[ABG^{+}05] Jakob R. Andersen, Lars Bak, Steffen Grarup, Kasper V. Lund, Toke Eskildsen, Klaus Marius Hansen, and Mads Torgersen. Design, implementation, and evaluation of the Resilient Smalltalk embedded platform. Computer Languages, Systems and Structures, 31(3–4):127– 141, October/December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Ashamalla:2017:MDA

[ABL17] Amir Ashamalla, Ghassan Beydoun, and Graham Low. Model driven approach for real-time requirement analysis of multi-agent systems. Computer Languages, Systems and Structures, 50(??):127– 139, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300823.

Arora:2017:STS

[ABS17] Vinay Arora, Rajesh Bhatia, and Maninder Singh. Synthesizing test scenarios in UML activity diagram using a bio-inspired approach. Computer Languages, Systems and Structures, 50(??):1–19, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300209.

Amarjeet:2017:HSB

[AC17] Amarjeet and Jitender Kumar Chhabra. Harmony search based remodularization for object-oriented software systems. Computer Languages, Systems and Structures, 47 (Part 2)(??):153– 169, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S1477842416301245.

Amarjeet:2018:FAF

[AC18] Amarjeet and Jitender Kumar Chhabra. FP-ABC: Fuzzy-Pareto dominance driven artificial bee colony algorithm for many-objective software module clustering. Computer Languages, Systems and Structures, 51(??):1–21, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416302044.

Amaral:2018:SIQ

[ACG18] Vasco Amaral, Jordi Cabot, and Miguel Goulão. Special issue on quality in modeldriven engineering. Computer Languages, Systems and Structures, 54(??):472–473, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300393.

Ambriola:1996:PMM

[ACS96] Vicenzo Ambriola, Giovanni A.

Cignoni, and Laura Semini. A proposal to merge multiple tuple spaces, object orientation, and logic programming. Computer Languages, 22(2–3):79– 93, July–October 1996. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Amandi:2005:JFB

[ACZ05] Analía Amandi, Marcelo Campo, and Alejandro Zunino. Java-Log: a framework-based integration of Java and Prolog for agent-oriented programming. Computer Languages, Systems and Structures, 31(1): 17–33, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Aldinucci:2007:SBP

[AD07] Marco Aldinucci and Marco Danelutto. Skeleton-based parallel programming: Functional and parallel semantics in a single shot. Computer Languages, Systems and Structures, 33(3–4):179–192, October/December 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842406000315.

Angel:2018:AMA

[ÁdLNW18] Mora Segura Ángel, Juan de Lara, Patrick Neubauer, and Manuel Wimmer. Automated modelling assistance by integrating heterogeneous information sources. Computer Languages, Systems and Structures, 53(??):90–120, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417301690.

Adje:2017:ASI

[AG17] Assalé Adjé and Pierre-Loïc Garoche. Automatic synthesis of k-inductive piecewise quadratic invariants for switched affine control programs. Computer Languages, Systems and Structures, 47 (part 1)(??):44– 61, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000937.

Axford:1993:LPP

[AJ93] Tom Axford and Mike Joy. List processing primitives for parallel computation. Computer Languages, 19(1):1–17, January 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Amtoft:2002:OCA

[AKPG02] Torben Amtoft, Assaf J.

Kfoury, and Santiago M. Pericas-Geertsen. Orderly communication in the Ambient Calculus. Computer Languages, Systems and Structures, 28(1): 29–60, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Ahson:1985:UFL

[AL85] S. I. Ahson and S. S. Lamba.

The use of FORTH language in process control. Computer Languages, 10(3-4):179–187, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Aljasser:2016:IDP

[Alj16] Khalid Aljasser. Implementing design patterns as parametric aspects using ParaAJ: the case of the singleton, observer, and decorator design patterns. Computer Languages, Systems and Structures, 45(??): 1–15, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000913.

Arusoaie:2015:SEB

[ALR15] Andrei Arusoaie, Dorel Lucanu, and Vlad Rusu. Symbolic execution based on language transformation. Computer Languages, Systems and Structures, 44 (part A)(??):48– 71, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241500055X.

Adesina:2018:IFA

[ALS^{+}18] Opeyemi O. Adesina, Timothy C. Lethbridge, Stéphane S. Somé, Vahdat Abdelzad, and Alvine Boaye Belle. Improving formal analysis of state machines with particular emphasis on and-cross transitions. Computer Languages, Systems and Structures, 54(??):544– 585, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300301.

Al-Mulhem:1997:VOS

[AMA97] Muhammed Al-Mulhem and Shahid Ali. Visual Occam: syntax and semantics. Computer Languages, 23(1):1–24, April 1997. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Al-Mulhem:1998:FSV

[AMA98] Muhammed Al-Mulhem and Shahid Ali. Formal semantics of Visual Occam. Computer Languages, 24(2):99– 114, July 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/18/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/18/19/article.
pdf.

Amir-Mohammadian:2013:NPP

[AMF13] Sepehr Amir-Mohammadian and Mehran S. Fallah. Noninterference in a predicative polymorphic calculus for access control. Computer Languages, Systems and Structures, 39(3):109–120, October 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241300016X.

Azzini:2018:DMP

[AMR18] Ivano Azzini, Ronal Muresano, and Marco Ratto. Dragonfly: a multi-platform parallel toolbox for MATLAB/ Octave. Computer Languages, Systems and Structures, 52(??): 21–42, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417300660.

Ancona:2013:RCP

[Anc13] Davide Ancona. Regular corecursion in Prolog. Computer Languages, Systems and Structures, 39(4):142–162, December 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842413000158.

Anonymous:1988:CCP

[Ano88] Anonymous. C code for the PC. Computer Languages, 5 (2):52–??, February 1988. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). advertisement.

Anonymous:1999:VI

[Ano99] Anonymous. Volume 25 index. Computer Languages, 25(4):??, December 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

                                        /www.elsevier.com/gej-ng/

10/15/18/28/30/26/abstract.
html.

Anonymous:2000:IV

[Ano00] Anonymous. Index to Volume 26. Computer Languages, 26(2–4):??, July 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

/www.elsevier.com/gej-ng/
10/15/18/50/31/31/abstract.
html.

Anonymous:2001:EIG

[Ano01a] Anonymous. Editorial: Introduction by the Guest Editor. Computer Languages, 27(1–3):1–2, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/54/27/27/
abstract.html.

Anonymous:2001:IV

[Ano01b] Anonymous. Index to Volume 27, 2001. Computer Languages, 27(4):??, December 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/54/31/29/
abstract.html.

Anonymous:2002:EBa

[Ano02a] Anonymous. Editorial Board.

Computer Languages, Systems and Structures, 28(2):CO2, December 15, 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2002:EBb

[Ano02b] Anonymous. Editorial Board. Computer Languages, Systems and Structures, 28(3):CO2, October 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2002:IFC

[Ano02c] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 28(4):CO2, December 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2002:VC

[Ano02d] Anonymous. Volume contents. Computer Languages, Systems and Structures, 28(4):??, December 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2003:IFCa

[Ano03a] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 29(1–2):CO2, April/July 2003. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2003:IFCb

[Ano03b] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 29(3):CO2, October 2003. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2003:IFCc

[Ano03c] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 29(4):CO2, December 2003. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2004:IEB

[Ano04a] Anonymous. IFC — Editorial Board. Computer Languages, Systems and Structures, 30(3–4):CO2, October/ December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2004:IFC

[Ano04b] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 30(1–2):CO2, April/July 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:AI

[Ano05a] Anonymous. Author index.

Computer Languages, Systems and Structures, 31(1):??, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:IFCa

[Ano05b] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 31(1):CO2, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:IFCb

[Ano05c] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 31(2):CO2, July 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:IFCc

[Ano05d] Anonymous. Inside front cover — Editorial Board. Computer Languages, Systems and Structures, 31(3–4):CO2, October/ December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:KI

[Ano05e] Anonymous. Keyword index. Computer Languages, Systems and Structures, 31(1):??, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:VC

[Ano05f] Anonymous. Volume contents. Computer Languages, Systems and Structures, 31(1):??, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2005:VCA

[Ano05g]  Anonymous.  Volume con-

tents, author and keyword index. Computer Languages, Systems and Structures, 31(3–4): ??, October/December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2006:A

[Ano06a] Anonymous. Announcement.

Computer Languages, Systems and Structures, 32(2–3):83, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000370.

Anonymous:2006:PN

[Ano06b] Anonymous. Publisher’s note.

Computer Languages, Systems and Structures, 32(1):1, April 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Anonymous:2006:VCA

[Ano06c]  Anonymous.  Volume con-

tents, author and keyword index. Computer Languages, Systems and Structures, 32(4): ??, December 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000108.

Anonymous:2007:CPE

[Ano07a]  Anonymous.  Call for pa-

pers: Embedded systems: Compiler-architecture interaction. Computer Languages, Systems and Structures, 33 (1):32, April 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000224.

Anonymous:2007:CPP

[Ano07b]  Anonymous.  Call for pa-

pers: Programming language and compiler support for secure and reliable computing. Computer Languages, Systems and Structures, 33(1): 33–34, April 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842406000236.

Anonymous:2008:R

[Ano08] Anonymous. Reviewers 2007. Computer Languages, Systems and Structures, 34(2–3):43, July/October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842407000346.

Anonymous:2009:EBP

[Ano09a] Anonymous. Editorial Board / publication information. Computer Languages, Systems and Structures, 35(4): ??, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842409000165.

Anonymous:2009:LR

[Ano09b] Anonymous. List of reviewers 2008. Computer Languages, Systems and Structures, 35 (2):99, July 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000432.

Anonymous:2010:EBP

[Ano10a] Anonymous. Editorial Board / publication information. Computer Languages, Systems and Structures, 36(4): ??, December 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000163.

Anonymous:2010:LR

[Ano10b] Anonymous. List of reviewers 2009. Computer Languages, Systems and Structures, 36 (1):1, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000232.

Anonymous:2011:LR

[Ano11a] Anonymous. List of reviewers 2010. Computer Languages, Systems and Structures, 37 (1):1, April 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000357.

Anonymous:2011:PN

[Ano11b] Anonymous. Publisher’s note.

Computer Languages, Systems and Structures, 37(4): 204, October 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842411000297.

Anonymous:2012:R

[Ano12] Anonymous. Reviewers 2011. Computer Languages, Systems and Structures, 38(1): 122, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842411000509.

Anonymous:2013:R

[Ano13] Anonymous. Reviewers 2012. Computer Languages, Systems and Structures, 39(1): 48, April 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842412000486.

Anonymous:2015:R

[Ano15] Anonymous. Reviewers 2014. Computer Languages, Systems and Structures, 41(??): 66, April 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000202.

Anonymous:2017:TYR

[Ano17]  Anonymous.  Thank you

reviewers: 2016. Computer Languages, Systems and Structures, 47 (part 1)(??): i–ii, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301592.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board.

Computer Languages, Systems and Structures, 52(??): ii, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301975.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board.

Computer Languages, Systems and Structures, 53(??): ii, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300563.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board.

Computer Languages, Systems and Structures, 54(??): ii, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418301209.

Barisic:2018:UDD

[BAG18] Ankica Barisić, Vasco Amaral, and Miguel Goulão. Usability driven DSL development with USE-ME. Computer Languages, Systems and Structures, 51(??):118– 157, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417300477.

Bailes:1986:SCV

[Bai86] Paul A. Bailes. Static checking of variable handling in Dijkstra’s Guarded Commands Language. Computer Languages, 11(3-4):123–142, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bailes:1987:GFL

[Bai87] Paul A. Bailes. G: a functional language with generic abstract data types. Computer Languages, 12(2):69–94, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bailes:1990:HDG

[Bai90] Paul A. Bailes. The hierarchical development of a generic type mechanism for functional languages. Computer Languages, 15(1):1–26, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bourbakis:1989:PIS

[BAK89] Nikolaos G. Bourbakis, Christos Alexopoulos, and Allen Klinger. A parallel implementation of the SCAN language. Computer Languages, 14(4):239–254, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bansal:2017:EAS

[Ban17] Ankita Bansal. Empirical analysis of search based algorithms to identify change prone classes of open source software. Computer Languages, Systems and Structures, 47 (Part 2)(??): 211–231, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301397.

Barman:1982:LLS

[Bar82] Mervyn Barman. Lass — a language for stochastic systems. Computer Languages, 7(3-4): 125–146, ???? 1982. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Basili:1975:SAL

[Bas75] V. R. Basili. A structured approach to language design. Computer Languages, 1(3): 255–273, September 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bays:1976:ADL

[Bay76] J. C. Bays. An animation description language. Computer Languages, 2(1-2):1–8, ???? 1976. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Baldassari:1991:POO

[BB91] Marco Baldassari and Giorgio Bruno. PROTOB: An object oriented methodology for developing discrete event dynamic systems. Computer Languages, 16(1):39–63, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bergel:2012:SFC

[BBRR12] Alexandre Bergel, Felipe Bañados, Romain Robbes, and David Röthlisberger. Spy: a flexible code profiling framework. Computer Languages, Systems and Structures, 38(1): 16–28, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000327.

Barrett:2015:AIC

[BBT15] Edd Barrett, Carl Friedrich Bolz, and Laurence Tratt. Approaches to interpreter composition. Computer Languages, Systems and Structures, 44 (Part C)(??):199– 217, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241500007X.

Barrett:2016:AIC

[BBT16] Edd Barrett, Carl Friedrich Bolz, and Laurence Tratt. Approaches to interpreter composition. Computer Languages, Systems and Structures, 45(??): 199–217, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241500007X.

Bossi:1984:UFQ

[BC84] Annalisa Bossi and C. Chezzi.

Using Fp as a query language for relational data-bases. Computer Languages, 9(1):25–37, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Barnard:1988:SPL

[BC88] David T. Barnard and James R.

Cordy. SL parses the LR languages. Computer Languages, 13(2):65–74, ???? 1988. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Barnard:1989:AGS

[BC89] David T. Barnard and James R.

Cordy. Automatically generating SL parsers from LL(1) grammars. Computer Languages, 14(2):93–98, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bailes:1993:FGT

[BC93] Paul A. Bailes and Trevor Chorvat. Facet Grammars: towards static semantic analysis by context-free parsing. Computer Languages, 18(4):251– 271, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bugliesi:2002:BTS

[BC02] Michele Bugliesi and Giuseppe Castagna. Behavioural typing for safe ambients. Computer Languages, Systems and Structures, 28(1):61–99, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Bruso:2010:NRA

[BC10] Mayla Brusò and Agostino Cortesi. Non-repudiation analysis using LySa with annotations. Computer Languages, Systems and Structures, 36(4):352–377, December 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241000014X.

Brandner:2013:EPC

[BC13] Florian Brandner and Quentin Colombet. Elimination of parallel copies using code motion on data dependence graphs. Computer Languages, Systems and Structures, 39(1): 25–47, April 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842412000346.

Braghin:2002:SBM

[BCF02] Chiara Braghin, Agostino Cortesi, and Riccardo Focardi. Security boundaries in mobile ambients. Computer Languages, Systems and Structures, 28(1):101–127, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Braghin:2004:NAM

[BCF^{+}04] Chiara Braghin, Agostino Cortesi, Riccardo Focardi, Flaminia L. Luccio, and Carla Piazza. Nesting analysis of mobile ambients. Computer Languages, Systems and Structures, 30(3–4):207–230, October/December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Bergenti:2011:PPS

[BCR11] F. Bergenti, L. Chiarabini, and G. Rossi. Programming with partially specified aggregates in Java. Computer Languages, Systems and Structures, 37(4):178–192, October 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000169.

Balachandran:1990:ERC

[BDB90] A. Balachandran, D. M.

Dhamdhere, and S. Biswas. Efficient retargetable code generation using bottom-up tree pattern matching. Computer Languages, 15(3):127–140, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Barbosa:2012:TPM

[BDL^{+}12] Jorge Barbosa, Fabiane Dillenburg, Gustavo Lermen, Alex Garzão, Cristiano Costa, and João Rosa. Towards a programming model for contextaware applications. Computer Languages, Systems and Structures, 38(3):199–213, October 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842412000115.

Bergel:2005:CCV

[BDNW05] Alexandre Bergel, Stéphane Ducasse, Oscar Nierstrasz, and Roel Wuyts. Classboxes: controlling visibility of class extensions. Computer Languages, Systems and Structures, 31(3– 4):107–126, October/December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Bergel:2008:STT

[BDNW08] Alexandre Bergel, Stéphane Ducasse, Oscar Nierstrasz, and Roel Wuyts. Stateful traits and their formalization. Computer Languages, Systems and Structures, 34(2–3):83–108, July/October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000140.

Bergel:2008:CSD

[BDPW08] Alexandre Bergel, Stéphane Ducasse, Colin Putney, and Roel Wuyts. Creating sophisticated development tools with OmniBrowser. Computer Languages, Systems and Structures, 34(2–3):109–129, July/October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000152.

Bryant:1986:TGI

[BEH86] Barrett R. Bryant, Balanjaninath Edupuganty, and Lee S. Hull. Two-level grammar as an implementable metalanguage for axiomatic semantics. Computer Languages, 11(3-4): 173–191, ???? 1986. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Berry:1977:PDAa

[BEL77] D. M. Berry, Z. Erlich, and C. J. Lucena. Pointers and data abstractions in high level languages. I. language proposals. Computer Languages, 2 (4):135–148, ???? 1977. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Berry:1977:PDAb

[Ber77] D. M. Berry. Pointers and data abstractions in high level languages. II. correctness proofs. Computer Languages, 2(4): 149–170, ???? 1977. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bermudez:1991:UML

[Ber91a] Manuel E. Bermudez. A unifying model for lookahead LR parsing. Computer Languages, 16(2):167–178, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Berzins:1991:BSS

[Ber91b] Valdis Berzins. Black-box specification in Spec. Computer Languages, 16(2):113–127, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bossi:2000:SWM

[BER00] Annalisa Bossi, Sandro Etalle, and Sabina Rossi. Semantics of well-moded input-consuming logic programs. Computer Languages, 26(1):1–25, April 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/15/18/50/25/25/
                                        abstract.html.

Bergel:2011:RMO

[Ber11] Alexandre Bergel. Reconciling method overloading and dynamically typed scripting languages. Computer Languages, Systems and Structures, 37(3): 132–150, July 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000030.

Barroso:1978:IDD

[BF78] P. B. Barroso and A. L. Furtado. Implementing a data definition facility driven by graph grammars. Computer Languages, 3(2):65–74, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bossi:2004:VPS

[BFPR04] Annalisa Bossi, Riccardo Focardi, Carla Piazza, and Sabina Rossi. Verifying persistent security properties. Computer Languages, Systems and Structures, 30(3–4):231–258, October/December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Bryant:1984:ILD

[BG84] Barrett R. Bryant and A. A.

Grau. An intermediate language to define dynamic semantics. Computer Languages, 9(3-4):149–159, ???? 1984. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Beringer:2013:VPS

[BGH13] Lennart Beringer, Robert Grabowski, and Martin Hofmann. Verifying pointer and string analyses with region type systems. Computer Languages, Systems and Structures, 39(2): 49–65, July 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241300002X.

Braz:2018:CAP

[BGM^{+}18] Larissa Braz, Rohit Gheyi, Melina Mongiovi, Márcio Ribeiro, Flávio Medeiros, Leopoldo Teixeira, and Sabrina Souto. A change-aware per-file analysis to compile configurable systems with ifdefs. Computer Languages, Systems and Structures, 54(??):427– 450, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300441.

Berry:1982:LCR

[BGMT82] D. M. Berry, C. Ghezzi, D. Mandrioli, and F. Tisato. Language constructs for realtime distributed systems. Computer Languages, 7(1):11–20, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bergenti:2017:AOM

[BIMP17] Federico Bergenti, Eleonora Iotti, Stefania Monica, and Agostino Poggi. Agent-oriented model-driven development for JADE with the JADEL programming language. Computer Languages, Systems and Structures, 50(??):142–158, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301075.

Bolot:1990:FSP

[BJ90] Jean-Chrysostome Bolot and Pankaj Jalote. Functional semantics of programs with exceptions. Computer Languages, 15(4):251–265, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Brandner:2014:RWC

[BJ14] Florian Brandner and Alexander Jordan. Refinement of worst-case execution time bounds by graph pruning. Computer Languages, Systems and Structures, 40(3–4): 155–170, October/December 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842414000359.

Bailes:1993:PGS

[BJS93] Paul A. Bailes, Dan B. Johnston, and Eric J. Salzman. A proposal for a genuinely-lazy streams facility for Ada. Computer Languages, 18(1):31–55, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Brichau:2008:ASM

[BKG^{+}08] Johan Brichau, Andy Kellens, Kris Gybels, Kim Mens, Robert Hirschfeld, and Theo D’Hondt. Application-specific models and pointcuts using a logic metalanguage. Computer Languages, Systems and Structures, 34(2–3):66–82, July/ October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842407000139.

Blanchard:2018:MCB

[BKL18] Allan Blanchard, Nikolai Kosmatov, and Frédéric Loulergue. MMFilter: A CHRbased solver for generation of executions under weak memory models. Computer Languages, Systems and Structures, 53(??):121–142, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417301665.

Bloom:2009:FPL

[BKSW09] Bard Bloom, Paul Keyser, Ian Simmonds, and Mark Wegman. Ferret: Programming language support for multiple dynamic classification. Computer Languages, Systems and Structures, 35(3):306– 321, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000225.

Berry:1980:TMV

[BKYV80] D. M. Berry, R. A. Kemmerer, S. Yemini, and A. Von Staa. Toward modular verifiable exception handling. Computer Languages, 5(2):77–101, ???? 1980. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bassiouni:1992:RQL

[BL92] M. A. Bassiouni and M. J.

Llewellyn. A relational-calculus query language for historical databases. Computer Languages, 17(3):185–197, July 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bassiouni:1999:ETQ

[BL99] M. A. Bassiouni and M. Llewellyn.

Extending temporal query languages to handle imprecise time intervals. Computer Languages, 25(1):39– 54, April 1, 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/17/18/article.
pdf.

Blieberger:1994:DLW

[Bli94] Johann Blieberger. Discrete loops and worst case performance. Computer Languages, 20(3):193–212, August 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bassiouni:1993:TOR

[BLM93] M. A. Bassiouni, M. J. Llewellyn, and A. Mukherjee. Time-based operators for relational algebra query languages. Computer Languages, 19(4): 261–276, October 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bryant:1995:GGF

[BM95] Kelvin S. Bryant and Jon Mauney. GRAS: a general framework for combining automatic code generation and register allocation. Computer Languages, 21(2):101–112, July 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Buga:2018:EBB

[BMN^{+}18] Andreea Buga, Atif Mashkoor, Sorana Tania Nemes, Klaus-Dieter Schewe, and Pornpan Songprasop. An Event-B-based approach to hybrid systems engineering and its application to a hemodialysis machine case study. Computer Languages, Systems and Structures, 54(??):297–315, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S147784241830037X.

Boudriga:1992:RMS

[BMZM92] N. Boudriga, A. Mili, R. Zalila, and Fatma Mili. A relational model for the specification of data types. Computer Languages, 17(2):101–131, April 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bouraqadi:2004:SMC

[Bou04] Noury Bouraqadi. Safe metaclass composition using mixinbased inheritance. Computer Languages, Systems and Structures, 30(1–2):49–61, April/ July 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Bourbakis:2008:GFL

[Bou08] Nikolaos G. Bourbakis. A generic, formal language-based methodology for hierarchical floorplanning-placement. Computer Languages, Systems and Structures, 34(1): 25–42, April 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000285.

Barthe:2007:STP

[BRB07] Gilles Barthe, Tamara Rezk, and Amitabh Basu. Security types preserving compilation. Computer Languages, Systems and Structures, 33(2):

35–59, July 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842405000230.

Breitner:2018:CA

[Bre18] Joachim Breitner. Call arity. Computer Languages, Systems and Structures, 52(??): 65–91, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416300756.

Broy:1988:BBC

[Bro88]  Manfred Broy.  Broadcast-

ing buffering communication. Computer Languages, 13(1): 31–47, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Biskup:1990:ESQ

[BRS90] Joachim Biskup, Uwe Räsch, and Holger Stiefeling. An extension of SQL for querying graph relations. Computer Languages, 15(2):65–82, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Brogi:1999:DCP

[BRT99] Antonio Brogi, Chiara Renso, and Franco Turini. Dynamic composition of parameterised logic modules. Computer Languages, 25(4):211–242, December 1999. CODEN COLADA.

ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/28/30/25/
abstract.html.

Bryant:2015:SIP

[Bry15] Barrett R. Bryant. Special issue on the programming languages track at the 29th ACM Symposium on Applied Computing. Computer Languages, Systems and Structures, 42 (??):1, ???? 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000251.

Bryant:2016:SIP

[Bry16] Barrett R. Bryant. Special issue on the programming languages track at the 30th ACM Symposium on Applied Computing. Computer Languages, Systems and Structures, 45(??): 163, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300598.

Bachmann:1978:SCL

[BS78] Wolfgang Bachmann and Herward Schwarze. Solmar — a computer language for utilizing sonar models. Computer Languages, 3(3):193–201, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Budd:1985:PTS

[BS85] Timothy A. Budd and Gopal Ajei S. Program testing by specification mutation. Computer Languages, 10(1):63–73, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Barnard:1992:CPP

[BS92] D. T. Barnard and D. B. Skillicorn. Context-free parsing on O(n) processors. Computer Languages, 17(1):61–66, ???? 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Burton:2018:OMD

[BS18] Eden Burton and Emil Sekerinski. An object model for dynamic mixins. Computer Languages, Systems and Structures, 51(??):90–101, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417300490.

Besova:2015:GBM

[BSW15] Galina Besova, Dominik Steenken, and Heike Wehrheim. Grammarbased model transformations: Definition, execution, and quality properties. Computer Languages, Systems and Structures, 43(??):116–138, October 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000287.

Bal:1986:LMG

[BT86] Henri E. Bal and Andrew S.

Tanenbaum. Language-and machine-independent global optimization on intermediate code. Computer Languages, 11 (2):105–121, ???? 1986. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bal:1991:DPS

[BT91] Henri E. Bal and Andrew S.

Tanenbaum. Distributed programming with shared data. Computer Languages, 16(2): 129–146, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Budd:1982:IGC

[Bud82] T. A. Budd. An implementation of generators in C. Computer Languages, 7(2):69–87, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Buffenbarger:2017:ACB

[Buf17] Jim Buffenbarger. Amake:

cached builds of top-level targets. Computer Languages, Systems and Structures, 50(??):20–30, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416301622.

Burns:1990:NPR

[BW90] A. Burns and A. J. Wellings. The notion of priority in realtime programming languages. Computer Languages, 15(3): 153–162, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bouraqadi:2005:I

[BW05] Noury Bouraqadi and Roel Wuyts. Introduction. Computer Languages, Systems and Structures, 31(3–4):103–105, October/December 2005. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Bail:1988:PCU

[BZ88] William G. Bail and Marvin V. Zelkowitz. Program complexity using hierarchical abstract computers. Computer Languages, 13(3-4):109–123, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chaitin:1981:RAC

[CAC^{+}81] Gregory J. Chaitin, Marc A. Auslander, Ashok K. Chandra, John Cocke, Martin E. Hopkins, and Peter W. Markstein. Register allocation via coloring. Computer Languages, 6 (1):47–57, ???? 1981. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Carvalho:1978:PGS

[Car78] Sergio E. R. Carvalho. Proposal for a generalized “For” statement. Computer Languages, 3(3):157–162, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Canedo:2008:NCG

[CAS08] Arquimedes Canedo, Ben Abderazek, and Masahiro Sowa. A new code generation algorithm for 2-offset producer order queue computation model. Computer Languages, Systems and Structures, 34(4):184–194, December 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000188.

Clarke:1993:EHP

[CB93a] Gwen Clarke and David T.

Barnard. Error handling in a parallel LR substring parser. Computer Languages, 19(4): 247–259, October 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cordy:1993:ISI

[CB93b] James R. Cordy and Mario Barbacci. Introduction to the special issue on ICCL ’92. Computer Languages, 19(2):63–64, April 1993. CODEN COLADA.

ISSN 0096-0551 (print), 1873-6742 (electronic).

Chadha:2017:FDC

[CBTR17] Sanchit Chadha, Antuan Byalik, Eli Tilevich, and Alla Rozovskaya. Facilitating the development of crossplatform software via automated code synthesis from web-based programming resources. Computer Languages, Systems and Structures, 48(??): 3–19, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300634.

Chebba:2018:AAR

[CBTR18] Asmaa Chebba, Thouraya Bouabana-Tebibel, and Stuart H. Rubin. Attributed and n-ary relations in OWL for knowledge modeling. Computer Languages, Systems and Structures, 54(??):183–198, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417302026.

Citrin:1995:CRL

[CC95] Wayne Citrin and Alistair Cockburn. Carla: a rule language for specifying communications architectures. Computer Languages, 21(3-4):165– 189, October–December 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Choi:2015:LAC

[CC15] Kwanghoon Choi and Byeong-Mo Chang. A lightweight approach to component-level exception mechanism for robust Android apps. Computer Languages, Systems and Structures, 44 (Part C)(??):283– 298, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300038.

Choi:2016:LAC

[CC16] Kwanghoon Choi and Byeong-Mo Chang. A lightweight approach to component-level exception mechanism for robust Android apps. Computer Languages, Systems and Structures, 45(??):283– 298, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300038.

Cadavid:2015:AMP

[CCB15] Juan José Cadavid, Benoit Combemale, and Benoit Baudry. An analysis of metamodeling practices for MOF and OCL. Computer Languages, Systems and Structures, 41(??): 42–65, April 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000068.

Cortesi:2015:ADT

[CCF15] Agostino Cortesi, Giulia Costantini, and Pietro Ferrara. The abstract domain of Trapezoid Step Functions. Computer Languages, Systems and Structures, 43(??):41–68, October 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S147784241500024X.

Chang:2012:CRC

[CCGC12] J. Morris Chang, Wei-Mei Chen, Paul A. Griffin, and Ho-Yuan Cheng. Cyclic reference counting by typed reference fields. Computer Languages, Systems and Structures, 38(1): 98–107, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842411000285.

Ching:1993:PBS

[CCJ93] Wai-Mee Ching, Paul Carini, and Dz-Ching Ju. A primitivebased strategy for producing efficient code for very high level programs. Computer Languages, 19(1):41–50, January 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cazzola:2018:DML

[CCRS18] Walter Cazzola, Ruzanna Chitchyan, Awais Rashid, and Albert Shaqiri. μ-DSU: a micro-language based approach to dynamic software updating. Computer Languages, Systems and Structures, 51(??):71– 89, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300416.

Cleereman:2008:MIC

[CCT08] Kevin Cleereman, Michelle Cheatham, and Krishnaprasad Thirunarayan. Mechanisms for improved covariant typechecking. Computer Languages, Systems and Structures, 34(1): 1–17, April 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000133.

Celentano:1981:ITP

[CD81] Augusto Celentano and Pierluigi Della Vigna. Interprocedural test path evaluation. Computer Languages, 6(3-4): 115–130, ???? 1981. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cocco:1982:MEH

[CD82] N. Cocco and S. Dulli. A mechanism for exception handling and its verification rules. Computer Languages, 7(2):89–102, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Costagliola:2018:ULC

[CDF18] Gennaro Costagliola, Mattia De Rosa, and Vittorio Fuccella. Using the local context for the definition and implementation of visual languages. Computer Languages, Systems and Structures, 54(??):20–38, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300307.

Celentano:1980:SPD

[CDGM80] Augusto Celentano, Pierluigi Della Vigna, Carlo Ghezzi, and Dino Mandrioli. Simple: a program development system. Computer Languages, 5 (2):103–114, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chis:2015:PDS

[CDGN15] Andrei Chis, Marcus Denker, Tudor Gîrba, and Oscar Nierstrasz. Practical domainspecific debuggers using the Moldable Debugger framework. Computer Languages, Systems and Structures, 44 (part A)(??):89–113, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000561.

Cassou:2009:TWD

[CDW09] Damien Cassou, Stéphane Ducasse, and Roel Wuyts. Traits at work: The design of a new trait-based stream library. Computer Languages, Systems and Structures, 35(1): 2–20, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000213.

Celentano:1981:LPT

[Cel81] Augusto Celentano. LR parsing technique for extended contextfree grammars. Computer Languages, 6(2):95–107, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chou:1979:ITN

[CF79] S. M. Chou and K. S. Fu. Inference for transition network grammars. Computer Languages, 4(2):83–92, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Casanova:1988:SPP

[CF88] Marco A. Casanova and Antonio L. Furtado. String patternmatching in Prolog. Computer Languages, 13(3-4):149– 170, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cortesi:2002:CLS

[CF02] Agostino Cortesi and Riccardo Focardi. Computer languages and security. Computer Languages, Systems and Structures, 28(1):1–2, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Ciancarini:2000:DCL

[CFG00] Paolo Ciancarini, Daniela Fogli, and Mauro Gaspari. A declarative coordination language. Computer Languages, 26(2–4):125–163, July 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/15/18/50/31/29/
                                        abstract.html.

Colombetti:1984:SCD

[CG84] Marco Colombetti and Giovanni Guida. Supporting control definition in programming languages. Computer Languages, 9(1):1–23, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Clematis:1993:SCO

[CG93] Andrea Clematis and Vittoria Gianuzzi. Structuring conversation in operation/procedure oriented programming languages. Computer Languages, 18(3):153–168, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ciancarini:1996:RCL

[CG96] Paolo Ciancarini and Mauro Gaspari. Rule-based coordination of logic programs. Computer Languages, 22(2–3):95– 113, July–October 1996. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chan:2009:GOC

[CGG^{+}09] Hiu Ning (Angela) Chan, Andrew J. Gallagher, Appu S. Goundan, Yi Lin William Au Yeung, Aaron W. Keen, and Ronald A. Olsson. Generic operations and capabilities in the JR concurrent programming language. Computer Languages, Systems and Structures, 35(3):293–305, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000183.

Clark:2002:IFA

[CHH02] David Clark, Chris Hankin, and Sebastian Hunt. Information flow for Algol-like languages. Computer Languages, Systems and Structures, 28(1): 3–28, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Cordy:1991:TRP

[CHHP91] James R. Cordy, Charles D.

Halpern-Hamu, and Eric Promislow. TXL: a rapid prototyping system for programming language dialects. Computer Languages, 16(1):97–107, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cooper:1993:MPC

[CHK93] Keith D. Cooper, Mary W. Hall, and Ken Kennedy. A methodology for procedure cloning. Computer Languages, 19(2):105–117, April 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Charousset:2016:RAP

[CHS16] Dominik Charousset, Raphael Hiesgen, and Thomas C. Schmidt. Revisiting actor programming in C++. Computer Languages, Systems and Structures, 45(??):105– 131, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416000038.

Ciancarini:1992:PPL

[Cia92] Paolo Ciancarini. Parallel programming with logic languages: a survey. Computer Languages, 17(4):213–239, October 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cater:1984:CLE

[CIF84] Steven C. Cater, S. Sitharama Iyengar, and John Fuller. Computation of logical effort in high level languages. Computer Languages, 9(3-4):133– 148, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cadoli:2000:NSE

[CIP^{+}00] Marco Cadoli, Giovambattista Ianni, Luigi Palopoli, Andrea Schaerf, and Domenico Vasile. NP-SPEC: an executable specification language for solving all problems in NP. Computer Languages, 26(2–4):165–195, July 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/50/31/30/
abstract.html.

Crawford:1980:NAC

[CJ80] John H. Crawford and Mehdi Jazayeri. New approach to code motion and its application to hoisting. Computer Languages, 5(1):29–36, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chavarriaga:2017:ABX

[CJD17] Enrique Chavarriaga, Francisco Jurado, and Fernando Díez. An approach to build XML-based domain specific languages solutions for clientside web applications. Computer Languages, Systems and Structures, 49(??):133–151, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301634.

Childers:2008:P

[CK08] Bruce Childers and Mahmut Kandemir. Preface. Computer Languages, Systems and Structures, 34(4):151–152, December 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000322.

Combemale:2018:COL

[CKM^{+}18] Benoit Combemale, Jörg Kienzle, Gunter Mussbacher, Olivier Barais, Erwan Bousse, Walter Cazzola, Philippe Collet, Thomas Degueule, Robert Heinrich, Jean-Marc Jézéquel, Manuel Leduc, Tanja Mayerhofer, Sébastien Mosser, Matthias Schöttle, Misha Strittmatter, and Andreas Wortmann. Concern-oriented language development (COLD): Fostering reuse in language engineering. Computer Languages, Systems and Structures, 54(??):139–155, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300496.

Coon:1983:CCI

[CKS83] Lawrence A. Coon, John P. Kearns, and Mary Lou Soffa. The contraction of control implementations. Computer Languages, 8(1):15–25, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Choen:1978:SFM

[CL78] A. Toni Choen and Leon S.

Levy. Structured flowcharts for multiprocessing. Computer Languages, 4(4):209–226, 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Coleman:1989:PIC

[CL89] Don M. Coleman and Ronald J.

Leach. Performance issues in C language fault-tolerant software. Computer Languages, 14(1):1–9, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chakravarty:1997:TUI

[CL97] Manuel M. T. Chakravarty and Hendrik C. R. Lock. Towards the uniform implementation of declarative languages. Computer Languages, 23(2-4):121– 160, July–December 1997. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cunha:1983:MOP

[CLM83] P. R. F. Cunha, C. J. Lucena, and T. S. E. Maibaum. Message oriented programming — a resource based methodology. Computer Languages, 8 (3-4):95–111, ???? 1983. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ciampolini:2001:LLC

[CLMT01] Anna Ciampolini, Evelina Lamma, Paola Mello, and Paolo Torroni. LAILA: a language for coordinating abductive reasoning among logic agents. Computer Languages, 27(4):137–161, December 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/54/31/27/
abstract.html.

Ciampolini:1996:DLO

[CLSM96] A. Ciampolini, E. Lamma, C. Stefanelli, and P. Mello. Distributed logic objects. Computer Languages, 22(4):237– 258, December 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chandra:1975:PPF

[CM75] A. K. Chandra and Z. Manna. On the power of programming features. Computer Languages, 1(3):219–232, September 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cockshott:2006:OPP

[CM06] Paul Cockshott and Greg Michaelson. Orthogonal parallel processing in vector Pascal. Computer Languages, Systems and Structures, 32(1):2– 41, April 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Cabral:2011:TMA

[CM11] Bruno Cabral and Paulo Marques. A transactional model for automatic exception handling. Computer Languages, Systems and Structures, 37(1): 43–61, April 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000266.

Cocco:1985:ATS

[CMM85] N. Cocco, D. Mandrioli, and V. Milanese. Ada task system and real-time applications: an implementation schema. Computer Languages, 10(3-4):189– 209, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Choe:2009:QGR

[CNGW09] Guija Choe, Young-Kwang Nam, Joseph Goguen, and Guilian Wang. Query generation for retrieving data from distributed semistructured documents using a metadata interface. Computer Languages, Systems and Structures, 35(4):422–434, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000407.

Coffin:1989:SAM

[CO89] Michael Coffin and Ronald A. Olsson. An SR approach to multiway rendezvous. Computer Languages, 14(4):255– 262, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chung:1998:NMI

[CO98] M. Chung and R. A. Olsson. New mechanisms for invocation handling in concurrent programming languages. Computer Languages, 24(4):245–270, December 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/15/18/27/20/
                                        19/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/15/18/27/20/19/article.
                                        pdf.

Cohen:1978:SFM

[Coh78]  A. T. Cohen.  Structured

flowcharts for multiprocessing. Computer Languages, 3 (4):209–226, ???? 1978. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Crawford:1995:SID

[COHW95] Richard H. Crawford, Ronald A. Olsson, W. Wilson Ho, and Christopher E. Wee. Semantic issues in the design of languages for debugging. Computer Languages, 21(1):17–37, April 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cook:1981:ADP

[Coo81] Robert P. Cook. Abstractions for distributed programming. Computer Languages, 6 (3-4):131–138, ???? 1981. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Cooke:1998:SPD

[Coo98] Daniel E. Cooke. SequenceL provides a different way to view programming. Computer Languages, 24(1):1–32, April 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/17/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/17/17/article.
pdf.

Coven:1993:AAP

[Cov93] H. Justin Coven. Altering and applying predicates. Computer Languages, 18(1):57–75, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Coen-Porisini:1993:ARS

[CPD93] Alberto Coen-Porisini and Flavio De Paoli. Array representation in symbolic execution. Computer Languages, 18 (3):197–216, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Combemale:2015:SII

[CPPV15] Benoit Combemale, David J. Pearce, Richard F. Paige, and Eric Van Wyk. Special issue on the 6th and 7th International Conferences on Software Language Engineering (SLE 2013 and SLE 2014). Computer Languages, Systems and Structures, 44 (part A)(??):1– 2, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000597.

Crowley:1979:PDP

[Cro79]  Charles Crowley.  Parallel

developments in programming languages and operating systems. Computer Languages, 4 (2):71–82, ???? 1979. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Crespi-Reghizzi:2000:ADP

[CRPP00] Stefano Crespi-Reghizzi, Matteo Pradella, and Pierluigi San Pietro. Associative definition of programming languages. Computer Languages, 26(2–4):105–123, July 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/50/31/28/
abstract.html.

Campos:2003:DRC

[CS03] Alvaro E. Campos and Dionel A.

Suazo. Data-race and concurrentwrite freedom are undecidable. Computer Languages, Systems and Structures, 29(1–2):1–13, April/July 2003. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Capelli:2016:FED

[CS16] Steven Capelli and Patrizia Scandurra. A framework for early design and prototyping of service-oriented applications with design patterns. Computer Languages, Systems and Structures, 46(??):140– 166, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300440.

Cordoba-Sanchez:2016:ADS

[CSdL16] Irene Córdoba-Sánchez and Juan de Lara. Ann: a domain-specific language for the effective design and validation of Java annotations. Computer Languages, Systems and Structures, 45(??):164– 190, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300318.

Cazzola:2014:JBR

[CV14] Walter Cazzola and Edoardo Vacchi. @Java: Bringing a richer annotation model to Java. Computer Languages, Systems and Structures, 40(1): 2–18, April 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842414000037.

Cazzola:2016:LCM

[CV16] Walter Cazzola and Edoardo Vacchi. Language components for modular DSLs using traits. Computer Languages, Systems and Structures, 45(??): 16–34, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300208.

Chan:2002:AGF

[CY02] Jien-Tsai Chan and Wuu Yang. An attribute-grammar framework for specifying the accessibility in Java programs. Computer Languages, Systems and Structures, 28(2):203–235, December 15, 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Chen:2015:AVB

[CYS^{+}15] Jiunn-Yeu Chen, Wuu Yang, Bor-Yeh Shen, Yuan-Jia Li, and Wei-Chung Hsu. Automatic validation for binary translation. Computer Languages, Systems and Structures, 43(??):96–115, October 2015. CODEN ????

ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000275.

Cortesi:2011:WNO

[CZ11] Agostino Cortesi and Matteo Zanioli. Widening and narrowing operators for abstract interpretation. Computer Languages, Systems and Structures, 37(1):24–42, April 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000254.

Costa:2014:JTV

[dACSAP14] Igor Rafael de Assis Costa, Henrique Nazaré Santos, Péricles Rafael Alves, and Fernando Magno Quintão Pereira. Just-in-time value specialization. Computer Languages, Systems and Structures, 40(2):37–52, ???? 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842413000286.

Dain:1994:PMD

[Dai94] J. A. Dain. A practical minimum distance method for syntax error handling. Computer Languages, 20(4):239– 252, November 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Dimitrieski:2015:CEE

[DCA^{+}15] Vladimir Dimitrieski, Milan Celiković, Slavica Aleksić, Sonja Ristić, Abdalla Alargt, and Ivan Luković. Concepts and evaluation of the extended entity-relationship approach to database design in a multiparadigm information system modeling tool. Computer Languages, Systems and Structures, 44 (Part C)(??):299– 318, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000615.

Dimitrieski:2016:CEE

[DCA^{+}16] Vladimir Dimitrieski, Milan Celiković, Slavica Aleksić, Sonja Ristić, Abdalla Alargt, and Ivan Luković. Concepts and evaluation of the extended entity-relationship approach to database design in a multi-paradigm information system modeling tool. Computer Languages, Systems and Structures, 45(??):299– 318, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000615.

Degueule:2017:SMP

[DCB^{+}17] Thomas Degueule, Benoit Combemale, Arnaud Blouin, Olivier Barais, and Jean-Marc Jézéquel. Safe model polymorphism for flexible modeling. Computer Languages, Systems and Structures, 49(??):176–195, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300264.

DiRuscio:2017:SIF

[DdLP17] Davide Di Ruscio, Juan de Lara, and Alfonso Pierantonio. Special issue on flexible model driven engineering. Computer Languages, Systems and Structures, 49(??):174–175, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416302068.

Denker:2006:RBT

[DDT06] Marcus Denker, Stéphane Ducasse, and Éric Tanter. Runtime bytecode transformation for Smalltalk. Computer Languages, Systems and Structures, 32(2–3):125–139, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000436.

DeMeuter:2008:P

[De 08] Wolfgang De Meuter. Preface.

Computer Languages, Systems and Structures, 34(2–3):45, July/October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842407000310.

Demers:1975:ESP

[Dem75] A. J. Demers. Elimination of single productions and merging nonterminal symbols of LR(1) grammars. Computer Languages, 1(2):105–119, June 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Drew:1994:EHE

[DG94] Steven J. Drew and K. John Gough. Exception handling: expecting the unexpected. Computer Languages, 20(2): 69–87, May 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Diaz-Gonzalez:1991:LAE

[DGU91] Jose P. Diaz-Gonzalez and Joseph E. Urban. Language aspects of ENVISAGER. an object-oriented environment for the specification of real-time systems. Computer Languages, 16(1):19–37, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Donnan:1986:PSR

[DH86] G. Donnan and M. Elizabeth C. Hull. On processes, synchronization and redundant code in communicating sequential processes. Computer Languages, 11(3-4):155–160, ???? 1986.

CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Dybvig:1989:EC

[DH89] R. Kent Dybvig and Robert Hieb. Engines from continuations. Computer Languages, 14 (2):109–123, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Dhamdhere:1988:RAU

[Dha88] D. M. Dhamdhere. Register assignment using code placement techniques. Computer Languages, 13(2):75–93, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Dhamdhere:1990:ULA

[Dha90] D. M. Dhamdhere. A usually linear algorithm for register assignment using edge placement of load and store instructions. Computer Languages, 15 (2):83–94, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Djakovic:1988:RLO

[Dja88] Dragan D. Djakovic. RASP — a language with operations on fuzzy set. Computer Languages, 13(3-4):143–147, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Dhamdhere:1983:CPL

[DK83] D. M. Dhamdhere and J. S. Keith. Characterization of program loops in code optimization. Computer Languages, 8 (2):69–76, ???? 1983. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

DaRosdeCarvalho:1992:OAV

[DK92] Claudio Sergio Da Ros de Carvalho and Tomasz Kowaltowski. On open arrays and variable number of parameters. Computer Languages, 17 (1):67–74, ???? 1992. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Dernehl:2018:CSA

[DKK18] Christian Dernehl, Jan Kühn, and Stefan Kowalewski. Case studies on automated verification with slope boundaries for block diagrams. Computer Languages, Systems and Structures, 54(??):528–543, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300246.

DSouza:2017:SII

[DL17] Deepak D’Souza and Akash Lal. Special issue on the 16th International Conference on Verification, Model Checking, and Abstract Interpretation. Computer Languages, Systems and Structures, 47 (part 1) (??):1, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300811.

DiCosmo:2007:CPC

[DLP07] Roberto Di Cosmo, Zheng Li, and Susanna Pelagatti. A calculus for parallel computations over multidimensional dense arrays. Computer Languages, Systems and Structures, 33(3–4):82–110, October/December 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000303.

Dabrowski:2015:FSN

[DLP15] Frédéric Dabrowski, Frédéric Loulergue, and Thomas Pinsard. A formal semantics of nested atomic sections with thread escape. Computer Languages, Systems and Structures, 42(??): 2–21, ???? 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000238.

delaVega:2018:FDF

[dlVGSZS18] Alfonso de la Vega, Diego García-Saiz, Marta Zorrilla, and Pablo Sánchez. FLANDM: a development framework of domain-specific languages for data mining democratisation. Computer Languages, Systems and Structures, 54(??):316– 336, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842418300381.

deLamadrid:2012:CFH

[dLZ12] James Gil de Lamadrid and Jill Zimmerman. Core FOBS: a hybrid functional and objectoriented language. Computer Languages, Systems and Structures, 38(3):214–241, October 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842412000127.

Deng:2010:CDA

[DMT10] Xiao Yan Deng, Greg Michaelson, and Phil Trinder. Costdriven autonomous mobility. Computer Languages, Systems and Structures, 36(1): 34–59, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842409000049.

Koster:2016:DSS

[DMVD16] Joeri De Koster, Stefan Marr, Tom Van Cutsem, and Theo D’Hondt. Domains: Sharing state in the communicating event-loop actor model. Computer Languages, Systems and Structures, 45(??):132– 160, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241600004X.

Dan:2017:EAV

[DMVY17] Andrei Dan, Yuri Meshman, Martin Vechev, and Eran Yahav. Effective abstractions for verification under relaxed memory models. Computer Languages, Systems and Structures, 47 (part 1)(??):62– 76, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300331.

DiSanto:1990:PCA

[DNR90] Michele Di Santo, Libero Nigro, and Wilma Russo. Programmer-defined control abstractions in Modula-2. Computer Languages, 15(3): 141–152, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Guimaraes:2006:GL

[dOG06] José de Oliveira Guimarães.

The Green language. Computer Languages, Systems and Structures, 32(4):203–215, December 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000369.

Guimaraes:2009:GLT

[dOG09] José de Oliveira Guimarães. The Green language type system. Computer Languages, Systems and Structures, 35(4):435–447, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842408000419.

Demaid:2006:AEO

[DOZ06] A. Demaid, S. Ogden, and J. Zucker. Access Enhancement Objects for data management in Smalltalk. Computer Languages, Systems and Structures, 32(4):185–202, December 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842405000217.

Degano:1998:LTH

[DP98] Pierpaolo Degano and Corrado Priami. LR techniques for handling syntax errors. Computer Languages, 24(2): 73–98, July 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/15/18/27/18/
                                        18/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/15/18/27/18/18/article.
                                        pdf.

Demeyer:2009:GEI

[DP09] Serge Demeyer and Jean-François Perrot. Guest Editor introduction. Computer Languages, Systems and Structures, 35(1):1, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842408000195.

Dimopoulos:2010:PAG

[DPP10] Alexandros C. Dimopoulos, Christos Pavlatos, and George Papakonstantinou. A platform for the automatic generation of attribute evaluation hardware systems. Computer Languages, Systems and Structures, 36(2): 203–222, July 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000359.

Dovier:2009:AMC

[DQ09] A. Dovier and E. Quintarelli.

Applying model-checking to solve queries on semistructured data. Computer Languages, Systems and Structures, 35(2): 143–172, July 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000339.

Drew:1996:FTT

[Dre96] S. J. Drew. Fail-safety techniques and their extensions to concurrent systems. Computer Languages, 22(4):193– 203, December 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Diaz:1997:DDR

[DRT97] M. Diaz, B. Rubio, and J. M. Troya. DRL: a distributed real-time logic language. Computer Languages, 23(2-4):87– 120, July–December 1997. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Doh:1993:ASD

[DS93] Kyung-Goo Doh and David A. Schmidt. Action semanticsdirected prototyping. Computer Languages, 19(4):213– 233, October 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ducasse:2005:USM

[DSW05] Stéphane Ducasse, Nathanael Schärli, and Roel Wuyts. Uniform and safe metaclass composition. Computer Languages, Systems and Structures, 31(3– 4):143–164, October/December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Dandan:2013:TSR

[DTXP13] Gong Dandan, Wang Tiantian, Su Xiaohong, and Ma Peijun. A test-suite reduction approach to improving fault-localization effectiveness.

Computer Languages, Systems and Structures, 39(3):95– 108, October 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842413000134.

Ducasse:2004:E

[DW04] Stephane Ducasse and Roel Wuyts. Editorial. Computer Languages, Systems and Structures, 30(1–2):1–3, April/July 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Earley:1975:HLI

[Ear75] J. Earley. High level iterators and a method for automatically designing data structure representation. Computer Languages, 1(4):321–342, ???? 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

English:2012:CSC

[ECB12] Michael English, Tony Cahill, and Jim Buckley. Construct specific coupling measurement for C++ software. Computer Languages, Systems and Structures, 38(4):300–319, December 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842412000243.

Ernst:1991:MVA

[EHMO91] George W. Ernst, Raymond J. Hookway, James A. Menegay, and William F. Ogden. Modular verification of Ada generics. Computer Languages, 16 (3-4):259–280, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Eckart:1987:OAL

[EL87] J. D. Eckart and R. J. LeBlanc. Overloading in the Ada language: is it too restrictive? Computer Languages, 12(3-4): 163–171, ???? 1987. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ellmenreich:2007:CSR

[EL07] Nils Ellmenreich and Christian Lengauer. Costing stepwise refinements of parallel programs. Computer Languages, Systems and Structures, 33(3–4):134–167, October/December 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842406000145.

Edelson:1989:CSC

[EP89] Daniel Edelson and Ira Phol. C++: solving C’s shortcomings? Computer Languages, 14 (3):137–152, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ergin:2016:DPO

[ESG16] Huseyin Ergin, Eugene Syriani, and Jeff Gray. Design pattern oriented development of model transformations. Computer Languages, Systems and Structures, 46(??):106– 139, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300148.

Erdweg:2015:ECL

[EvdSV^{+}15] Sebastian Erdweg, Tijs van der Storm, Markus Völter, Laurence Tratt, Remi Bosman, William R. Cook, Albert Gerritsen, Angelo Hulshout, Steven Kelly, Alex Loh, Gabriël Konat, Pedro J. Molina, Martin Palatnik, Risto Pohjonen, Eugen Schindler, Klemens Schindler, Riccardo Solmi, Vlad Vergu, Eelco Visser, Kevin van der Vlist, Guido Wachsmuth, and Jimi van der Woning. Evaluating and comparing language workbenches: Existing results and benchmarks for the future. Computer Languages, Systems and Structures, 44 (part A)(??):24– 47, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000573.

Font:2017:LVM

[FAHC17] Jaime Font, Lorena Arcega, Øystein Haugen, and Carlos Cetina. Leveraging variability modeling to address metamodel revisions in Modelbased Software Product Lines. Computer Languages, Systems and Structures, 48(??): 20–38, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S147784241630001X.

Falkman:1997:PSD

[Fal97] G. Falkman. Program separation and definitional higher order programming. Computer Languages, 23(2-4):179– 206, July–December 1997. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fabresse:2012:LBG

[FBDH12] Luc Fabresse, Noury Bouraqadi, Christophe Dony, and Marianne Huchard. A language to bridge the gap between component-based design and implementation. Computer Languages, Systems and Structures, 38(1):29–43, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842411000339.

Faddegon:2018:TGO

[FC18] Maarten Faddegon and Olaf Chitil. Type generic observation of intermediate data structures for debugging lazy functional programs. Computer Languages, Systems and Structures, 52(??):92– 110, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300732.

Fabresse:2008:FSU

[FDH08] Luc Fabresse, Christophe Dony, and Marianne Huchard. Foundations of a simple and unified component-oriented language. Computer Languages, Systems and Structures, 34(2–3):130–149, July/ October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000164.

Felleisen:1987:RLJ

[Fel87] M. Felleisen. Reflections on Landin’s J-operator: a partly historical note. Computer Languages, 12(3-4):197–207, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Flores:1975:SSL

[FF75] I. Flores and M. Feuerman.

Source statement libraries and IBM System/370. Computer Languages, 1(2):139–150, June 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Felleisen:1986:CLE

[FF86] Matthias Felleisen and Daniel P. Friedman. A closer look at export and import statements. Computer Languages, 11(1): 29–37, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Franco:1989:CEP

[FF89] John Franco and Daniel P. Friedman. Creating efficient programs by exchanging data for procedures. Computer Languages, 14(1):11–23, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Franco:1990:TFL

[FF90] John Franco and Daniel P. Friedman. Towards a facility for lexically scoped, dynamic mutual recursion in Scheme. Computer Languages, 15(1): 55–64, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Franco:1990:MSS

[FFJ90] John Franco, Daniel P. Friedman, and Steven D. Johnson. Multi-way streams in Scheme. Computer Languages, 15(2):109–125, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fister:2011:DID

[FFMB11] Iztok Fister, Jr., Iztok Fister, Marjan Mernik, and Janez Brest. Design and implementation of domain-specific language easytime. Computer Languages, Systems and Structures, 37(4):151–167, October 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000145.

Fisher:1988:COC

[Fis88] A. J. Fisher. A critique of Occam channel types. Computer Languages, 13(2):95–105, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Flores:1975:LEF

[FKR75] I. Flores, P. Kaminsky, and D. Ryan. List and execute forms of macros. Computer Languages, 1(1):45–60, January 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Feeley:1987:UCC

[FL87] Marc Feeley and Guy Lapalme.

Using closures for code generation. Computer Languages, 12 (1):47–66, ???? 1987. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Feeley:1992:CGB

[FL92] Marc Feeley and Guy Lapalme.

Closure generation based on viewing lambda as epsilon plus compile. Computer Languages, 17(4):251–267, October 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Franzle:2001:VTL

[FL01] Martin Fränzle and Karsten Lüth. Visual temporal logic as a rapid prototyping tool. Computer Languages, 27(1–3):93–113, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/54/27/32/
abstract.html.

Fleck:1978:RDT

[Fle78] A. C. Fleck. Recent developments in the theory of data structures. Computer Languages, 3(1):37–52, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fleck:1984:PCT

[Fle84] A. C. Fleck. Proposal for comparison of types in Pascal and associated semantic models. Computer Languages, 9 (2):71–87, ???? 1984. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fleck:1986:SFF

[Fle86]  A. C. Fleck.  Structuring

Fp-Style functional programs. Computer Languages, 11(2): 55–63, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Flores:1978:MME

[Flo78] Ivan Flores. Making macrospace effective. Computer Languages, 3(2):95–113, ???? 1978. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fischer:1980:REP

[FM80] C. N. Fischer and J. Mauney.

On the role of error productions in syntactic error correction. Computer Languages, 5 (3-4):131–139, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fabry:2004:LID

[FM04] Johan Fabry and Tom Mens.

Language-independent detection of object-oriented design patterns. Computer Languages, Systems and Structures, 30(1– 2):21–33, April/July 2004. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Farhad:2016:MSP

[FNRR16] S. M. Farhad, Muhammad Ali Nayeem, Md. Khaledur Rahman, and M. Sohel Rahman. Mapping stream programs onto multicore platforms by local search and genetic algorithm. Computer Languages, Systems and Structures, 46(??):182– 205, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241530052X.

Fodor:2002:SIP

[FO02] Eugene F. Fodor and Ronald A. Olsson. SIR: inter-program concurrency support for SR programs. Computer Languages, Systems and Structures, 28(4):307–325, December 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Fong:2010:IUS

[FO10] Philip W. L. Fong and Simon Orr. Isolating untrusted software extensions by custom scoping rules. Computer Languages, Systems and Structures, 36(3):268–287, October 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S147784240900044X.

Friedman:1992:BBB

[Fri92] Linda Weiser Friedman. From Babbage to Babel and beyond: a brief history of programming languages. Computer Languages, 17(1):1–17, ???? 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Fruja:2010:TPT

[Fru10]  Nicu G. Fruja.  Towards

proving type safety of C#. Computer Languages, Systems and Structures, 36(1): 60–95, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000128.

Fordo:2015:SCP

[FT15] Viktória Fördős and Melinda Tóth. Supporting comprehensible presentation of clone candidates through two-dimensional maximisation. Computer Languages, Systems and Structures, 44 (Part C)(??):355– 365, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000731.

Fordos:2016:SCP

[FT16] Viktória Fördős and Melinda Tóth. Supporting comprehensible presentation of clone candidates through two-dimensional maximisation. Computer Languages, Systems and Structures, 45(??):355–365, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842415000731.

Friedman:1978:FC

[FW78] Daniel P. Friedman and David S.

Wise. Functional combination. Computer Languages, 3 (1):31–35, ???? 1978. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ford:1987:PEM

[FW87] R. Ford and M. Wagner. Performance evaluation methodologies for abstract data type implementation techniques. Computer Languages, 12(3-4): 145–162, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Feng:1996:BLD

[FWY96] M. D. Feng, W. F. Wong, and C. K. Yuen. BaLinda lisp: design and implementation. Computer Languages, 22(4):205– 214, December 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gomez-Abajo:2017:DSL

[GAGdL17] Pablo Gómez-Abajo, Esther Guerra, and Juan de Lara. A domain-specific language for model mutation and its application to the automated generation of exercises. Computer Languages, Systems and Structures, 49(??):152–173, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S147784241630094X.

Ganapathi:1989:PBR

[Gan89a] Mahadevan Ganapathi. Prolog based retargetable code generation. Computer Languages, 14 (3):193–204, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ganapathi:1989:SPP

[Gan89b] Mahadevan Ganapathi. Semantic predicates in parser generators. Computer Languages, 14(1):25–33, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gokhale:2017:SII

[GAS17] Aniruddha Gokhale, Kenichi Asai, and Ulrik Pagh Schultz. Special issue on the 2015 International Conference on Generative Programming: Concepts & Experiences (GPCE). Computer Languages, Systems and Structures, 48(??): 1–2, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301968.

Greiner:2009:ZBS

[GBZ09] Saso Greiner, Janez Brest, and Viljem Zumer. Zero — a blend of static typing and dynamic metaprogramming. Computer Languages, Systems and Structures, 35(3):241– 251, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000146.

Gumzej:2009:SSS

[GCH09] Roman Gumzej, Matjaz Colnaric, and Wolfgang A. Halang.

Safety shell for specification-PEARL oriented UML realtime projects. Computer Languages, Systems and Structures, 35(3):277–292, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000171.

Gomez:2012:RUM

[GDD12] Verónica Uquillas Gómez, Stéphane Ducasse, and Theo D’Hondt. Ring: a unifying meta-model and infrastructure for Smalltalk source code analysis tools. Computer Languages, Systems and Structures, 38(1):44–60, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000443.

Gehani:1977:UMD

[Geh77] N. Gehani. Units of measure as a data attribute. Computer Languages, 2(3):93–111, ???? 1977. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gehani:1979:HLD

[Geh79] Narain Gehani. High level data structure — the grid. Computer Languages, 4(2):93–98, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gehani:1980:GPI

[Geh80] Narain Gehani. Generic procedures: an implementation and an undecidability result. Computer Languages, 5(3-4):155– 161, ???? 1980. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gehani:1982:CAM

[Geh82] N. H. Gehani. Concurrency in Ada and multicomputers. Computer Languages, 7(1):21– 23, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Georgeff:1981:EOP

[GFK81] M. P. Georgeff, I. Fris, and J. Kautsky. Effect of operators on parsing and evaluation in APL. Computer Languages, 6(2):67–78, ???? 1981. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gini:1982:IDO

[GG82] Giuseppina C. Gini and Maria L.

Gini. Interactive development of object handling programs. Computer Languages, 7(1):1– 10, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Guo:2009:DRA

[GG09] Hai-Feng Guo and Gopal Gupta. Dynamic reordering of alternatives for definite logic programs. Computer Languages, Systems and Structures, 35(3):252–265, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000158.

Groppe:2011:TXS

[GGK^{+}11] Sven Groppe, Jinghua Groppe, Niklas Klein, Ralf Bettentrupp, Stefan Böttcher, and Le Gruenwald. Transforming XSLT stylesheets into XQuery expressions and vice versa. Computer Languages, Systems and Structures, 37(2): 76–111, July 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000394.

Gautier:2007:RSI

[GH07] Thierry Gautier and Hamid-Reza Hamidi. Re-scheduling invocations of services for RPC grids. Computer Languages, Systems and Structures, 33(3–4):168–178, October/December 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000297.

Gogolla:2018:AMQ

[GHD18] Martin Gogolla, Frank Hilken, and Khanh-Hoang Doan. Achieving model quality through model validation, verification and exploration. Computer Languages, Systems and Structures, 54(??):474–511, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300118.

Golubski:1995:CSS

[GL95] Wolfgang Golubski and Wolfram-M. Lippe. A complete semantics for SMALLTALK-80. Computer Languages, 21(2): 67–79, July 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ghezzi:1989:SEC

[GMMP89] Carlo Ghezzi, Dino Mandrioli, Sandro Morasca, and Mauro Pezze. Symbolic execution of concurrent systems using Petri nets. Computer Languages, 14 (4):263–281, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Griswold:1988:SPL

[GO88] Ralph E. Griswold and Janalee O’Bagy. Seque: a programming language for manipulating sequences. Computer Languages, 13(1):13–22, ???? 1988. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Garg:1991:CLC

[GR91] Vijay K. Garg and C. V. Ramamoorthy. ConC. A language for concurrent programming. Computer Languages, 16(1):5–18, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Griswold:1983:ISP

[Gri83] Ralph E. Griswold. Implementing Snobol4 pattern matching in Icon. Computer Languages, 8(2):77–92, ???? 1983. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Gehlot:1986:ISA

[GS86] Vijay Gehlot and Y. N. Srikant. Interpreter for SLIPS — an applicative language based on lambda-calculus. Computer Languages, 11(1):1–13, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ghorbal:2017:HPR

[GSP17] Khalil Ghorbal, Andrew Sogokon, and André Platzer. A hierarchy of proof rules for checking positive invariance of algebraic and semialgebraic sets. Computer Languages, Systems and Structures, 47 (part 1)(??):19– 43, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000925.

Gedela:1999:CPN

[GSX99] Ravi K. Gedela, Sol M. Shatz, and Haiping Xu. Compositional Petri net models of advanced tasking in Ada-95. Computer Languages, 25(2): 55–87, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/25/article.
pdf.

Guo:2016:SAA

[Guo16]  Hai-Feng Guo.  A seman-

tic approach for automated test oracle generation. Computer Languages, Systems and Structures, 45(??):204– 219, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241530021X.

Geilen:2001:OOM

[GVvdP^{+}01] M. C. W. Geilen, J. P. M.

Voeten, P. H. A. van der Putten, L. J. van Bokhoven, and M. P. J. Stevens. Objectoriented modelling and specification using SHE. Computer Languages, 27(1–3): 19–38, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

/www.elsevier.com/gej-ng/
10/15/18/54/27/29/abstract.
html.

Gybels:2006:ILR

[GWDD06] Kris Gybels, Roel Wuyts, Stéphane Ducasse, and Maja D’Hondt. Inter-language reflection: a conceptual model and its implementation. Computer Languages, Systems and Structures, 32(2–3):109–124, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842405000448.

Gannon:1987:TIM

[GZ87] John D. Gannon and Marvin V. Zelkowitz. Two implementation models of abstract data types. Computer Languages, 12 (1):21–25, ???? 1987. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hanson:1978:DSS

[Han78]  David R. Hanson.  Data

structures in SL5. Computer Languages, 3(3):181–192, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hanus:1997:LNS

[Han97] Michael Hanus. Lazy narrowing with simplification. Computer Languages, 23(2-4):61– 85, July–December 1997. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Harland:1997:GPC

[Har97] James Harland. On goaldirected provability in classical logic. Computer Languages, 23 (2-4):161–178, July–December 1997. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hatcher:1991:ESE

[Hat91] P. J. Hatcher. The equational specification of efficient compiler code generation. Computer Languages, 16(1):81–95, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Haller:2016:SIA

[HB16] Philipp Haller and Elisa Gonzalez Boix. Special issue on the 3rd and 4th ACM SIGPLAN workshops on programming based on actors, agents, and decentralized control (AGERE! 2013/2014). Computer Languages, Systems and Structures, 45(??):161– 162, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300227.

Hermenegildo:1996:RDA

[HC96] Manuel V. Hermenegildo and Manuel Carro. Relating data-parallelism and (AND- )parallelism in logic programs. Computer Languages, 22(2–3): 143–163, July–October 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hasan:2005:SBF

[HC05] Yusuf Hasan and Morris Chang. A study of best-fit memory allocators. Computer Languages, Systems and Structures, 31(1):35–48, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Halder:2012:AID

[HC12] Raju Halder and Agostino Cortesi. Abstract interpretation of database query languages. Computer Languages, Systems and Structures, 38(2): 123–157, July 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842411000510.

Haldiman:2009:PPT

[HDN09] Niklaus Haldiman, Marcus Denker, and Oscar Nierstrasz. Practical, pluggable types for a dynamic language. Computer Languages, Systems and Structures, 35(1): 48–62, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842408000262.

Herriot:1976:SSD

[Her76] R. G. Herriot. Structured syntax diagrams. Computer Languages, 2(1-2):9–19, ???? 1976. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Haynes:1987:ATP

[HF87] Christopher T. Haynes and Daniel P. Friedman. Abstracting timed preemption with engines. Computer Languages, 12 (2):109–121, ???? 1987. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Haynes:1986:OCC

[HFW86] Christopher T. Haynes, Daniel P.

Friedman, and Mitchell Wand. Obtaining coroutines with continuations. Computer Languages, 11(3-4):143–153, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hendren:1993:DPL

[HG93] Laurie J. Hendren and Guang R.

Gao. Designing programming languages for the analyzability of pointer data structures. Computer Languages, 19(2): 119–134, April 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Herzeel:2009:FCH

[HGC^{+}09] Charlotte Herzeel, Kris Gybels, Pascal Costanza, Coen De Roover, and Theo D’Hondt. Forward chaining in HALO: An implementation strategy for history-based logic pointcuts. Computer Languages, Systems and Structures, 35(1): 31–47, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842408000250.

Hirschfeld:2006:OA

[HH06] Robert Hirschfeld and Stefan Hanenberg. Open aspects. Computer Languages, Systems and Structures, 32 (2–3):87–108, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1477842405000400.

Hansen:1989:IRD

[HHLv89] Michael R. Hansen, Bo S. Hansen, Peter Lucas, and Peter van Emde Boas. Integrating relational databases and constraint languages. Computer Languages, 14(2):63–82, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hayes:1990:IES

[HHS90] Roger Hayes, Norman C. Hutchinson, and Richard D. Schlichting. Integrating Emerald into a system for mixedlanguage programming. Computer Languages, 15(2):95–108, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hsiung:2008:ASV

[HL08] Pao-Ann Hsiung and Shang-Wei Lin. Automatic synthesis and verification of realtime embedded software for mobile and ubiquitous systems. Computer Languages, Systems and Structures, 34(4):153–169, December 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000206.

Huang:1976:DIL

[HLJ76] H. K. Huang, R. S. Ledley, and P. Johnson. DILOGdigitaled intelligence logic machine. Computer Languages, 2 (1-2):27–43, ???? 1976. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hage:2018:SIT

[HM18] Jurriaan Hage and Jay Mc-Carthy. Special issue on Trends in Functional Programming 2013/14. Computer Languages, Systems and Structures, 52(??): 63–64, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300775.

Hassan:2018:EIM

[HMHS18] Somaia A. Hassan, Mountasser M. M. Mahmoud, A. M. Hemeida, and Mahmoud A. Saber. Effective implementation of matrix–vector multiplication on Intel’s AVX multicore processor. Computer Languages, Systems and Structures, 51(??):158–175, January 2018. CODEN ????

ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417300428.

Huang:1990:EHM

[HO90] Daniel T. Huang and Ronald A. Olsson. An exception handling mechanism for SR. Computer Languages, 15(3):163–176, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hoare:1975:PPA

[Hoa75] C. A. R. Hoare. Parallel programming: an axiomatic approach. Computer Languages, 1(2):151–160, June 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hooper:1987:LFD

[Hoo87] James W. Hooper. Language features for discrete simulation. Computer Languages, 12 (1):39–46, ???? 1987. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hooper:1989:LFP

[Hoo89] James W. Hooper. Language features for prototyping and simulation support of the software life cycle. Computer Languages, 14(2):83–92, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Horspool:1990:IGL

[Hor90] R. Nigel Horspool. Incremental generation of LR parsers. Computer Languages, 15(4): 205–223, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Horspool:1993:RAP

[Hor93] R. Nigel Horspool. Recursive ascent-descent parsing. Computer Languages, 18(1):1–15, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Horry:2017:FID

[Hor17] Youichi Horry. Financial information description language and visualization/analysis tools. Computer Languages, Systems and Structures, 50(??):31–52, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301993.

Hsia:1991:IDC

[HR91] Pei Hsia and David C. Rine.

Introduction: Development of computer languages and software. Computer Languages, 16 (2):109–111, 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hsia:1992:ISP

[HR92] Pei Hsia and David C. Rine.

Introduction: Special-purpose languages. Computer Languages, 17(3):155–156, July 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hammond:1984:SSE

[HRS84] K. Hammond and V. J. Rayward-Smith. A survey on syntactic error recovery and repair. Computer Languages, 9 (1):51–67, ???? 1984. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Holldobler:2018:SLE

[HRW18] Katrin Hölldobler, Bernhard Rumpe, and Andreas Wortmann. Software language engineering in the large: towards composing and deriving languages. Computer Languages, Systems and Structures, 54(??):386–405, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300459.

Hill:2003:LPC

[HS03] Patricia M. Hill and Fausto Spoto. Logic programs as compact denotations. Computer Languages, Systems and Structures, 29(3):45–73, October 2003. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Haripriyan:1988:CWS

[HSS88] H. K. Haripriyan, Y. N.

Srikant, and Priti Shankar. A compiler writing system based on affix grammars. Computer Languages, 13(1):1–11, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Haraburda:2013:BTC

[HT13] David Haraburda and Paul Tarau. Binary trees as a computational framework. Computer Languages, Systems and Structures, 39(4):163–181, December 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842413000183.

Hughes:1985:PLE

[Hug85] Herman D. Hughes. A programming language engineered for beginners. Computer Languages, 10(1):23–36, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hull:1987:OPL

[Hul87] M. Elizabeth C. Hull. Occam — a programming language for multiprocessor systems. Computer Languages, 12(1):27–37, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Horspool:1993:SAP

[HV93] R. Nigel Horspool and Jan Vitek. Static analysis of PostScript code. Computer Languages, 19(2):65–78, April 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Hartel:1994:EDU

[HV94] Pieter H. Hartel and Willem G. Vree. Experiments with destructive updates in a lazy functional language. Computer Languages, 20(3):177–192, August 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Haubl:2013:CST

[HWM13] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Context-sensitive trace inlining for Java. Computer Languages, Systems and Structures, 39(4):123–141, December 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842413000146.

Heping:1996:ESL

[HZ96] H. Heping and H. Zedan. An executable specification language for fast prototyping parallel responsive systems. Computer Languages, 22(1):1–13, April 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ierusalimschy:1993:DAT

[Ier93] Roberto Ierusalimschy. A denotational approach for typechecking in object-oriented programming languages. Computer Languages, 19(1):19– 40, January 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Iranmanesh:2016:SSE

[IF16] Zeinab Iranmanesh and Mehran S.

Fallah. Specification and static enforcement of schedulerindependent noninterference in a middleweight Java. Computer Languages, Systems and Structures, 46(??):20–43, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300300.

Imanimehr:2018:TVS

[IF18] Fatemeh Imanimehr and Mehran S.

Fallah. On transparent valuesensitive run-time monitoring for information flow policies. Computer Languages, Systems and Structures, 54(??):273– 296, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300290.

Iyengar:1982:MLC

[IPF82] S. Sitharama Iyengar, N. Parameswaran, and John Fuller. A measure of logical complexity of programs. Computer Languages, 7 (3-4):147–160, ???? 1982. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ierusalimschy:1995:SFF

[IR95] Roberto Ierusalimschy and Noemi Rodriguez. Sideeffect free functions in objectoriented languages. Computer Languages, 21(3-4):129–146, October–December 1995. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Insa:2018:AAJ

[IS18] David Insa and Josep Silva. Automatic assessment of Java code. Computer Languages, Systems and Structures, 53 (??):59–72, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842417301045.

Inostroza:2017:MII

[IvdS17] Pablo Inostroza and Tijs van der Storm. Modular interpreters with implicit context propagation. Computer Languages, Systems and Structures, 48(??):39– 67, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300021.

Jalote:1992:STA

[Jal92] Pankaj Jalote. Specification and testing of abstract data types. Computer Languages, 17 (1):75–82, ???? 1992. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Jayaraman:1992:SAL

[Jay92] Bharat Jayaraman. Sublist assertions for listless and lazy evaluation. Computer Languages, 17(2):133–146, April 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Johnson:1994:FSP

[JD94] Gregory F. Johnson and Dominic Duggan. First-class stores and partial continuations in a programming language and environment. Computer Languages, 20(1):53–68, March 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Jenkins:1989:LBN

[JG89] Michael A. Jenkins and Janice I. Glasgow. A logical basis for nested array data structures. Computer Languages, 14 (1):35–51, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Jayaraman:1998:PLG

[JGM98] Bharat Jayaraman, Kannan Govindarajan, and Surya Mantha. Preference logic grammars. Computer Languages, 24(3):179–196, October 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/19/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/19/19/article.
pdf.

Jafari:2016:SMC

[JKK^{+}16] Ali Jafari, Ehsan Khamespanah, Haukur Kristinsson, Marjan Sirjani, and Brynjar Magnusson. Statistical model checking of Timed Rebeca models. Computer Languages, Systems and Structures, 45(??): 53–79, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416000051.

Jenkins:1996:PTS

[JL96] S. L. Jenkins and G. T. Leavens. Polymorphic typechecking in Scheme. Computer Languages, 22(4):215– 223, December 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Li:2009:UDS

[jLtCxH09] Zhi jie Li, Chun tian Cheng, and Fei xue Huang. Utilitydriven solution for optimal resource allocation in computational grid. Computer Languages, Systems and Structures, 35(4):406–421, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000390.

Jacquet:1996:TRH

[JM96] Jean-Marie Jacquet and Luís Monteiro. Towards resource handling in logic programming: the PPL framework and its semantics. Computer Languages, 22(2–3):51–77, July–October 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Jouvelot:2011:DVT

[JO11] Pierre Jouvelot and Yann Orlarey. Dependent vector types for data structuring in multirate Faust. Computer Languages, Systems and Structures, 37(3): 113–131, July 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000029.

Johnson:1981:DRD

[Joh81] Mark Scott Johnson. Dispel — a run-time debugging language. Computer Languages, 6(2):79–94, ???? 1981. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Joseloff:1978:CPC

[Jos78] S. H. Joseloff. Copyright protection for computer programs and other computer readable works. Computer Languages, 3 (4):265–266, ???? 1978. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Jung:2008:EEH

[JPB^{+}08] Dong-Heon Jung, JongKuk Park, Sung-Hwan Bae, Jaemok Lee, and Soo-Mook Moon. Efficient exception handling in Java bytecode-to-C aheadof-time compiler for embedded systems. Computer Languages, Systems and Structures, 34(4):170–183, December 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842407000176.

Joy:1985:ECC

[JRSB85] M. S. Joy, V. J. Rayward-Smith, and F. W. Burton. Efficient combinator code. Computer Languages, 10(3-4):211– 224, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kumar:2007:MDR

[KA07] Rajeev Kumar and Vikram Agrawal. Multiple dispatch in reflective runtime environment. Computer Languages, Systems and Structures, 33(2): 60–78, July 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000382.

Kowalski:2017:OLI

[KA17] Tomasz Marek Kowalski and Radosław Adamus. Optimisation of language-integrated queries by query unnesting. Computer Languages, Systems and Structures, 47 (Part 2)(??): 131–150, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300240.

Kampen:1975:FDS

[KB75] G. Kampen and J. L. Baer.

The formal definition of semantics by string automata. Computer Languages, 1(2):121–138, June 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Khedker:2003:BDF

[KDM03] Uday P. Khedker, Dhananjay M. Dhamdhere, and Alan Mycroft. Bidirectional data flow analysis for type inferencing. Computer Languages, Systems and Structures, 29(1–2): 15–44, April/July 2003. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Kennedy:1978:UCA

[Ken78] Ken Kennedy. Use-definition chains with applications. Computer Languages, 3(3):163–179, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kessler:1998:SED

[Kes98] Christoph W. Kessler. Scheduling expression DAGs for minimal register need. Computer Languages, 24(1):33– 53, April 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/17/18/article.
pdf.

Kafle:2017:HCV

[KG17] Bishoksan Kafle and John P. Gallagher. Horn clause verification with convex polyhedral abstraction and tree automatabased refinement. Computer Languages, Systems and Structures, 47 (part 1)(??):2– 18, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000822.

Kardas:2017:SIM

[KGS17] Geylani Kardas and Jorge J. Gomez-Sanz. Special issue on model-driven engineering of multi-agent systems in theory and practice. Computer Languages, Systems and Structures, 50(??):140– 141, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241730101X.

Koutavas:2012:FOR

[KH12] Vasileios Koutavas and Matthew Hennessy. First-order reasoning for higher-order concurrency. Computer Languages, Systems and Structures, 38(3):242– 277, October 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842412000139.

Khan:2010:FDS

[Kha10] Minhaj Ahmad Khan. Feedbackdirected specialization of code. Computer Languages, Systems and Structures, 36(1): 2–15, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000025.

Khan:2011:IPT

[Kha11] Minhaj Ahmad Khan. Improving performance through deep value profiling and specialization with code transformation. Computer Languages, Systems and Structures, 37(4):193– 203, October 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000170.

Ko:2014:SET

[KHO14] Yoonseok Ko, Kihong Heo, and Hakjoo Oh. A sparse evaluation technique for detailed semantic analyses. Computer Languages, Systems and Structures, 40(3–4):99–111, October/December 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842414000323.

Kirli:2002:DCT

[Kır02] Dilsun Kırlı. Distributed calltracking for security. Computer Languages, Systems and Structures, 28(1):129–154, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Krizevnik:2012:DBV

[KJ12] Marcel Krizevnik and Matjaz B. Juric. Data-bound variables for WS–BPEL executable processes. Computer Languages, Systems and Structures, 38(4):279–299, December 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842412000231.

Kannimoola:2017:TCO

[KJTA17] Jinesh M. Kannimoola, Bharat Jayaraman, Pallavi Tambay, and Krishnashree Achuthan. Temporal constrained objects: Application and implementation. Computer Languages, Systems and Structures, 49(??):82–100, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416302135.

Kislal:2018:DAS

[KK18] Orhan Kislal and Mahmut T.

Kandemir. Data access skipping for recursive partitioning methods. Computer Languages, Systems and Structures, 53(??):143–162, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300058.

Klerer:1992:LAP

[KKG92] Robert J. Klerer, Melvin Klerer, and Fred Grossman. A language for automated programming of mathematical applications. Computer Languages, 17(3):169–184, July 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kelefouras:2015:MSL

[KKG15] Vasilios Kelefouras, Angeliki Kritikakou, and Costas Goutis. A methodology for speeding up loop kernels by exploiting the software information and the memory architecture. Computer Languages, Systems and Structures, 41(??):

21–41, April 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000044.

Kanovich:2014:BMP

[KKNS14] Max Kanovich, Tajana Ban Kirigin, Vivek Nigam, and Andre Scedrov. Bounded memory protocols. Computer Languages, Systems and Structures, 40(3–4):137–154, October/December 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842414000347.

Kaufmann:2015:IIC

[KKP^{+}15] Petra Kaufmann, Martin Kronegger, Andreas Pfandler, Martina Seidl, and Magdalena Widl. Intra- and interdiagram consistency checking of behavioral multiview models. Computer Languages, Systems and Structures, 44 (part A)(??):72– 88, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000548.

Kurs:2015:BS

[KLIN15] Jan Kurs, Mircea Lungu, Rathesan Iyadurai, and Oscar Nierstrasz. Bounded seas. Computer Languages, Systems and Structures, 44 (part A)(??):114–140, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000536.

Karakoidas:2015:TSE

[KMLS15] Vassilios Karakoidas, Dimitris Mitropoulos, Panagiotis Louridas, and Diomidis Spinellis. A type-safe embedding of SQL into Java using the extensible compiler framework J%. Computer Languages, Systems and Structures, 41(??): 1–20, April 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000020.

Kieburtz:1985:DAL

[KN85] Richard B. Kieburtz and Bengt Nordstrom. Design of Apple — a language for modular programs. Computer Languages, 10(1):1–22, ???? 1985. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kwon:1994:IPT

[KNW94] Keehang Kwon, Gopalan Nadathur, and Debra Sue Wilson. Implementing polymorphic typing in a logic programming language. Computer Languages, 20(1):25–42, March 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kornilowicz:2015:FCM

[Kor15] Artur Kornilowicz. Flexary connectives in Mizar. Computer Languages, Systems and Structures, 44 (Part C) (??):238–250, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1477842415000421.

Kornilowicz:2016:FCM

[Kor16] Artur Kornilowicz. Flexary connectives in Mizar. Computer Languages, Systems and Structures, 45(??):238– 250, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000421.

Konopasek:1978:QAS

[KP78] Milos Konopasek and Christos Papaconstadopoulos. Question answering system on mathematical models (QAS): Description of the language. Computer Languages, 3(3):145–155, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kolesnichenko:2017:SCL

[KPN17] Alexey Kolesnichenko, Christopher M. Poskitt, and Sebastian Nanz. SafeGPU: Contractand library-based GPGPU for object-oriented languages. Computer Languages, Systems and Structures, 48(??):

68–88, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300057.

Kuhn:1993:CBV

[KPP93] Eva Kuhn, Herbert Pohlai, and Franz Puntigam. Concurrency and backtracking in Vienna Parallel Logic. Computer Languages, 19(3):185–203, July 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kessler:1995:GOC

[KR95] Christoph W. Kessler and Thomas Rauber. Generating optimal contiguous evaluations for expression DAGs. Computer Languages, 21(2): 113–127, July 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Kaser:1998:EIT

[KR98] Owen Kaser and C. R. Ramakrishnan. Evaluating inlining techniques. Computer Languages, 24(2):55– 72, July 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/18/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/18/17/article.
pdf.

Karimpour:2017:ERO

[KR17] Reza Karimpour and Guenther Ruhe. Evolutionary robust optimization for software product line scoping: an explorative study. Computer Languages, Systems and Structures, 47 (Part 2)(??):189– 210, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416301063.

Kreutzer:1990:CSF

[KS90] Wolfgang Kreutzer and Malcolm Stairmand. C-Flavours: a Scheme-based flavour system with coroutines and its application to the design of object-oriented simulation software. Computer Languages, 15 (4):225–249, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Knobe:1975:SST

[KY75] B. Knobe and G. Yuval. Some steps towards a better PAS-CAL. Computer Languages, 1 (4):277–286, ???? 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Leszczylowski:1989:PLS

[LB89] Jacek Leszczylowski and James M. Bieman. Prosper. A language for specification by prototyping. Computer Languages, 14 (3):165–180, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Lusth:2006:MAO

[LB06] John C. Lusth and R. Sean Bowman. A minimalist approach to objects. Computer Languages, Systems and Structures, 32(1):42–55, April 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

LeMoulec:2018:APE

[LBGA18] Gwendal Le Moulec, Arnaud Blouin, Valérie Gouranton, and Bruno Arnaldi. Automatic production of end user documentation for DSLs. Computer Languages, Systems and Structures, 54(??):337– 357, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301811.

Lepage:1981:OHD

[LBR81] Mary T. Lepage, David T.

Barnard, and Andres Rudmik. Optimization of hierarchical directed graphs. Computer Languages, 6(1):19–34, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Lee:2002:SDM

[LC02] Woo Hyong Lee and Morris Chang. A study of dynamic memory management in C++ programs. Computer Languages, Systems and Structures, 28(3):237–272, October 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Li:2007:PRC

[LCC07] Hui-Xian Li, Chun-Tian Cheng, and K. W. Chau. Parallel resource co-allocation for the computational grid. Computer Languages, Systems and Structures, 33(1): 1–10, April 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000412.

Lanvin:2010:EOO

[LCFÁ10] Daniel Fernández Lanvin, Raúl Izquierdo Castanedo, Aquilino Adolfo Juan Fuente, and Alberto Manuel Fernández Álvarez. Extending objectoriented languages with backward error recovery integrated support. Computer Languages, Systems and Structures, 36(2): 123–141, July 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000141.

Lienhard:2009:TOC

[LDG09] Adrian Lienhard, Stéphane Ducasse, and Tudor Gîrba. Taking an object-centric view on dynamic information with object flow analysis. Computer Languages, Systems and Structures, 35(1):63–79, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000201.

Le:2018:DDD

[LDN18] Duc Minh Le, Duc-Hanh Dang, and Viet-Ha Nguyen. On domain driven design using annotation-based domain specific language. Computer Languages, Systems and Structures, 54(??):199–235, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241730204X.

Ledley:1999:TCS

[Led99a] R. S. Ledley. 3-Tier Client/ Server at Work, Revised Edition; By Jeri Edwards. John Wiley & Sons, Inc. ISBN: 0-471-31502-8. Computer Languages, 25(2):116– 117, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/31/article.
pdf.

Ledley:1999:BMS

[Led99b] R. S. Ledley. Building Microsoft SQL Server 7 Applications with COM ; By Sanjiv Purba. John Wiley & Sons, Inc. ISBN: 0-471-19233-3. Computer Languages, 25(2):114– 115, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/29/article.
pdf.

Ledley:1999:BTA

[Led99c] R. S. Ledley. Building N-Tier Applications with COM And Visual Basic 6.0 ; Ash Rofail, Tony Martin. John Wiley & Sons, Inc. ISBN: 0-471-29549-3. Computer Languages, 25(2): 115–116, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/30/article.
pdf.

Ledley:1999:CYP

[Led99d]  R. S. Ledley.  Countdown

Y2K ; Peter de Jager, Richard Bergeon. John Wiley & Sons, Inc. ISBN: 0-471-32734-4. Computer Languages, 25(2): 113, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/

27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/27/article.
pdf.

Ledley:1999:MPD

[Led99e] R. S. Ledley. Mobile Processing in Distributed and Open Environments; Peter Sapaty. John Wiley and Sons, Inc. ISBN: 0-471-19572-3. Computer Languages, 25(2):113– 114, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/28/article.
pdf.

Lee:2005:PRG

[Lee05] Gyung-Ok Lee. On the prediction of reduction goals: a deterministic approach. Computer Languages, Systems and Structures, 31(2):49–70, July 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Lee:2000:ESI

[LfL00] Jimmy H. M. Lee and Ho fung Leung. An execution scheme for interactive problem-solving in concurrent constraint logic programming languages. Computer Languages, 25(3):119– 144, October 1, 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/29/

                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/15/18/28/29/25/article.
                                        pdf.

Li:1996:BEI

[Li96] W. X. Li. Building efficient incremental LL parsers by augmenting LL tables and threading parse trees. Computer Languages, 22(4):225–235, December 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Liao:1992:RPP

[Lia92] Husheng Liao. Removing partial parametrization for efficient implementation of functional languages. Computer Languages, 17(4):241–250, October 1992. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Liu:1988:SPM

[Liu88] Ken-Chih Liu. On string pattern matching: a quantitative analysis and a proposal. Computer Languages, 13(1):23–29, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Liu:1993:APL

[Liu93] Shaoying Liu. An abstract programming language and correctness proofs. Computer Languages, 18(4):273–282, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Luttighuis:2001:VBP

[LLvdW^{+}01] Paul Oude Luttighuis, Marc Lankhorst, Rob van de Wetering, René Bal, and Harmen van den Berg. Visualising business processes. Computer Languages, 27(1– 3):39–59, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

/www.elsevier.com/gej-ng/
10/15/18/54/27/30/abstract.
html.

Lamma:1993:PCM

[LMR93] Evelina Lamma, Paola Mello, and Gianfranco Rossi. Parametric composable modules in a logic programming language. Computer Languages, 18(2): 105–123, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Lozinskii:1986:PP

[LN86] Eliezer L. Lozinskii and Sergei Nirenburg. Parsing in parallel. Computer Languages, 11 (1):39–51, ???? 1986. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Lin:1991:FTF

[LN91] Kwei-Jay Lin and Swaminathan Natarajan. FLEX: Towards flexible real-time programs. Computer Languages, 16(1):65–79, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Logozzo:2009:CIA

[Log09]  Francesco Logozzo.  Class

invariants as abstract interpretation of trace semantics. Computer Languages, Systems and Structures, 35(2): 100–142, July 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000345.

Loulergue:2007:ISI

[Lou07]  Frédéric Loulergue.  Intro-

duction to the special issue on semantics and costs models for high-level parallel programming. Computer Languages, Systems and Structures, 33(3–4):79–81, October/December 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000340.

Lee:1997:OLI

[LP97] J. H. M. Lee and P. K. C. Pun. Object logic integration: a multiparadigm design methodology and a programming language. Computer Languages, 23(1):25–42, April 1997. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Loidl:2016:ESI

[LP16] Hans-Wolfgang Loidl and Ricardo Peña. Editorial of special issue trends in functional programming 2011/12. Computer Languages, Systems and Structures, 44 (part B):141– 142, ???? 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000603.

Liu:2017:ACS

[LR17] Jiangchao Liu and Xavier Rival. An array content static analysis based on noncontiguous partitions. Computer Languages, Systems and Structures, 47 (part 1)(??):104– 129, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416000063.

Lam:2011:MOE

[LRB^{+}11] Chi-Chung Lam, Thomas Rauber, Gerald Baumgartner, Daniel Cociorva, and P. Sadayappan. Memoryoptimal evaluation of expression trees involving large objects. Computer Languages, Systems and Structures, 37(2): 63–75, July 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000278.

Lafora:1984:REG

[LS84] Fernando Lafora and Mary Lou Soffa. Reverse execution in a generalized control regime. Computer Languages, 9(3-4): 183–192, ???? 1984. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Liu:1990:EHR

[LS90] Leo Y. Liu and R. K. Shyamasundar. Exception handling in RT-CDL. Computer Languages, 15(3):177–192, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Liu:1994:RCD

[LS94] Leo YuHsiang Liu and R. K. Shyamasundar. RT-CDL: a distributed real-time design language and its operational semantics. Computer Languages, 20(1):1–23, March 1994. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Luqi:1993:RTC

[Luq93] Luqi. Real-time constraints in a rapid prototyping language. Computer Languages, 18(2):77–103, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Lusth:2002:USL

[Lus02] John C. Lusth. Unified selection from lists, arrays, and objects. Computer Languages, Systems and Structures, 28(3): 289–305, October 2002. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Ledley:1975:PHQ

[LW75] R. S. Ledley and J. B. Wilson.

The precise handling of qualitative relationships. Computer Languages, 1(1):83–99, January 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mendez-Acuna:2016:LSP

[MAGD^{+}16] David Méndez-Acuña, José A.

Galindo, Thomas Degueule, Benoît Combemale, and Benoît Baudry. Leveraging software product lines engineering in the development of external DSLs: a systematic literature review. Computer Languages, Systems and Structures, 46(??):206– 235, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300768.

Malton:1993:DSF

[Mal93] Andrew Malton. The denotational semantics of a functional tree-manipulation language. Computer Languages, 19(3):157–168, July 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Malkov:2010:CFP

[Mal10] Sasa N. Malkov. Customizing a functional programming language for Web development. Computer Languages, Systems and Structures, 36(4):345–351, December 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842410000138.

Malhotra:2017:SIS

[Mal17] Ruchika Malhotra. Special issue on search-based techniques and their hybridizations in software engineering. Computer Languages, Systems and Structures, 47 (Part 2)(??):151– 152, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416301658.

Manacher:1978:IVC

[Man78] Glenn K. Manacher. Improved version of the Cocke-Younger-Kasami algorithm. Computer Languages, 3(2):127–133, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mansurov:2001:ASS

[Man01] Nikolai Mansurov. Automatic synthesis of SDL from MSC and its applications in forward and reverse engineering. Computer Languages, 27 (1–3):115–136, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/54/27/33/
abstract.html.

Merwin:1975:DME

[MB75] R. E. Merwin and F. R. Broca.

Direct microprogrammed execution of the intermediate text from a high-level language compiler. Computer Languages, 1 (1):17–28, January 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mazaher:1985:DCO

[MB85] Shahrzade Mazaher and Daniel M.

Berry. Deriving a compiler from an operational semantics written in VDL. Computer Languages, 10(2):147–164, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mernik:2013:SIP

[MB13] Marjan Mernik and Barrett R.

Bryant. Special issue on the Programming Languages track at the 27th ACM Symposium on Applied Computing. Computer Languages, Systems and Structures, 39(4):121–122, December 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842413000171.

Mernik:2014:SIP

[MB14] Marjan Mernik and Barrett R. Bryant. Special issue on the programming languages track at the 28th ACM Symposium on Applied Computing. Computer Languages, Systems and Structures, 40 (1):1, April 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842414000050.

McNamee:1996:ISC

[MC96] C. M. McNamee and W. A. Crow. Inter-entry selection control mechanisms: implementation and evaluation. Computer Languages, 22(4):259– 278, December 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

McCrosky:1991:ICR

[McC91] Carl McCrosky. Intermediate container removal. Computer Languages, 16(2):179–195, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Martinez:2017:MBA

[MCC17] Salvador Martínez, Valerio Cosentino, and Jordi Cabot. Model-based analysis of Java EE web security misconfigurations. Computer Languages, Systems and Structures, 49(??):36–61, September 2017. CODEN ????

ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301348.

McKeeman:1975:MBM

[McK75] W. M. McKeeman. Mechanizing bankers’ morality. Computer Languages, 1(1):73–82, January 1975. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

McLeod:1977:HLD

[McL77] D. J. McLeod. High level definition of abstract domains in a relational data base system. Computer Languages, 2 (3):61–73, ???? 1977. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mondejar:2012:TPT

[MGLFCP12] Rubén Mondéjar, Pedro García-López, Enrique Fernández-Casado, and Carles Pairot. TaKo: Providing transparent collaboration on singleuser applications. Computer Languages, Systems and Structures, 38(1):108– 121, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000303.

Merlin:2007:BSP

[MH07] Armelle Merlin and Gaétan Hains. A bulk-synchronous parallel process algebra. Computer Languages, Systems and Structures, 33(3–4):111–133, October/December 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842406000327.

Michaelson:1986:IFG

[Mic86] Greg Michaelson. Interpreters from functions and grammars. Computer Languages, 11(2): 85–104, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Michel:1996:DID

[Mic96] Olivier Michel. Design and implementation of 8_{1/2}: a declarative data-parallel language. Computer Languages, 22(2–3):165–179, July–October 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mens:2006:CEC

[MKPW06] Kim Mens, Andy Kellens, Frédéric Pluquet, and Roel Wuyts. Co-evolving code and design with intensional views: a case study. Computer Languages, Systems and Structures, 32(2–3):140–156, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842405000394.

Miranda:2005:PFF

[MLW05] Eliot Miranda, David Leibs, and Roel Wuyts. Parcels: a fast and feature-rich binary deployment technology. Computer Languages, Systems and Structures, 31(3–4):165–181, October/December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

McDonald:1982:QLF

[MM82] Nancy H. McDonald and John P. McNally. Query language feature analysis by usability. Computer Languages, 7 (3-4):103–124, ???? 1982. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Marand:2015:DDS

[MMC15] Elaheh Azadi Marand, Elham Azadi Marand, and Moharram Challenger. DSML4CP: a Domain-specific Modeling Language for Concurrent Programming. Computer Languages, Systems and Structures, 44 (Part C)(??):319– 341, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000627.

Marand:2016:DDS

[MMC16] Elaheh Azadi Marand, Elham Azadi Marand, and Moharram Challenger. DSML4CP: a domain-specific modeling language for concurrent programming. Computer Languages, Systems and Structures, 45(??): 319–341, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000627.

Maurer:1983:UCT

[MO83] Peter M. Maurer and Arthur E. Oldehoeft. Use of combinators in translating a purely functional language to lowlevel data-flow graphs. Computer Languages, 8(1):27–45, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Morazan:2016:GAR

[Mor16] Marco T. Morazán. Generative and accumulative recursion made fun for beginners. Computer Languages, Systems and Structures, 44 (part B): 181–197, ???? 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000524.

Mano:1984:NPE

[MOT84] Yoshihisa Mano, Kazuhito Ohmaki, and Koji Torii. New programming environment with a multi-display terminal and early experiences with it. Computer Languages, 9(1):39–49, ???? 1984. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Matwin:1985:PPR

[MP85] Stanislaw Matwin and Tomasz Pietrzykowski. Prograph: a preliminary report. Computer Languages, 10(2):91–126, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Myers:1992:ITC

[MP92] Heidi E. Myers and James M.

Purtilo. Interface type checking of large C applications. Computer Languages, 17(2): 147–154, April 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mosconi:2000:ICD

[MP00] M. Mosconi and M. Porta.

Iteration constructs in dataflow visual programming languages. Computer Languages, 26(2–4):67–104, July 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/50/31/27/
abstract.html.

Milewicz:2017:RSH

[MP17] Reed Milewicz and Peter Pirkelbauer. Refinement of structural heuristics for model checking of concurrent programs through data mining. Computer Languages, Systems and Structures, 47 (Part 2)(??): 170–188, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300793.

Murching:1990:IRD

[MPS90] Arvind M. Murching, Y. V. Prasad, and Y. N. Srikant. Incremental recursive descent parsing. Computer Languages, 15(4):193–204, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Maraninchi:2001:AAB

[MR01] Florence Maraninchi and Yann Rémond. Argos: an automatonbased synchronous language. Computer Languages, 27(1– 3):61–92, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

/www.elsevier.com/gej-ng/
10/15/18/54/27/31/abstract.
html.

Meenakshi:2004:RAL

[MR04] B. Meenakshi and R. Ramanujam. Reasoning about layered message passing systems. Computer Languages, Systems and Structures, 30(3–4):171– 206, October/December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Maris:2003:DRT

[MRO03] Justin T. Maris, Matthew D. Roper, and Ronald A. Olsson. DescaRTeS: a run-time system with SR-like functionality for programming a network of embedded systems. Computer Languages, Systems and Structures, 29(4):75–100, December 2003. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Murching:1989:IAE

[MS89] Arvind M. Murching and Y. N.

Srikant. Incremental attribute evaluation through recursive procedures. Computer Languages, 14(4):225–237, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

McCrosky:1993:STP

[MS93] Carl McCrosky and Ken Sailor.

A synthesis of type-checking and parsing. Computer Languages, 18(4):241–250, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Malik:2010:SGL

[MSRG10] Avinash Malik, Zoran Salcic, Partha S. Roop, and Alain Girault. SystemJ: a GALS language for system level design. Computer Languages, Systems and Structures, 36(4):317–344, December 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000023.

Maier:2014:RSS

[MST14] P. Maier, R. Stewart, and P. W. Trinder. Reliable scalable symbolic computation: the design of SymGrid-Par2. Computer Languages, Systems and Structures, 40(1): 19–35, April 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842414000049.

Magnenat-Thalmann:1982:CIL

[MT82] Nadia Magnenat-Thalmann. Choosing an implementation language for automatic translation. Computer Languages, 7 (3-4):161–170, ???? 1982. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mens:2005:DSC

[MT05] Kim Mens and Tom Tourwé. Delving source code with formal concept analysis. Computer Languages, Systems and Structures, 31(3–4):183–197, October/December 2005. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Milewicz:2016:LRC

[MVT^{+}16] Reed Milewicz, Rajesh Vanka, James Tuck, Daniel Quinlan, and Peter Pirkelbauer. Lightweight runtime checking of C programs with RTC. Computer Languages, Systems and Structures, 45(??):191– 203, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S1477842416000026.

Messerschmidt:1982:CCO

[MW82] J. Messerschmidt and R. Wilhelm. Constructors from composed objects. Computer Languages, 7(2):53–59, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mitchell:1996:SCO

[MW96] S. E. Mitchell and A. J.

Wellings. Synchronisation, concurrent object-oriented programming and the inheritance anomaly. Computer Languages, 22(1):15–26, April 1996. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Mejri:2017:FSI

[MY17] Mohamed Mejri and Hamdi Yahyaoui. Formal specification and integration of distributed security policies. Computer Languages, Systems and Structures, 49(??):1–35, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300707.

Mernik:2005:IPL

[MŽ05] Marjan Mernik and Viljem Žumer. Incremental programming language development. Computer Languages, Systems and Structures, 31(1):1–16, April 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Mateos:2007:EMS

[MZC07] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. Extending movilog for supporting Web services. Computer Languages, Systems and Structures, 33(1):11–31, April 2007. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000029.

Mateos:2010:ANI

[MZC10] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. An approach for non-intrusively adding malleable fork/join parallelism into ordinary JavaBean compliant applications. Computer Languages, Systems and Structures, 36(3):288– 315, October 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842410000035.

Mandrioli:1985:MAT

[MZGT85] Dino Mandrioli, Roberto Zicari, Carlo Ghezzi, and Francesco Tisato. Modeling the Ada task system by Petri nets. Computer Languages, 10(1):43–61, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nagata:1979:ELN

[Nag79] Hiroyasu Nagata. EQL: a language for numerical computation. Computer Languages, 4 (1):17–27, ???? 1979. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nagata:1980:FLM

[Nag80] Hiroyasu Nagata. Formal: a language with a macro-oriented extension facility. Computer Languages, 5(2):65–76, ???? 1980. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nirenburg:1984:HHU

[NB84] Sergei Nirenburg and Yosef Ben Asher. Huhu: the Hebrew University Hebrew Understander. Computer Languages, 9(3-4): 161–182, ???? 1984. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Newton:1977:SLS

[ND77] G. E. Newton and J. Denbigh Starkey. SESPOOL, a language for systems programming. Computer Languages, 2 (3):75–91, ???? 1977. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nam:1993:CSP

[NH93] Young K. Nam and Lawrence J.

Henschen. A controlling scheme for Prolog through count terms.

Computer Languages, 19(4): 235–245, October 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nilsen:1990:SDT

[Nil90]  Kelvin Nilsen.  A stream

data type that supports goaldirected pattern matching on unbounded sequences of values. Computer Languages, 15 (1):41–54, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nigam:2012:MDL

[NJLS12] Vivek Nigam, Limin Jia, Boon Thau Loo, and Andre Scedrov. Maintaining distributed logic programs incrementally. Computer Languages, Systems and Structures, 38(2): 158–180, July 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842412000024.

Nawrocki:1990:DOA

[NK90] J. R. Nawrocki and C. H. A. Koster. On display optimization for Algol-like languages. Computer Languages, 15(1): 27–39, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ng:1995:MLO

[NL95] K. W. Ng and C. K. Luk. I^{+}: a multiparadigm language for object-oriented declarative programming. Computer Languages, 21(2):81–100, July 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Naumchev:2017:SR

[NM17] A. Naumchev and B. Meyer.

Seamless requirements. Computer Languages, Systems and Structures, 49(??):119–132, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301981.

Nielson:2009:MFC

[NN09] Hanne Riis Nielson and Flemming Nielson. A monotone framework for CCS. Computer Languages, Systems and Structures, 35(4):365–394, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000286.

Nielson:2017:AGI

[NN17] Flemming Nielson and Hanne Riis Nielson. Atomistic Galois insertions for flow sensitive integrity. Computer Languages, Systems and Structures, 50(??):82–107, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300052.

Noonan:1985:AGA

[Noo85] Robert E. Noonan. An algorithm for generating abstract syntax trees. Computer Languages, 10(3-4):225–236, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nosal:2017:CHI

[NPS17] Milan Nosál, Jaroslav Porubän, and Matús Sulír. Customizing host IDE for nonprogramming users of pure embedded DSLs: a case study. Computer Languages, Systems and Structures, 49(??):101–118, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416301269.

Nederhof:1993:PEG

[NS93] Mark-Jan Nederhof and Janos J. Sarbo. Partial evaluation grammars. Computer Languages, 18 (2):137–151, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Nymeyer:1995:GSH

[Nym95] Albert Nymeyer. A grammatical specification of humancomputer dialogue. Computer Languages, 21(1):1–16, April 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Oubelli:2018:SMB

[OAB^{+}18] Lynda Ait Oubelli, Yamine Aït Ameur, Judicaël Bedouet, Romain Kervarc, Benoît Chausserie-Laprée, and Béatrice Larzul. A scalable model based approach for data model evolution: Application to space missions data models. Computer Languages, Systems and Structures, 54(??):358–385, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300447.

Olsson:2002:FSI

[OBGK02] Ronald A. Olsson, Gregory D.

Benson, Tingjian Ge, and Aaron W. Keen. Fairness in shared invocation servicing. Computer Languages, Systems and Structures, 28(4):327–351, December 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Osterbye:2000:SAB

[ØK00] Kasper Østerbye and Wolfgang Kreutzer. Synchronization abstraction in the BETA programming language. Computer Languages, 25(3):165– 187, October 1, 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/29/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/15/18/28/29/27/article.
                                        pdf.

Ozkaya:2018:SUB

[OK18] Mert Ozkaya and Mehmet Alp Kose. SAwUML — UMLbased, contractual software architectures and their formal analysis using SPIN. Computer Languages, Systems and Structures, 54(??):71–94, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417301550.

Oudshoorn:1991:LOM

[OM91] Michael J. Oudshoorn and Chris D. Marlin. A layered, operational model of data control in programming languages. Computer Languages, 16(2): 147–165, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Olsson:1992:ISN

[OM92] Ronald A. Olsson and Carole M. McNamee. Inter-entry selection: non-determinism and explicit control mechanisms. Computer Languages, 17(4): 269–282, October 1992. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Oliveira:2018:CNC

[ONB^{+}18] Pedro Oliveira, Pedro Santos Neto, Ricardo Britto, Ricardo Rabêlo, Ronyerison Braga, and Matheus Souza. CIaaS — computational intelligence as a service with Athena. Computer Languages, Systems and Structures, 54(??):95–118, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301458.

Ouared:2018:QQM

[OOB18] Abdelkader Ouared, Yassine Ouhammou, and Ladjel Bellatreche. QoSMOS: QoS metrics management tool suite. Computer Languages, Systems and Structures, 54(??):236– 251, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300162.

Orman:1983:FSL

[Orm83] Levent Orman. Familial specification language for database application systems. Computer Languages, 8(3-4):113– 124, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

OBagy:1993:OSI

[OWG93] Janalee O’Bagy, Kenneth Walker, and Ralph E. Griswold. An operational semantics for Icon: implementation of a procedural goal-directed language. Computer Languages, 18(4):217–239, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ozturk:2011:RMS

[Ozt11]  Ozcan Ozturk.  Reducing

memory space consumption through dataflow analysis. Computer Languages, Systems and Structures, 37(4):168– 177, October 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842411000157.

Pagan:1978:FSS

[Pag78] F. G. Pagan. Formal semantics of a Snobol4 subset. Computer Languages, 3(1):13–30, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pagan:1979:SSU

[Pag79]  Frank G. Pagan.  Seman-

tic specification using twolevel grammars: Blocks, procedures and parameters. Computer Languages, 4(3-4):171– 185, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pai:2016:DRE

[Pai16] Rekha R. Pai. Detection of redundant expressions: a precise, efficient, and pragmatic algorithm in SSA. Computer Languages, Systems and Structures, 46(??):167–181, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300622.

Patnaik:1984:ICD

[PB84] L. M. Patnaik and B. R. Badrinath. Implementation of CSP-S for description of distributed algorithms. Computer Languages, 9(3-4):193–202, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Peck:2012:OSC

[PBDF12] Mariano Martinez Peck, Noury Bouraqadi, Stéphane Ducasse, and Luc Fabresse. Object swapping challenges: an evaluation of imagesegment. Computer Languages, Systems and Structures, 38(1): 1–15, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000315.

Patnaik:1984:DDF

[PBG84] L. M. Patnaik, Prabal Bhattacharya, and R. Ganesh. DFL: a data flow language. Computer Languages, 9(2):97–106, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pao:1978:SSI

[PC78] Tsyh-Wen Pao and W. Carr, III. A solution of the syntactical induction-inference problem for regular languages. Computer Languages, 3(1):53–64, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Patnaik:1985:GQH

[PC85] L. M. Patnaik and D. M. Chowdhary. Generalized query-by-rule: a heterogeneous database query language. Computer Languages, 10(3-4):165– 178, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Patrignani:2015:FAT

[PC15] Marco Patrignani and Dave Clarke. Fully abstract trace semantics for protected module architectures. Computer Languages, Systems and Structures, 42(??):22– 45, ???? 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000081.

Pinto:2018:ACM

[PCB^{+}18] Pedro Pinto, Tiago Carvalho, João Bispo, Miguel António Ramalho, and João M. P. Cardoso. Aspect composition for multiple target languages using LARA. Computer Languages, Systems and Structures, 53(??):1–26, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S147784241730115X.

Planas:2016:LSV

[PCG16] Elena Planas, Jordi Cabot, and Cristina Gómez. Lightweight and static verification of UML executable models. Computer Languages, Systems and Structures, 46(??):66–90, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300361.

Pavlatos:2009:ERE

[PDK^{+}09] Christos Pavlatos, Alexandros C. Dimopoulos, Andrew Koulouris, Theodore Andronikos, Ioannis Panagopoulos, and George Papakonstantinou. Efficient reconfigurable embedded parsers. Computer Languages, Systems and Structures, 35(2): 196–215, July 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000334.

Pohl:1988:ZCL

[PE88] Ira Pohl and Daniel Edelson.

A to Z: C language shortcomings. Computer Languages, 13 (2):51–64, ???? 1988. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Penna:2005:TSS

[Pen05] Giuseppe Della Penna. A type system for static and dynamic checking of C++ pointers. Computer Languages, Systems and Structures, 31(2):71– 101, July 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Penna:2014:MCX

[Pen14] Giuseppe Della Penna. Model checking XSL transformations. Computer Languages, Systems and Structures, 40(2): 73–97, ???? 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842414000062.

Peterson:1978:ESA

[Pet78] Norman D. Peterson. Elements of style among machinegenerated Cobol flowcharts. Computer Languages, 3(1):1– 11, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pereira:2016:OAD

[PFH16] Maria João Varanda Pereira, João Fonseca, and Pedro Rangel Henriques. Ontological approach for DSL development. Computer Languages, Systems and Structures, 45(??): 35–52, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S1477842415300270.

Papazoglou:1984:OPL

[PGM84] M. P. Papazoglou, P. I. Georgiadis, and D. G. Maritsas. An outline of the programming language SIMULA. Computer Languages, 9(2):107–131, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pontelli:1996:IEN

[PGT^{+}96] Enrico Pontelli, Gopal Gupta, Dongxing Tang, Manuel Carro, and Manuel V. Hermenegildo. Improving the efficiency of nondeterministic independent AND-parallel systems. Computer Languages, 22(2–3):115– 142, July–October 1996. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Purtilo:1991:EPD

[PJ91] James M. Purtilo and Pankaj Jalote. An environment for prototyping distributed applications. Computer Languages, 16 (3-4):197–207, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Placer:1991:MLG

[Pla91]  John Placer.  The multi-

paradigm language G. Computer Languages, 16(3-4):235– 258, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Popovic:2015:DMA

[PLDD15] Aleksandar Popovic, Ivan Lukovic, Vladimir Dimitrieski, and Verislav Djukic. A DSL for modeling application-specific functionalities of business applications. Computer Languages, Systems and Structures, 43(??):69–95, October 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000263.

Pontelli:2010:IPE

[PLS10] Enrico Pontelli, Hung Viet Le, and Tran Cao Son. An investigation in parallel execution of answer set programs on distributed memory platforms: Task sharing and dynamic scheduling. Computer Languages, Systems and Structures, 36(2): 158–202, July 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000244.

Papenhausen:2018:CAO

[PM18] Eric Papenhausen and Klaus Mueller. Coding ants: Optimization of GPU code using ant colony optimization. Computer Languages, Systems and Structures, 54(??):119– 138, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300137.

Pereira:2018:PRS

[PMK^{+}18] Juliana Alves Pereira, Pawel Matuszyk, Sebastian Krieter, Myra Spiliopoulou, and Gunter Saake. Personalized recommender systems for productline configuration processes. Computer Languages, Systems and Structures, 54(??):451– 471, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241730043X.

Park:2015:CVS

[PMS15] Heejong Park, Avinash Malik, and Zoran Salcic. Compiling and verifying SC-SystemJ programs for safety-critical reactive systems. Computer Languages, Systems and Structures, 44 (Part C)(??):251– 282, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241530004X.

Park:2016:CVS

[PMS16] Heejong Park, Avinash Malik, and Zoran Salcic. Compiling and verifying SC-SystemJ programs for safety-critical reactive systems. Computer Languages, Systems and Structures, 45(??):251– 282, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S147784241530004X.

Przigoda:2018:FCA

[PNF^{+}18] Nils Przigoda, Philipp Niemann, Jonas Gomes Filho, Robert Wille, and Rolf Drechsler. Frame conditions in the automatic validation and verification of UML/OCL models: a symbolic formulation of modifies only statements. Computer Languages, Systems and Structures, 54(??):512– 527, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417300131.

Pastrana:2011:QES

[PPK11] J. L. Pastrana, E. Pimentel, and M. Katrib. QoSenabled and self-adaptive connectors for Web services composition and coordination. Computer Languages, Systems and Structures, 37(1): 2–23, April 2011. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842410000242.

Perugini:2010:PTI

[PR10] Saverio Perugini and Naren Ramakrishnan. Program transformations for information personalization. Computer Languages, Systems and Structures, 36(3):223–249, October 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842409000256.

Pontelli:2002:ODS

[PRD02] E. Pontelli, D. Ranjan, and A. Dal Palú. An optimal data structure to handle dynamic environments in nondeterministic computations. Computer Languages, Systems and Structures, 28(2):181–201, December 15, 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Perin:2012:LSC

[PRR12] Fabrizio Perin, Lukas Renggli, and Jorge Ressia. Linguistic style checking with program checking tools. Computer Languages, Systems and Structures, 38(1):61–72, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842411000455.

Peterssen:1986:PDT

[PS86] Kent Petersson and Jan M.

Smith. Program derivation in type theory: a partitioning problem. Computer Languages, 11(3-4):161–172, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pradeep:1994:PAE

[PS94a] B. Pradeep and C. Siva Ram Murthy. Parallel arithmetic expression evaluation on reconfigurable meshes. Computer Languages, 20(4):267–277, November 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Pradeep:1994:PRP

[PS94b] B. Pradeep and C. Siva Ram Murthy. Parallel recognition and parsing on mesh connected computers with multiple broadcasting. Computer Languages, 20(1):43–51, March 1994. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Philippidis:2010:MRU

[PS10] Cesar J. Philippidis and Weijia Shang. On minimizing register usage of linearly scheduled algorithms with uniform dependencies. Computer Languages, Systems and Structures, 36(3):250–267, October 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842409000438.

Purtilo:1995:EPS

[PSW95] James M. Purtilo, Thomas M. Swiss, and Elizabeth L. White. Extracting program structure for packaging in a componentbased environment. Computer Languages, 21(1):39–48, April 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Prahofer:2013:MDS

[PSW^{+}13] Herbert Prähofer, Roland Schatz, Christian Wirth, Dominik Hurnaus, and Hanspeter Mössenböck. Monaco — a domain-specific language solution for reactive process control programming with hierarchical components. Computer Languages, Systems and Structures, 39(3):67–94, October 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842413000031.

Papadimitriou:2009:JIS

[PT09] Stergios Papadimitriou and Konstantinos Terzidis. jLab: Integrating a scripting interpreter with Java technology for flexible and efficient scientific computation. Computer Languages, Systems and Structures, 35(3):217–240, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784240800002X.

Polach:2016:EDV

[PTJM16] Radomír Polách, Jan Trávnícek, Jan Janousek, and Borivoj Melichar. Efficient determinization of visibly and heightdeterministic pushdown automata. Computer Languages, Systems and Structures, 46(??):91–105, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416300136.

Puntigam:2001:SID

[Pun01] Franz Puntigam. State inference for dynamically changing interfaces. Computer Languages, 27(4):163–202, December 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/15/18/54/31/28/
                                        abstract.html.

Razavi:2005:LSA

[RBY^{+}05] Reza Razavi, Noury Bouraqadi, Joseph Yoder, Jean-François Perrot, and Ralph Johnson. Language support for adaptive object-models using metaclasses. Computer Languages, Systems and Structures, 31(3– 4):199–218, October/December 2005. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Ripley:1978:SAS

[RD78] G. David Ripley and Frederick C. Druseikis. A statistical analysis of syntax errors. Computer Languages, 3(4):227–240, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Reis:2015:FGM

[RDB15] Leonardo V. S. Reis, Vladimir O.

Di Iorio, and Roberto S. Bigonha. An on-the-fly grammar modification mechanism for composing and defining extensible languages. Computer Languages, Systems and Structures, 42(??):46– 59, ???? 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000032.

Rothlisberger:2008:UPB

[RDT08] David Röthlisberger, Marcus Denker, and Éric Tanter. Unanticipated partial behavioral reflection: Adapting applications at runtime. Computer Languages, Systems and Structures, 34(2–3):46–65, July/October 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000127.

Reed:1984:ATA

[Ree84] Alan Reed. Anatomy of a text analysis package. Computer Languages, 9(2):89–96, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Reza:2012:JS

[Rez12] Juan Rolando Reza. Java supervenience. Computer Languages, Systems and Structures, 38(1):73–97, April 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842411000182.

Rondogiannis:1998:BTL

[RGP98] P. Rondogiannis, M. Gergatsoulis, and T. Panayiotopoulos. Branching-time logic programming: the language Cactus and its applications. Computer Languages, 24(3):155– 178, October 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/19/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/19/18/article.
pdf.

Rus:1994:ATL

[RH94] Teodor Rus and Tom Halverson. Algebraic tools for language processing. Computer Languages, 20(4):213– 238, November 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Radakovic:2018:TCE

[RH18] Davorka Radaković and ore Herceg. Towards a completely extensible dynamic geometry software with metadata. Computer Languages, Systems and Structures, 52(??): 1–20, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241730057X.

Rich:1980:MPT

[Ric80] Robert P. Rich. Mechanical proof testing. Computer Languages, 5(1):1–28, ???? 1980. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ricci:2016:PEL

[Ric16] Alessandro Ricci. Programming with event loops and control loops — from actors to agents. Computer Languages, Systems and Structures, 45(??): 80–104, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000949.

Riddle:1979:ASSa

[Rid79a] William E. Riddle. An approach to software system behavior description. Computer Languages, 4(1):29–47, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Riddle:1979:ASSb

[Rid79b] William E. Riddle. An approach to software system modelling and analysis. Computer Languages, 4(1):49–66, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Rine:1991:ICL

[Rin91] David C. Rine. Introduction: Computer language: a perspective. Computer Languages, 16(1):1–3, 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Reddy:1993:PAI

[RK93] Uday S. Reddy and Samuel N. Kamin. On the power of abstract interpretation. Computer Languages, 19(2):79–89, April 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Rieger:2018:POM

[RK18] Christoph Rieger and Herbert Kuchen. A process-oriented modeling approach for graphical development of mobile business apps. Computer Languages, Systems and Structures, 53(??):43–58, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301215.

Radha:1993:PIU

[RM93] S. Radha and C. R. Muthukrishnan. A portable implementation of UNITY on von Neumann machines. Computer Languages, 18(1):17–30, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Renggli:2009:TMD

[RN09] Lukas Renggli and Oscar Nierstrasz. Transactional memory in a dynamic language. Computer Languages, Systems and Structures, 35(1): 21–30, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000237.

RodriguesdaSilva:2015:MDE

[Rod15] Alberto Rodrigues da Silva.

Model-driven engineering: a survey supported by the unified conceptual model. Computer Languages, Systems and Structures, 43(??):139–155, October 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000408.

Romanovsky:1995:CO

[Rom95]  A. Romanovsky.  Conver-

sations of objects. Computer Languages, 21(3-4):147– 163, October–December 1995. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Romanovsky:1997:PEH

[Rom97] Alexander Romanovsky. Practical exception handling and resolution in concurrent programs. Computer Languages, 23(1):43–58, April 1997. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Rotenstreich:1992:OLF

[Rot92] Shmuel Rotenstreich. Orthogonal languages: a framework for program development. Computer Languages, 17(2):83–100, April 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Reghizzi:1998:GPM

[RP98] Stefano Crespi Reghizzi and Giuseppe Psaila. Grammar partitioning and modular deterministic parsing1. Computer Languages, 24(4):197–227, December 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/15/18/27/20/
                                        17/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/15/18/27/20/17/article.
                                        pdf.

Roldan:2009:SCL

[RPB09] Ana M. Roldan, Ernesto Pimentel, and Antonio Brogi. Software composition with Linda. Computer Languages, Systems and Structures, 35(4):395–405, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842408000389.

Reeves:1999:SBT

[RR99] A. C. Reeves and C. Rattray. A sketch-based technique for the derivation of self-interpreters. Computer Languages, 25(1):1– 37, April 1, 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/17/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/17/17/article.
pdf.

Ramanath:1982:OCF

[RS82] M. V. S. Ramanath and Marvin Solomon. Optimal code from flow graphs. Computer Languages, 7(1):41–52, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Ramanathan:1983:DIA

[RS83] J. Ramanathan and D. Soni.

Design and implementation of an adaptable software environment. Computer Languages, 8 (3-4):139–159, ???? 1983. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Rubenstein:1987:CEL

[RS87] Michael C. Rubenstein and Richard M. Salter. Computationally extended logic programming. Computer Languages, 12(1):1–7, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Runger:1994:POS

[RS94] Gudula Rünger and Kurt Sieber. A process oriented semantics of the PRAM-language FORK. Computer Languages, 20(4):253–265, November 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Rodrigues:2018:HDD

[RT18] Elder Rodrigues and Ricardo Terra. How do developers use dynamic features? The case of Ruby. Computer Languages, Systems and Structures, 53(??):73–89, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842417300945.

Ribic:2018:RRD

[RTMRK18] Samir Ribić, Razija Turcinhodzić, Amela Muratović-Ribić, and Tomaz Kosar. REDOSPLAT: a readable domain-specific language for timetabling requirements definition. Computer Languages, Systems and Structures, 54(??):252–272, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842418300368.

Runciman:1989:WAN

[Run89] Colin Runciman. What about the natural numbers? (in language design). Computer Languages, 14(3):181–191, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Rus:1987:AMP

[Rus87] T. Rus. An algebraic model for programming languages. Computer Languages, 12(3-4):173– 195, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Resler:2009:HOS

[RW09] R. Daniel Resler and Victor Winter. A higher-order strategy for eliminating common subexpressions. Computer Languages, Systems and Structures, 35(4):341–364, December 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000274.

Svenningsson:2016:CDS

[SA16] Josef Svenningsson and Emil Axelsson. Combining deep and shallow embedding of domainspecific languages. Computer Languages, Systems and Structures, 44 (part B):143– 165, ???? 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000500.

Salter:1983:CAI

[Sal83] Richard Salter. Concurrent applicative implementations of nondeterministic algorithms. Computer Languages, 8(2):61– 68, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Salzman:1992:ASM

[Sal92] Eric Salzman. Activity simulation in Modula-2: an exercise in language extension. Computer Languages, 17(1):39–60, ???? 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Samet:1979:DSB

[Sam79] Hanan Samet. Deep and shallow binding: the assignment operation. Computer Languages, 4(3-4):187–198, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sarwar:1993:RBS

[Sar93] S. Mansoor Sarwar. Run-time behavior of SASL programs: a performance study. Computer Languages, 19(3):205–211, July 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sarbo:1994:GTO

[Sar94] Janos J. Sarbo. Grammar transformations for optimizing backtrack parsers. Computer Languages, 20(2):89–100, May 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schwartz:1979:SVA

[SB79] Richard L. Schwartz and Daniel M. Berry. A semantic view of ALGOL 68. Computer Languages, 4(1):1–15, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Scharli:2004:BIP

[SB04] Nathanael Schärli and Andrew P. Black. A browser for incremental programming. Computer Languages, Systems and Structures, 30(1–2):79–95, April/July 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Salter:1980:CLC

[SBF80] Richard M. Salter, Terence J.

Brennan, and Daniel P. Friedman. Concur: a language for continuous, concurrent processes. Computer Languages, 5 (3-4):163–189, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Strothotte:1987:SPL

[SC87] Thomas W. Strothotte and Gordon V. Cormack. Structured program lookahead. Computer Languages, 12(2): 95–108, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Shen:1994:ACP

[SC94] Jun Shen and Gordon V. Cormack. Access control for private declarations in Ada. Computer Languages, 20(2):117– 126, May 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schwartz:1975:OVHb

[Sch75a] J. T. Schwartz. Optimisation of very high level languages. II. deducing relationships of inclusion and membership. Computer Languages, 1(3):197–218, September 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schwartz:1975:OVHa

[Sch75b] J. T. Schwartz. Optimization of very high level languages-value transmission and its corollaries. Computer Languages, 1 (2):161–194, June 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schwartz:1976:WPS

[Sch76] J. T. Schwartz. What programmers should know. Computer Languages, 2(1-2):21–25, ???? 1976. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schwartz:1978:PCD

[Sch78] Richard L. Schwartz. Parallel compilation: a design and its application to Simula 67. Computer Languages, 3(2):75–94, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Scott:1991:LDP

[Sco91] Michael L. Scott. The Lynx distributed programming language: motivation, design and experience. Computer Languages, 16(3-4):209–233, ???? 1991. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Stinckwich:2006:ISS

[SD06] Serge Stinckwich and Stéphane Ducasse. Introduction to the Smalltalk special issue. Computer Languages, Systems and Structures, 32(2–3):85–86, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784240500045X.

Sebesta:1989:CPG

[Seb89]  Robert W. Sebesta.  On

context-free programmed grammars. Computer Languages, 14 (2):99–108, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Szabo:1989:PAL

[SF89] M. E. Szabo and E. J. Farkas.

A probabilistic analysis of loop programs. Computer Languages, 14(2):125–136, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Soderberg:2015:DRT

[SH15] Emma Söderberg and Görel Hedin. Declarative rewriting through circular nonterminal attributes. Computer Languages, Systems and Structures, 44 (part A)(??):3–23, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842415000585.

Sharma:1975:SOP

[Sha75] O. P. Sharma. Syntax optimisation for and parsing of patterns. Computer Languages, 1 (3):233–253, September 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sharir:1980:SAN

[Sha80] M. Sharir. Structural analysis: a new approach to flow analysis in optimizing compilers. Computer Languages, 5(3-4):141–153, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sharir:1981:FIP

[Sha81]  Micha Sharir.  Formal in-

tegration: a program transformation technique. Computer Languages, 6(1):35–46, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sassa:2009:CEB

[SIK09] Masataka Sassa, Yo Ito, and Masaki Kohama. Comparison and evaluation of back-translation algorithms for static single assignment forms. Computer Languages, Systems and Structures, 35(2): 173–195, July 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842407000115.

Sistla:2004:ESR

[Sis04] A. Prasad Sistla. Employing symmetry reductions in model checking. Computer Languages, Systems and Structures, 30(3– 4):99–137, October/December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Sarwar:1994:ESR

[SJW94] S. Mansoor Sarwar, Mansour H. A. Jaragh, and Mike Wind. An empirical study of the run-time behavior of quicksort, Shellsort and mergesort for medium to large size data. Computer Languages, 20(2): 127–134, May 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Saritas:2014:MDA

[SK14] Hidayet Burak Saritas and Geylani Kardas. A model driven architecture for the development of smart card software. Computer Languages, Systems and Structures, 40(2): 53–72, ???? 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842414000025.

Sicak:2018:SSA

[SK18] Michal Sicák and Ján Kollár. Supercombinator set acquired from context-free grammar samples. Computer Languages, Systems and Structures, 54(??): 1–19, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301689.

Slivnik:2017:DLL

[Sli17] Bostjan Slivnik. On different LL and LR parsers used in LLLR parsing. Computer Languages, Systems and Structures, 50(??):108–126, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301853.

Syriani:2018:SMS

[SLS18] Eugene Syriani, Lechanceux Luhunu, and Houari Sahraoui. Systematic mapping study of template-based code generation. Computer Languages, Systems and Structures, 52(??): 43–62, June 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301239.

Sarbo:1989:TI

[SM89] Janos J. Sarbo and Matthias Moritz. Translator inversion. Computer Languages, 14(3): 205–224, ???? 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sailor:1994:PAT

[SM94] Ken Sailor and Carl McCrosky.

A practical approach to typesensitive parsing. Computer Languages, 20(2):101–116, May 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sanchez:2015:VAR

[SMB15] Alejandro Sanchez, Alexandre Madeira, and Luís S. Barbosa. On the verification of architectural reconfigurations. Computer Languages, Systems and Structures, 44 (Part C)(??):218– 237, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241500041X.

Sanchez:2016:VAR

[SMB16] Alejandro Sanchez, Alexandre Madeira, and Luís S. Barbosa. On the verification of architectural reconfigurations. Computer Languages, Systems and Structures, 45(??): 218–237, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241500041X.

Simao:2009:TLM

[SMdSB09] Adenilso Simão, José Carlos Maldonado, and Roberto da Silva Bigonha. A transformational language for mutant description. Computer Languages, Systems and Structures, 35(3):322–339, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000420.

Seipel:2018:DSL

[SNA18] Dietmar Seipel, Falco Nogatz, and Salvador Abreu. Domainspecific languages in Prolog for declarative expert knowledge in rules and ontologies. Computer Languages, Systems and Structures, 51(??):102– 117, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301804.

Sulir:2016:RCS

[SNP16] Matús Sulír, Milan Nosál, and Jaroslav Porubän. Recording concerns in source code using annotations. Computer Languages, Systems and Structures, 46(??):44–65, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241630015X.

Simonin:2018:AIC

[SP18] Jacques Simonin and John Puentes. Automatized integration of a contextual model into a process with data variability. Computer Languages, Systems and Structures, 54(??):156– 182, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241730177X.

Spragins:1979:ATM

[Spr79] John Spragins. Approximate techniques for modeling the performance of complex systems. Computer Languages, 4 (2):99–129, ???? 1979. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Silvestre:2010:FCE

[SRRB10] B. Silvestre, S. Rossetto, N. Rodriguez, and J.-P. Briot. Flexibility and coordination in event-based, loosely coupled, distributed systems. Computer Languages, Systems and Structures, 36(2):142– 157, July 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842409000220.

Struber:2017:TBV

[SRT17] Daniel Strüber, Felix Rieger, and Gabriele Taentzer. A text-based visual notation for the unit testing of modeldriven tools. Computer Languages, Systems and Structures, 49(??):196–215, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842416300276.

Schiffner:1979:MVA

[SS79] Gerd Schiffner and Peter Scheuermann. Multiple views and abstractions with an extended-entity-relationship model. Computer Languages, 4(3-4): 139–154, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sheard:1992:ITA

[SS92] Tim Sheard and David Stemple. Inheritance of theory in the ADABTPL language. Computer Languages, 17(3): 157–167, July 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Shekhar:1993:LSS

[SS93] K. H. Shekhar and Y. N. Srikant. LINDA sub system on Transputers. Computer Languages, 18(2):125–136, ????

1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schaeckeler:2009:OSS

[SS09] Stefan Schaeckeler and Weijia Shang. Optimizing the stack size of recursive functions. Computer Languages, Systems and Structures, 35(3):266– 276, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784240800016X.

Shenoy:1994:APF

[SSB94] U. Nagaraj Shenoy, Y. N.

Srikant, and V. P. Bhatkar. An automatic parallelization framework for multicomputers. Computer Languages, 20(3): 135–150, August 1994. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sarwar:1996:EQ

[SSJB96] S. Mansoor Sarwar, Syed Aqeel Sarwar, Mansour H. A. Jaragh, and Jesse Brandeburg. Engineering Quicksort. Computer Languages, 22(1):39–47, April 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sukumaran:2010:DCG

[SSM10] Srihari Sukumaran, Ashok Sreenivas, and Ravindra Metta. The dependence condition graph: Precise conditions for dependence between program points. Computer Languages, Systems and Structures, 36(1): 96–121, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784240900013X.

Seidl:2017:GSP

[SSS17] Christoph Seidl, Sven Schuster, and Ina Schaefer. Generative software product line development using variabilityaware design patterns. Computer Languages, Systems and Structures, 48(??):89– 111, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300609.

Stewart:1975:SES

[Ste75] S. L. Stewart. STAPLE, an experimental structured programming language. Computer Languages, 1(1):61–71, January 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Stetter:1984:MPC

[Ste84] Franz Stetter. A measure of program complexity. Computer Languages, 9(3-4):203– 208, ???? 1984. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sullivan:1975:EPS

[Sul75] J. E. Sullivan. Extending PL/1 for structured programming. Computer Languages, 1 (1):29–43, January 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Sutii:2018:EMR

[SvdBV18] Ana Maria Sutîi, Mark van den Brand, and Tom Verhoeff. Exploration of modularity and reusability of domainspecific languages: an expression DSL in MetaMod. Computer Languages, Systems and Structures, 51(??): 48–70, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417300404.

Smetsers:2016:DIH

[SvE16] Sjaak Smetsers and Marko van Eekelen. Derivation and inference of higher-order strictness types. Computer Languages, Systems and Structures, 44 (part B):166–180, ???? 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000512.

Saal:1977:ESA

[SW77] H. J. Saal and Z. Weiss. An empirical study of APL programs. Computer Languages, 2(3):47–59, ???? 1977. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Symes:1985:POC

[Sym85] D. Michael Symes. Procedural operators considered as fundamental programming devices. Computer Languages, 10 (2):75–89, ???? 1985. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Tai:1979:CFW

[Tai79] Kuo Chung Tai. Constant folding within an expression by semantic attributes. Computer Languages, 4(3-4):131– 137, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Talia:1993:DTC

[Tal93a] Domenico Talia. Distributed termination of concurrent processes in Occam. Computer Languages, 19(1):51–62, January 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Talia:1993:SPC

[Tal93b] Domenico Talia. A survey of PARLOG and Concurrent Prolog: the integration of logic and parallelism. Computer Languages, 18(3):185–196, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Taylor:1996:ARM

[Tay96] Hamish Taylor. Assembling a resolution multiprocessor from interface, programming and distributed processing components. Computer Languages, 22(2–3):181–192, July–October 1996. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Tourwe:2004:IIS

[TBKG04] Tom Tourwé, Johan Brichau, Andy Kellens, and Kris Gybels. Induced intentional software views. Computer Languages, Systems and Structures, 30(1– 2):35–47, April/July 2004. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Tsay:1981:DIC

[TC81] Jong Chuang Tsay and Yu Ming Chang. Design and implementation of a Chinese terminal controller. Computer Languages, 6(3-4):155–163, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Tennent:1983:SPI

[Ten83] R. D. Tennent. Some proposals for improving Pascal. Computer Languages, 8(3-4):125– 137, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Tharp:1977:CCF

[Tha77] A. L. Tharp. A comparison of COBOL, FORTRAN, PL-I and SPITBOL. Computer Languages, 2(4):171–178, ???? 1977. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Thimbleby:1982:TEI

[Thi82] Harold Thimbleby. A text editing interface: definition and use. Computer Languages, 7 (1):25–40, ???? 1982. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Thiemann:1993:OSR

[Thi93] Peter J. Thiemann. Optimizing structural recursion in functional programs. Computer Languages, 19(2):91–103, April 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Thirunarayan:1999:SMI

[TKH99] Krishnaprasad Thirunarayan, Günter Kniesel, and Haripriyan Hampapuram. Simulating multiple inheritance and generics in Java. Computer Languages, 25(4):189–210, December 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/28/30/24/
abstract.html.

Tremblay:2000:LEP

[TM00] G. Tremblay and B. Malenfant. Lenient evaluation and parallelism. Computer Languages, 26(1):27–41, April 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/18/50/25/26/
abstract.html.

Tremblay:2000:LEN

[Tre00] G. Tremblay. Lenient evaluation is neither strict nor lazy. Computer Languages, 26 (1):43–66, April 2000. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

/www.elsevier.com/gej-ng/
10/15/18/50/25/27/abstract.
html.

Torii:1987:LPT

[TSF^{+}87] Koji Torii, Yuji Sugiyama, Mamoru Fujii, Tadao Kasami, and Yoshitomi Morisawa. Logical programming for the telegram analysis problem. Computer Languages, 12(1):9–20, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Tucker:1975:VHL

[Tuc75] A. Tucker. Very high-level language design: a viewpoint. Computer Languages, 1(1):3– 16, January 1975. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Tzevelekos:2012:PES

[Tze12] Nikos Tzevelekos. Program equivalence in a simple language with state. Computer Languages, Systems and Structures, 38(2):181– 198, July 2012. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842412000036.

Ulgen:2015:IMA

[ÜA15] Onur Ülgen and Mutlu Avci. The intelligent memory allocator selector. Computer Languages, Systems and Structures, 44 (Part C)(??):342– 354, December 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300075.

Ulgen:2016:IMA

[ÜA16] Onur Ülgen and Mutlu Avci. The intelligent memory allocator selector. Computer Languages, Systems and Structures, 45(??):342–354, April 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1477842415300075.

Urban:2017:IRF

[UM17] Caterina Urban and Antoine Miné. Inference of ranking functions for proving temporal properties by abstract interpretation. Computer Languages, Systems and Structures, 47 (part 1)(??):77– 103, ???? 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000743.

Ururahy:2002:AFP

[URI02] C. Ururahy, N. Rodriguez, and R. Ierusalimschy. ALua: flexibility for parallel programming. Computer Languages, Systems and Structures, 28(2):155–180, December 15, 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Vainsencher:2004:MLB

[Vai04] Daniel Vainsencher. MudPie:

layers in the ball of mud. Computer Languages, Systems and Structures, 30(1–2):5–19, April/July 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Vidal:2016:ECJ

[VBDPM16] S. Vidal, A. Bergel, J. A.

Díaz-Pace, and C. Marcos. Over-exposed classes in Java: an empirical study. Computer Languages, Systems and Structures, 46(??):1–19, November 2016. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415300531.

VanCutsem:2014:APR

[VBS^{+}14] Tom Van Cutsem, Elisa Gonzalez Boix, Christophe Scholliers, Andoni Lombide Carreton, Dries Harnie, Kevin Pinte, and Wolfgang De Meuter. AmbientTalk: programming responsive mobile peer-to-peer applications with actors. Computer Languages, Systems and Structures, 40(3–4):112–136, October/December 2014. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842414000335.

Vacchi:2015:NFF

[VC15] Edoardo Vacchi and Walter Cazzola. Neverlang: a framework for feature-oriented language development. Computer Languages, Systems and Structures, 43(??):1–40, October 2015. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842415000056.

Virgilio:1982:BSS

[VF82] T. R. Virgilio and R. Finkel. Binding strategies and scope rules are independent. Computer Languages, 7(2):61–67, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Vialle:1998:DIP

[VLC98] Stéphane Vialle, Yannick Lallement, and Thierry Cornu. Design and implementation of a parallel cellular language for MIMD architectures. Computer Languages, 24(3):125– 153, October 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/19/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/19/17/article.
pdf.

VanCutsem:2009:LSB

[VMD09] Tom Van Cutsem, Stijn Mostinckx, and Wolfgang De Meuter. Linguistic symbiosis between event loop actors and threads. Computer Languages, Systems and Structures, 35(1): 80–98, April 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842408000249.

Vandercammen:2018:FFS

[VMD18] Maarten Vandercammen, Stefan Marr, and Coen De Roover. A flexible framework for studying trace-based justin-time compilation. Computer Languages, Systems and Structures, 51(??):22–47, January 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416302093.

vanOmmering:2001:LFV

[vOKF01] Rob van Ommering, René Krikhaar, and Loe Feijs. Languages for formalizing, visualizing and verifying software architectures. Computer Languages, 27(1–3): 3–18, April–October 2001. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL http:/

                                        /www.elsevier.com/gej-ng/
                                        10/15/18/54/27/28/abstract.
                                        html.

Venugopal:1993:HCD

[VS93] R. Venugopal and Y. N. Srikant. Heuristic chaining in directed acyclic graphs. Computer Languages, 19(3):169– 184, July 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Viswanathan:1994:PIL

[VS94] N. Viswanathan and Y. N. Srikant. Parallel incremental LR parsing. Computer Languages, 20(3):151–175, August 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Venugopal:1995:SET

[VS95] R. Venugopal and Y. N. Srikant. Scheduling expression trees with reusable registers on delayed-load architectures. Computer Languages, 21(1):49–65, April 1995. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

VanEs:2017:IPS

[VSN^{+}17] Noah Van Es, Quentin Stievenart, Jens Nicolay, Theo D’Hondt, and Coen De Roover. Implementing a performant Scheme interpreter for the web in asm.js. Computer Languages, Systems and Structures, 49(??):62–81, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300951.

Wadia:1980:GNL

[Wad80] Aspi B. Wadia. Generation of node lists using segment analysis. Computer Languages, 5 (2):115–129, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Walker:1989:FPI

[Wal89] Kenneth W. Walker. First-class patterns for Icon. Computer Languages, 14(3):153–163, ???? 1989. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wang:1992:CBL

[Wan92] Ching-Lin Wang. A continuationbased language embedded in Scheme. Computer Languages, 17(1):19–37, ???? 1992. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Warren:1978:TSN

[War78] Henry S. Warren, Jr. Three simple node list algorithms. Computer Languages, 3(2): 115–126, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wasserman:1979:TVA

[Was79] Anthony I. Wasserman. Testing and verification aspects of Pascal-Like languages. Computer Languages, 4(3-4):155– 169, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wu:2010:CBL

[WBGM10] Xiaoqing Wu, Barrett R. Bryant, Jeff Gray, and Marjan Mernik. Component-based LR parsing. Computer Languages, Systems and Structures, 36(1): 16–33, April 2010. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842409000037.

Wuyts:2004:UID

[WD04] Roel Wuyts and Stéphane Ducasse. Unanticipated integration of development tools using the classification model. Computer Languages, Systems and Structures, 30(1–2):63–77, April/July 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Wang:2008:PWI

[WDCL08] Shuangbao Wang, Zegang Dong, Jim X. Chen, and Robert S. Ledley. PPL: a whole-image processing language. Computer Languages, Systems and Structures, 34(1): 18–24, April 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842406000248.

Weiser:1985:EDD

[Wei85] Mark Weiser. Experience with a dataflow datatype. Computer Languages, 10(1):37–42, ???? 1985. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wetherell:1977:WAE

[Wet77] C. Wetherell. Why automatic error correctors fail. Computer Languages, 2(4):179–186, ???? 1977. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wand:1978:CLU

[WF78] Mitchell Wand and Daniel P.

Friedman. Compiling lambdaexpressions using continuations and factorizations. Computer Languages, 3(4):241–263, ???? 1978. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wampler:1983:RS

[WG83] Stephen B. Wampler and Ralph E. Griswold. Result sequences. Computer Languages, 8(1):1–14, ???? 1983. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

White:1977:EDP

[Whi77] J. E. White. Elements of a distributed programming system. Computer Languages, 2 (4):117–134, ???? 1977. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Wautelet:2017:USD

[WHKK17] Yves Wautelet, Samedi Heng, Soreangsey Kiv, and Manuel Kolp. User-story driven development of multi-agent systems: a process fragment for agile methods. Computer Languages, Systems and Structures, 50(??):159–176, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416301178.

Williams:1980:FNS

[Wil80] M. H. Williams. A formal notation for specifying static semantic rules. Computer Languages, 5(1):37–55, ???? 1980. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Williams:1981:MSS

[Wil81] Morgan Howard Williams. Methods for specifying static semantics. Computer Languages, 6(1):1–17, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Walters:2008:CRP

[WMP^{+}08] Edward K. Walters II, J. Eliot B. Moss, Trek Palmer, Timothy Richards, and Charles C.

Weems. CASL: a rapidprototyping language for modern micro-architectures. Computer Languages, Systems and Structures, 34(4):195–211, December 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784240700019X.

Wong:1999:OFP

[Won99] W. F. Wong. Optimizing floating point operations in Scheme. Computer Languages, 25(2): 89–112, July 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/28/27/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/28/27/26/article.
pdf.

Wilkinson:2006:PBM

[WPR06] Hernán Wilkinson, Máximo Prieto, and Luciano Romeo. A point based model of the Gregorian Calendar. Computer Languages, Systems and Structures, 32(2–3):157–183, July/October 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842405000424.

Yang:1996:MMB

[Yan96] Wuu Yang. Mealy machines are a better model of lexical analyzers. Computer Languages, 22(1):27–38, April 1996. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Yang:2000:FPA

[Yan00] Wuu Yang. A finest partitioning algorithm for attribute grammars. Computer Languages, 25(3):145–164, October 1, 2000. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/15/18/28/29/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/15/18/28/29/26/article.
                                        pdf.

Chen:1992:MMD

[yCH92] Jen yen Chen and Pei Hsia. MDL (Methodology Definition Language): a language for defining and automating software development processes. Computer Languages, 17(3): 199–211, July 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Yelowitz:1978:DSP

[YD78] Lawrence Yelowitz and Arthur G. Duncan. Data structures and program correctness: Bridging the gap. Computer Languages, 3(2):135–142, ???? 1978. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Yen:1983:DSM

[YF83] W. C. Yen and K. S. Fu. Distributed synchronization mechanism for interacting processes. Computer Languages, 8(2):51– 60, ???? 1983. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Yuen:1998:AOA

[YF98] C. K. Yuen and M. D.

Feng. Active objects as atomic control structures in BaLinda K. Computer Languages, 24(4):229–244, December 1, 1998. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/15/18/27/20/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/15/18/27/20/18/article.
pdf.

Yau:1993:CPS

[YG93] Stephen S. Yau and Jacob V. Gore. Constraint-driven programming in stronglytyped object-oriented languages. Computer Languages, 18(3):169–183, ???? 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Yang:2002:ALM

[YTC02] Wuu Yang, Chey-Woei Tsay, and Jien-Tsai Chan. On the applicability of the longest-match rule in lexical analysis. Computer Languages, Systems and Structures, 28(3):273–288, October 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Zaki:1987:FDA

[ZA87] M. Zaki and Al. H. Albarhamtoshy. Formal design of an Arabic text formatter for microcomputers. Computer Languages, 12(2):123–143, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zaki:1988:DGI

[Zak88] M. Zaki. Design of a graphics interface for computer-based biomedical applications. Computer Languages, 13(3-4):125– 141, ???? 1988. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zave:1986:CSP

[Zav86] Pamela Zave. Case study: The Paisley approach applied to its own software tools. Computer Languages, 11(1):15–28, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zolotas:2017:CPT

[ZCM^{+}17] Athanasios Zolotas, Robert Clarisó, Nicholas Matragkas, Dimitrios S. Kolovos, and Richard F. Paige. Constraint programming for type inference in flexible model-driven engineering. Computer Languages, Systems and Structures, 49(??):216–230, September 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842416300288.

Zdun:2006:TLB

[Zdu06] Uwe Zdun. Tailorable language for behavioral composition and configuration of software components. Computer Languages, Systems and Structures, 32(1): 56–82, April 2006. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Zaki:1985:PSA

[ZGE85] Mohamed Zaki and S. A.

Gamal-Eldin. A portable syntax analyzer for microcomputers. Computer Languages, 10 (2):127–146, ???? 1985. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zhao:2018:DID

[ZH18] Tian Zhao and Xiaobing Huang. Design and implementation of DeepDSL: a DSL for deep learning. Computer Languages, Systems and Structures, 54(??):39–70, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842418300319.

Zima:1986:CLI

[Zim86] Hans P. Zima. A constraint language and its interpreter. Computer Languages, 11(2):65–83, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zelkowitz:1981:ILE

[ZL81] Marvin V. Zelkowitz and James R. Lyle. Implementation of language enhancements. Computer Languages, 6(3-4): 139–153, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zhou:2018:SMA

[ZLZ^{+}18] Wenbo Zhou, Lei Liu, Peng Zhang, Shuai Lü, and Jingyao Li. SDAC: a model for analysis of the execution semantics of data processing framework in cloud. Computer Languages, Systems and Structures, 54(??):406–426, December 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842418300046.

Zobel:1993:PSB

[Zob93] Angelika Zobel. Program structure as basis for the parallelization of global register allocation. Computer Languages, 19 (2):135–155, April 1993. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zuck:2004:MCA

[ZP04] Lenore Zuck and Amir Pnueli. Model checking and abstraction to the aid of parameterized systems (a survey). Computer Languages, Systems and Structures, 30(3–4):139–169, October/December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Zhu:2018:TPS

[ZP18] Mengmeng Zhu and Hoang Pham. A two-phase software reliability modeling involving with software fault dependency and imperfect fault removal. Computer Languages, Systems and Structures, 53(??):27–42, September 2018. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1477842417301598.

Zouaoui:2017:CNG

[ZT17] Chakib Mustapha Anouar Zouaoui and Nasreddine Taleb. CL ARRAY: a new generic library of multidimensional containers for C++ compilers with extension for OpenCL framework. Computer Languages, Systems and Structures, 50(??):53–81, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S147784241630135X.

Zhou:2013:DSD

[ZTLM13] Wenchao Zhou, Tao Tao, Boon Thau Loo, and Yun Mao. Declarative secure distributed information systems. Computer Languages, Systems and Structures, 39(1): 1–24, April 2013. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1477842412000358.

Zuck:2004:SIV

[Zuc04] L. Zuck. Special issue of VMCAI’03. Computer Languages, Systems and Structures, 30(3–4):97–98, October/ December 2004. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).